### v1

# Global Dynamics of McKean-Vlasov SDEs via Stochastic Order

Abstract
This paper studies the rich dynamics of general one-dimensional McKean-Vlasov stochastic differential equations based on the preservation of stochastic order. The existing work is limited to the case of additive noises, while our framework allows multiplicative noises. When the equation has multiple invariant measures, we show basins of attraction of certain invariant measures contain unbounded open sets in the $2$-Wasserstein space, which is vacant in previous research even for additive noises. Also, our main results target on total orderedness and finiteness of invariant measures, global convergence to order interval enclosed by two order-related invariant measures, alternating arrangement of invariant measures in terms of stability (locally attracting) and instability (connecting orbits). Our theorems cover a wide range of classical granular media equations, such as symmetric, flat-bottom and asymmetric double-well confinement potentials with quadratic interaction, multi-well landscapes, and double-well landscapes with perturbations. Specific values for the parameter ranges, explicit descriptions of attracting sets and phase diagrams are provided.
MSC2020 subject classifications: Primary 60H10, 60B10; secondary 37C65, 60E15
Keywords: McKean-Vlasov SDEs, order-preserving semigroups, invariant measures, connecting orbits, basins of attraction, stochastic order
1 Introduction
This paper focuses on one-dimensional McKean-Vlasov processes governed by the following stochastic differential equation (SDE),
$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mathcal{L}(X_{t}))(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},$ (1.1)
where $\{B_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Brownian motion on $\mathbb{R}$, $\mathcal{L}(X_{t})$ is the law of the random variable $X_{t}$, $V:\mathbb{R}\to\mathbb{R}$ is a potential function, $W:\mathbb{R}\to\mathbb{R}$ is a local interaction, and $\sigma:\mathbb{R}\to\mathbb{R}$ is a diffusion term. The category of McKean-Vlasov SDEs refers to a class of SDEs whose coefficients depend on laws of solutions. McKean [52] introduced it as a generalization of Itô’s diffusion processes by considering a nonlinear version of Fokker-Planck equations. For (1.1), the corresponding Fokker-Planck partial differential equation (PDE) is given by
$\partial_{t}\rho_{t}=\frac{1}{2}\partial_{xx}\big{(}\sigma^{2}\rho_{t}\big{)}+\partial_{x}\big{[}(V^{\prime}+W^{\prime}*\rho_{t})\rho_{t}\big{]},$ (1.2)
where $\rho_{t}$ is the probability density function of $X_{t}$. After that, the aspect of particle approximation property was examined for McKean-Vlasov SDEs. The equation (1.1) can be deduced as the mean-filed limit (or hydrodynamic limit) of the following interacting particle system,
$\mathop{}\!\mathrm{d}X_{t}^{i,N}=-V^{\prime}\big{(}X_{t}^{i,N}\big{)}\mathop{}\!\mathrm{d}t-\frac{1}{N}\sum_{j=1}^{N}W^{\prime}\big{(}X_{t}^{i,N}-X_{t}^{j,N}\big{)}\mathop{}\!\mathrm{d}t+\sigma\big{(}X_{t}^{i,N}\big{)}\mathop{}\!\mathrm{d}B_{t}^{i},\quad i=1,2,\dots,N,$ (1.3)
where $\{B_{t}^{i}\}_{i=1}^{\infty}$ is a sequence of independent Brownian motions. As the particle number $N\to\infty$, a single particle $X_{t}^{i,N}$ or the empirical measure $\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{t}^{i,N}}$ converges to the continuum process $X_{t}$ in some sense. This is called “propagation of chaos”, and we refer to [21, 22, 53, 61] on this topic. Due to the origin from infinite-particle limits, McKean-Vlasov SDEs are natural mathematical models for large-scale phenomena. See [9] for galactic dynamics, [10] for plasma physics, and [59] for molecular and fluid dynamics. [11, 13, 45] use this kind of SDEs to model collective behaviour of different species. Another application on mathematical biology is [26, 28, 35, 49] to neuroscience. For optimal controls and differential games of McKean-Vlasov type, as well as study on machine learning and neural networks, see [15, 24, 70].
The equation (1.1) is a model for the time evolution of granular materials (see Benedetto-Caglioti-Pulvirenti [8]), so it is also called a granular media equation. A typical setting for the local interaction is attractive quadratic, namely
$W(x)=\frac{\theta}{2}x^{2},\quad\theta>0.$
As for the potential function, it has more choices ranging from double-well to multi-well shapes. Depending on the noise intensity and the number of critical points of the potential, the number of invariant measures of the equation (1.1) may vary. We refer to Dawson [27], Tugaut [66] for phase transitions on the number of invariant measures in double-well landscapes, and Alecio [1] for multi-well cases. Roughly speaking, if $V^{\prime}$ has $(2n-1)$ simple zeros, the McKean-Vlasov SDE (1.1) admits a unique invariant measure when $\sigma>\overline{\sigma}_{c}$ above some upper-critical value, and it admits $(2n-1)$ invariant measures when $\sigma<\underline{\sigma}_{c}$ below a lower-critical value. For more related work on phase transitions, see also [16, 29, 30].
Besides phase transitions on the number of invariant measures, convergence to invariant measures has been an active research direction. The study for the dynamics of (1.1) can be divided into two different cases according to uniqueness and non-uniqueness of invariant measures. If there is a unique invariant measure for (1.1), since the interacting particle system (1.3) also has a unique invariant measure, uniform-in-time propagation of chaos can give the convergence result. Malrieu [51] goes through this way for uniformly convex interaction terms, and obtains exponential convergence rate via a log-Sobolev inequality. Cattiaux-Guillin-Malrieu [20] relaxes to the non-uniformly convex case, and instead uses the method of synchronous couplings. We refer to [32, 46] and literature therein for more on uniform propagation of chaos, and [38] for uniform log-Sobolev inequalities of mean-filed particle systems.
Investigation of the associated PDE (1.2) is an alternative path to prove convergence to the unique invariant measure for (1.1). From this perspective, [5, 6] shows convergence to the unique invariant measure with vanishing potential ($V=0$) by Bakry-Émery’s $\Gamma$-calculus, while Benedetto-Caglioti-Carrillo-Pulvirenti [7] utilizes free energy as a Lyapunov functional when $V$ is convex. Actually, when $\sigma$ is a constant, the PDE (1.2) is a Wasserstein gradient flow of the free energy (see Ambrosio-Gigli-Savare [3]), which is a sum of entropy, potential energy and interaction energy,
$\mathcal{F}(\rho):=\frac{\sigma^{2}}{2}\int_{\mathbb{R}}\rho(x)\log\rho(x)\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}V(x)\rho(x)\mathop{}\!\mathrm{d}x+\frac{1}{2}\int_{\mathbb{R}\times\mathbb{R}}W(x-y)\rho(x)\rho(y)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y.$
By the “Riemannian structure” on Wasserstein spaces, also known as Otto’s calculus (see Otto [58]), free energy attains its global minimum at the unique invariant measure. Then nonlinear log-Sobolev inequalities, that link free energy and its dissipation, guarantee the exponential convergence. Carrillo-McCann-Villani [18] achieves this by checking Bakry-Émery criterion, and Carrillo-McCann-Villani [19] develops an abstract theory of gradient flows in length spaces to handle degenerate convexity. For other relevant work, see [4, 31] and refenrence therein for Bakry-Émery analysis, [23, 50] for log-Sobolev inequalties, and [17, 64] for entropy dissipation.
In the case of non-unique invariant measures, the dynamics is not as intuitive as the case of unique invariant measures. When the equation (1.1) has three invariant measures with a double-well potential, one of the predictions is local convergence. Now, propagation of chaos is hard to resort to, as the particle system (1.3) still has a unique invariant measure. Nonetheless, the gradient flow (1.2) remains workable, where free energy has two local minima and one local maximum, attained at three invariant measures respectively. Assuming qualitative smallness of the noise intensity $\sigma$, Tugaut [67] uses WJ-inequality to show the local convergence. See Bolley-Gentil-Guillin [12] for the derivation of WJ-inequalities and consequent spectral gaps in Wasserstein distance. Meanwhile, Monmarché-Reygner [56] executes a local version of log-Sobolev argument to obtain the local convergence in a sharp range of $\sigma$ for double-well landscapes. We refer to Tamura [62, 63] for early works using PDE techniques, and Cormier [25] by Lions derivatives. As for other directions than local convergence for the double-well case, the convergence to one of invariant measures is carried out by Tugaut [65], and on the other hand, Liu-Qu-Yao-Zhi [48] indicates that the central invariant measure is unstable and has connecting orbits to the other two.
The present paper is a continuation of [48], which establishes a general framework to study unstable dynamics in multi-dimensional cooperative McKean-Vlasov SDEs. In this paper, we investigate the richer dynamical behaviour of the one-dimensional equation (1.1) with non-unique invariant measures, from the stochastic-order-preserving perspective. For general case, Theorem 2.1 and Corollary 2.2 show the set of invariant measures is totally ordered and finite, and its infimum and supremum enclose a $\mathcal{P}_{2}$-compact order interval $[\underline{\nu},\overline{\nu}]$, which attracts any bounded set in the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$. The infimum $\underline{\nu}$ attracts all probability measures smaller than $\underline{\nu}$, and the supremum $\overline{\nu}$ attracts those greater than $\overline{\nu}$. Given the zero-crossing number of the derivative of the potential function, Theorem 2.3 and Corollary 2.4 unveil an alternating arrangement of invariant measures in terms of stability and instability, with stable ones $\underline{\nu}$, $\overline{\nu}$ at the beginning and the end. A stable invariant measure attracts an open neighbourhood, and an unstable invariant measure means there exists a connecting orbit away from it. We can further prove the infimum $\underline{\nu}$ and the supremum $\overline{\nu}$ attract unbounded open neighbourhoods in $\mathcal{P}_{2}(\mathbb{R})$. Additionally, Theorem 2.5-2.9 provide some applications of the main results to classical examples, and their phase diagrams are presented in Figure 1, 2.
Our approach is based on order-preservation of the semigroup generated by the equation (1.1) on the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$. To be precise, take an attractive quadratic interaction, and let $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ be the semigroup of (1.1). Consider the stochastic order “$\leq_{\mathrm{st}}$” defined by
$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\ \ \text{for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
Then $P_{t}^{*}$ is order-preserving on $\mathcal{P}_{2}(\mathbb{R})$ (see Theorem 4.4), that is
$\mu\leq_{\mathrm{st}}\nu\ \text{ implies }\ P_{t}^{*}\mu\leq_{\mathrm{st}}P_{t}^{*}\nu\ \text{ for $t\geq 0$}.$
Generally speaking, our method is parallel to the gradient flow approach. When the invariant measure is unique, to show global convergence, the gradient flow approach relies on the decreasing free energy along each single orbit, and a functional inequality involving free energy and its dissipation. By contrast, we develop the global attractor structures of order-preserving semigroups based on the characterization of order-boundedness and the coming down from infinity property. For the case of multiple invariant measures, the gradient flow approach needs a local functional inequality to show local convergence, while we establish the local attractor structures by tracing solutions backward along connecting orbits. To the best of our knowledge, the gradient flow approach concerning the basin of attraction of a single invariant measure is limited to the case of additive noises and local convergence, when invariant measures are non-unique. The main advantage of the order-preserving method is that we step further to give unbounded basins of attraction, even for multiplicative noises.
To illustrate that the basins of attraction of $\underline{\nu}$ (infimum invariant measure), $\overline{\nu}$ (supremum invariant measure) in Theorem 2.1 are unbounded, we exhibit inside a class of measures with explicit densities. Precisely, there are initial conditions arbitrarily far from $\underline{\nu}$ (resp. $\overline{\nu}$), for which the solutions converge to $\underline{\nu}$ (resp. $\overline{\nu}$). Consider a family of probability measures $\{\mu_{m}\}_{m\in\mathbb{R}}$ on $\mathbb{R}$, whose densities are given by
$\frac{\mathop{}\!\mathrm{d}\mu_{m}}{\mathop{}\!\mathrm{d}x}\propto\frac{1}{\sigma^{2}(x)}\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}.$ (1.4)
From a calculation of Fokker-Planck equations, the invariant measures of the equation (1.1) are those $\mu_{m}$ satisfying the following self-consistency equation,
$\int_{\mathbb{R}}(x-m)\mathop{}\!\mathrm{d}\mu_{m}(x)=0.$
Since $\mu_{m_{1}}\leq_{\mathrm{st}}\mu_{m_{2}}$ for $m_{1}\leq m_{2}$ (see Remark 4.3), it means $\underline{\nu}=\mu_{\underline{m}}$ and $\overline{\nu}=\mu_{\overline{m}}$ for some $\underline{m}\leq\overline{m}$. In this case, the solution starting from $\mu_{m}$ converges to $\underline{\nu}$ if $m\leq\underline{m}$, and converges to $\overline{\nu}$ if $m\geq\overline{m}$. Corollary 2.4 also gives a complete partition on limiting behaviour of solutions starting from the totally ordered arc $\{\mu_{m}\}_{m\in\mathbb{R}}$ in (1.4).
Our method can even show the basins of attraction contain unbounded open sets in $\mathcal{P}_{2}(\mathbb{R})$. By translating probability measures, we obtain order related pairs under the stochastic order. Given a local neighbourhood attracted by the infimum invariant measure, its basin of attraction actually contains all translated copies of the neighbourhood along the descending direction of the stochastic order. Correspondingly, the basin of attraction of the supremum invariant measure contains every ascending copy of its attracting neighbourhood. It is worth mentioning the stochastic order is inconsistent with the Wasserstein topology. In fact, any order interval has empty interior in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition A.1), and resembles a “thin slice” as shown in Figure 1, 2. So attracting open sets by the method of preserving such an unsatisfactory partial order relation is counter-intuitive.
As another necessary ingredient in our approach, we give the maximal number of invariant measures (see Proposition 5.3). Due to the analyticity and convolution form of the self-consistency function, we show the number of invariant measures of the equation (1.1) is finite and at most the zero-crossing number of the potential derivative $V^{\prime}$, regardless of the ranges of $\theta$ and $\sigma$. This helps to make our main theorems quantitative, and overcomes difficulties caused by qualitative results and missing parameter in the previous literature on phase transitions (see e.g., [1, 66]). One the other hand, this proposition allows $V^{\prime}$ vanishes on some interval, which is out of the settings in [67]. Also, it provides an insight that the essential factor in the maximal number of invariant measures is the zero-crossing number of $V^{\prime}$, instead of the number of simple zeros of $V^{\prime}$.
To demonstrate the versatility of our theory, we apply our main theorems to the following examples, including symmetric, flat-bottom, asymmetric double-well potentials, multi-well potentials, and double-well with perturbations (see Theorem 2.5-2.9). Take $W(x)=\frac{\theta}{2}x^{2}$ with $\theta>0$, and take the potential derivatives respectively as
[…]
In all these equations, all the dynamical results from Theorem 2.1, Corollary 2.2, Theorem 2.3 and Corollary 2.4 hold. The specific values for the ranges of $\theta$, $\sigma$, and explicit descriptions of attracting sets are given. For the phase diagrams, see Figure 1, 2.
This paper is organized as follows. In Section 2, we state our main results. Section 3 is about properties of the stochastic order, including an equivalent characterization of order boundedness in $\mathcal{P}_{2}(\mathbb{R})$ and a generalization of the lattice property. Section 4 is devoted to the proof of Theorem 2.1 and Corollary 2.2. We obtain the structure of global attractors of order-preserving semigroups, the comparison theorem for extrinsic-law-dependent SDEs, and the coming down from infinity property of McKean-Vlasov SDEs. Section 5 serves the proof of Theorem 2.3 and Corollary 2.4. We deduce the information on the number of invariant measures, the existence of positively invariant neighbourhoods of invariant measures, and the structure of local attractors of order-preserving semigroups. Also, the order bounds of limit sets are found by tracing solutions backward along connecting orbits. In Section 6, we apply the main theorems to examples stated in Theorem 2.5-2.9. In Appendix A, we show that, in $\mathcal{P}_{2}(\mathbb{R})$, any probability measure has no order relation with any open set, which denies the existence of strong ordered pairs for the stochastic order.
Notations
[…]
[type=symbols,style=mylong]
2 Assumptions and Main Results
In this section, we state our assumptions, some necessary definitions and main results.
Assumption 1.
(i) The continuous function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is one-sided Lipschitz, i.e., there exists $L>0$ such that, for all $x,y\in\mathbb{R}$,
$-(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq L\left\lvert x-y\right\rvert^{2}.$
(ii) The function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ has polynomial growth, i.e., there exist $L>0$, $\kappa\geq 1$ such that, for all $x\in\mathbb{R}$,
$\left\lvert V^{\prime}(x)\right\rvert\leq L(1+|x|^{\kappa}).$
(iii) The function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is hyper-dissipative, i.e., there exist $\alpha>0$, $\beta>0$, $\delta>0$ such that, for all $x,y\in\mathbb{R}$,
$-(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq-\alpha\left\lvert x-y\right\rvert^{2+\delta}+\beta.$
(iv) $W(x)=\frac{\theta}{2}x^{2}$ with $\theta>0$.
(v) The function $\sigma:\mathbb{R}\to\mathbb{R}$ is Lipschitz continuous and non-degenerate, i.e., there exist $0<\underline{\sigma}<\overline{\sigma}$, such that, for all $x\in\mathbb{R}$,
$\underline{\sigma}^{2}\leq\sigma^{2}(x)\leq\overline{\sigma}^{2}.$
Next, we emphasize the following definitions that are aforementioned.
Definition 2.1.
The stochastic order “$\leq_{\mathrm{st}}$” on $\mathcal{P}_{2}(\mathbb{R})$ is a partial order, defined by
$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\text{ for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
Definition 2.2.
The zero-crossing number of a continuous function $f:\mathbb{R}\to\mathbb{R}$, denoted by $Z(f)$, is the maximal integer $N$ such that, there exists an increasing sequence $x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}$ satisfying $f(x_{n})f(x_{n+1})<0$ for all $n=1,2,\dots,N$.
Definition 2.3.
The equation (1.1) is called locally dissipative at $a\in\mathbb{R}$ with configuration $(r,g)$, if
  1. (i)
    the function $g:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}$ satisfies, for all $x\in\mathbb{R}$, $\mu\in\mathcal{P}_{2}(\mathbb{R})$,
    $-2xV^{\prime}(x+a)-2x(W^{\prime}*\mu)(x)+\sigma^{2}(x+a)\leq-g(x^{2},\|\mu\|_{2}^{2});$
  2. (ii)
    the number $r>0$ and the function $g$ satisfy
    $\begin{split}&g(\cdot,r^{2})\text{ is continuous and convex};\\ &\inf_{0\leq w\leq r^{2}}g(z,w)=g(z,r^{2});\\ &\inf_{z\geq r^{2}}g(z,r^{2})>0.\end{split}$
Definition 2.4.
For a semigroup $P^{\ast}_{t}$ on $\mathcal{P}_{2}(\mathbb{R})$,
  1. (i)
    a measure $\nu\in\mathcal{P}_{2}(\mathbb{R})$ is called an invariant measure if $P_{t}^{*}\nu=\nu$ for all $t\geq 0$;
  2. (ii)
    a family $\{\nu_{t}\}_{t\in\mathbb{R}}\subset\mathcal{P}_{2}(\mathbb{R})$ is called a connecting orbit from an invariant measure $\nu$ to another invariant measure $\tilde{\nu}$ if
    $P^{\ast}_{t}\nu_{s}=\nu_{t+s}\text{ for any $t\geq 0$, $s\in\mathbb{R}$};$
    $\mathcal{W}_{2}(\nu_{t},\nu)\to 0\text{ as $t\to-\infty$ and }\mathcal{W}_{2}(\nu_{t},\tilde{\nu})\to 0\text{ as $t\to\infty$}.$
    Furthermore, if
    $\nu_{s}\leq_{\mathrm{st}}\nu_{t}\ \ (\nu_{s}\geq_{\mathrm{st}}\nu_{t})\text{ for all }s\leq t,$
    then we call $\{\nu_{t}\}_{t\in\mathbb{R}}$ an increasing (decreasing) connecting orbit from $\nu$ to $\tilde{\nu}$.
Remark 2.1.
A family $\{\nu_{t}\}_{t\in\mathbb{R}}\subset\mathcal{P}_{2}(\mathbb{R})$ is called an entrance measure if $P_{t}^{*}\nu_{s}=\nu_{t+s}$ for all $t\geq 0$, $s\in\mathbb{R}$, and particularly, a connecting orbit from $\nu$ to $\tilde{\nu}$ is an entrance measure.
Now we state our main theorems. Theorem 2.1, 2.3 and Corollary 2.2, 2.4 are abstract results for the equation (1.1), while Theorem 2.5-2.9 deal with concrete equations. In the sequel, we denote $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by the semigroup generated by the equation (1.1), see also (4.5).
Theorem 2.1.
Under Assumption 1, the following statements hold for the equation (1.1).
  1. (i)
    (Total Orderedness and Finiteness) The set of invariant measures $\mathcal{M}$ is a finite set, totally ordered with respect to the stochastic order, and $\mathcal{M}\subset\mathcal{P}_{\infty}(\mathbb{R})$;
  2. (ii)
    (Global Convergence to Order Interval) $\inf_{\leq_{\mathrm{st}}}\mathcal{M}$ and $\sup_{\leq_{\mathrm{st}}}\mathcal{M}$ exist. We denote them by $\underline{\nu}$ and $\overline{\nu}$ respectively. Then for all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$,
    $\mathcal{W}_{2}(P_{t}^{*}B,[\underline{\nu},\overline{\nu}])\to 0\ \text{ as $t\to\infty$};$
  3. (iii)
    (Globally Attracting from Below/Above) For all $\mu\leq_{\mathrm{st}}\underline{\nu}$,
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\underline{\nu})\to 0\text{ as }t\to\infty,$
    and for all $\mu\geq_{\mathrm{st}}\overline{\nu}$,
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\overline{\nu})\to 0\text{ as }t\to\infty.$
Corollary 2.2.
(Global Convergence) Under Assumption 1, if the equation (1.1) has a unique invariant measure $\nu$, then for all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,\nu)\to 0$ as $t\to\infty$.
Theorem 2.3.
Suppose Assumption 1 holds, and suppose the zero-crossing number $Z(V^{\prime})$ of $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is $(2n-1)$. If the equation (1.1) is locally dissipative at $a_{1}<a_{2}<\cdots<a_{n}$ with configurations $\{(r_{k},g_{k})\}_{k=1}^{n}$, and if
$r_{k}^{2}+r_{k+1}^{2}\leq\frac{1}{2}(a_{k+1}-a_{k})^{2}\ \text{ for all $k=1,2,\dots,n-1$},$ (2.1)
then
  1. (i)
    (Total Orderedness) there are exactly $(2n-1)$ invariant measures $\{\nu_{k}\}_{k=1}^{2n-1}$ with $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$;
  2. (ii)
    (Positively Invariant Open Neighbourhoods) the open balls $\{B(\delta_{a_{k}},r_{k})\}_{k=1}^{n}$ are pairwise disjoint, and there hold
    $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k}),\ \text{ for all }k=1,2,\dots,n,$ (2.2)
    and
    $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}\bigcup\overline{B(\delta_{a_{k+1}},r_{k+1})},\ \text{ for all }k=1,2,\dots,n-1.$ (2.3)
    Moreover, $P^{*}_{t}B(\delta_{a_{k}},r_{k})\subset B(\delta_{a_{k}},r_{k})$ for all $k=1,2,\dots,n$ and $t\geq 0$;
  3. (iii)
    (Connecting Orbits) for all $k=1,2,\dots,n-1$, there are a decreasing connecting orbit $\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$ from $\nu_{2k}$ to $\nu_{2k-1}$ and an increasing connecting orbit $\{\nu^{k,\uparrow}_{t}\}_{t\in\mathbb{R}}$ from $\nu_{2k}$ to $\nu_{2k+1}$;
  4. (iv)
    (Locally Attracting) for all $k=1,2,\dots,n$,
    $\mathcal{W}_{2}(P_{t}^{*}B(\delta_{a_{k}},r_{k}),\nu_{2k-1})=\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0\text{ as }t\to\infty;$
  5. (v)
    (Attracting Unbounded Open Neighbourhoods) if $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r_{1})$, then
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\text{ as }t\to\infty.$
    If $\mu\geq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{n}},r_{n})$, then
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0\text{ as }t\to\infty.$
    In particular,
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\text{ as }t\to\infty\ \text{ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{a_{1}+a},r_{1})$},$
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0\text{ as }t\to\infty\ \text{ forall $\mu\in\bigcup_{a\geq 0}B(\delta_{a_{n}+a},r_{n})$}.$
Remark 2.2.
(i). By Assumption 1 (i)(iii), the function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ must have an odd zero-crossing number whenever finite.
(ii). By definition, a connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\nu$ to $\tilde{\nu}$ as a bounded set itself cannot be attracted by its forward limit $\tilde{\nu}$, i.e., $\mathcal{W}_{2}(P_{t}^{*}\{\nu_{s}:s\in\mathbb{R}\},\tilde{\nu})\nrightarrow 0$, although $\mathcal{W}_{2}(P_{t}^{*}\nu_{s},\tilde{\nu})\to 0$ as $t\to\infty$ for any $s\in\mathbb{R}$. So, the set of invariant measures cannot attract all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. In other word, the order interval $[\underline{\nu},\overline{\nu}]$ in Theorem 2.1 (ii) cannot be replaced by the set of invariant measures.
(iii). By the alternating arrangement of the invariant measures in Theorem 2.3, we have more information on the basins of attraction (see Theorem 4.2).
  • if $\mu<_{\mathrm{st}}\nu^{1,\downarrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$ as $t\to\infty$;
  • if $\nu^{k-1,\uparrow}_{t_{0}}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu^{k,\downarrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, $k=2,3,\dots,n-1$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$;
  • if $\mu>_{\mathrm{st}}\nu^{n-1,\uparrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0$ as $t\to\infty$.
Corollary 2.4.
Suppose the assumptions in Theorem 2.3 hold, and let $\{\mu_{m}\}_{m\in\mathbb{R}}$ be probability measures given in (1.4). Then
  1. (i)
    there are increasing values $m_{k}$, $k=1,2,\dots,2n-1$, such that $\nu_{k}=\mu_{m_{k}}$;
  2. (ii)
    set $m_{0}=-\infty$ and $m_{2n}=\infty$. For all $k=1,2,\dots,n$, and for all $m\in(m_{2k-2},m_{2k})$,
    $\mathcal{W}_{2}(P_{t}^{*}\mu_{m},\nu_{2k-1})\to 0\text{ as }t\to\infty.$
Remark 2.3.
(i). For each $k=1,2,\dots,2n-2$, although the set $\{\mu_{m}\}_{m\in(m_{k},m_{k+1})}$ is a totally ordered arc with endpoints $\nu_{k}$, $\nu_{k+1}$, it is not a connecting orbit between invariant measures. One can check it from the Fokker-Planck equation.
(ii). Corollary 2.4 also provides more information on basins of attraction. For each $k=1,2,\dots,n$, if $\mu_{m}\leq_{\mathrm{st}}\mu\leq_{\mathrm{st}}\mu_{\tilde{m}}$ for some $m_{2k-2}<m<\tilde{m}<m_{2k}$, then by the order-preservation of $P^{*}$ and the squeezing principle, we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$.
Besides abstract results, we have the following applications.
Theorem 2.5 (Double-well landscapes, Figure 1).
Consider the following one-dimensional McKean-Vlasov SDE,
$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+1)+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.4)
If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{51\sqrt{17}-107}{256}\ \text{ for all $x\in\mathbb{R}$},$ (2.5)
and let $r=\frac{9-\sqrt{17}}{8}$, then the following statements are true.
  1. (i)
    There are exactly three invariant measures, $\nu_{-1},\nu_{0},\nu_{1}\in\mathcal{P}_{\infty}(\mathbb{R})$ with $\nu_{-1}<_{\mathrm{st}}\nu_{0}<_{\mathrm{st}}\nu_{1}$;
  2. (ii)
    For all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\nu_{-1},\nu_{1}])\to 0$ as $t\to\infty$;
  3. (iii)
    There are a decreasing connecting orbit from $\nu_{0}$ to $\nu_{-1}$ and an increasing connecting orbit from $\nu_{0}$ to $\nu_{1}$;
  4. (iv)
    There hold $\nu_{-1}\in B(\delta_{-1},r)$, $\nu_{1}\in B(\delta_{1},r)$, and
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{-1})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{-1+a},r)$},$
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{1+a},r)$}.$
Figure 1: Phase diagram for double-well landscapes
Theorem 2.6 (Double-well with flat bottoms, Figure 1).
Consider the following one-dimensional McKean-Vlasov SDE,
$\mathop{}\!\mathrm{d}X_{t}=-\left[V^{\prime}(X_{t})+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.6)
where
$V^{\prime}(x)=\begin{cases}x\big{(}x-\frac{31}{32}\big{)}\big{(}x+\frac{31}{32}\big{)},&x\in\big{(}-\frac{31}{32},\frac{31}{32}\big{)},\\ 0,&x\in\big{(}-1,-\frac{31}{32}\big{]}\cup\big{[}\frac{31}{32},1\big{)},\\ x(x-1)(x+1),&x\in(-\infty,-1]\cup[1,\infty).\end{cases}$
If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{51\sqrt{17}-171}{256}\ \text{ for all $x\in\mathbb{R}$},$ (2.7)
and let $r=\frac{9-\sqrt{17}}{8}$, then the exactly same results as in Theorem 2.5 are valid.
Theorem 2.7 (Asymmetric double-well, Figure 1).
Consider the following one-dimensional McKean-Vlasov SDE,
$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+2)+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.8)
If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
$\theta\geq\frac{4(3+\sqrt{3})}{3},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{3(39+55\sqrt{33})}{256}\ \text{ for all $x\in\mathbb{R}$},$
and let $r_{-2}=\frac{(\sqrt{33}-15)^{2}}{64}$, $r_{1}=\frac{6-3\sqrt{3}}{2}$, then the following statements are true.
  1. (i)
    There are exactly three invariant measures, $\nu_{-2},\nu_{0},\nu_{1}\in\mathcal{P}_{\infty}(\mathbb{R})$ with $\nu_{-2}<_{\mathrm{st}}\nu_{0}<_{\mathrm{st}}\nu_{1}$;
  2. (ii)
    For all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\nu_{-2},\nu_{1}])\to 0$ as $t\to\infty$;
  3. (iii)
    There are a decreasing connecting orbit from $\nu_{0}$ to $\nu_{-2}$ and an increasing connecting orbit from $\nu_{0}$ to $\nu_{1}$;
  4. (iv)
    There hold $\nu_{-2}\in B(\delta_{-2},r_{-2})$, $\nu_{1}\in B(\delta_{1},r_{1})$, and
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{-2})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{-2+a},r_{-2})$},$
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{1+a},r_{1})$}.$
Figure 2: Phase diagram for multi-well landscapes
Theorem 2.8 (Multi-well landscapes, Figure 2).
Consider the following one-dimensional McKean-Vlasov SDE,
$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+1)(X_{t}-2)(X_{t}+2)+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.9)
If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
$\theta\geq 8\sqrt{5+\sqrt{13}},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{4(13\sqrt{13}-35)}{27}\ \text{ for all $x\in\mathbb{R}$},$
and let $r=\frac{5-\sqrt{13}}{3}$, then the following statements are true.
  1. (i)
    There are exactly five invariant measures, $\nu_{-2},\nu_{-1},\nu_{0},\nu_{1},\nu_{2}\in\mathcal{P}_{\infty}(\mathbb{R})$ with $\nu_{-2}<_{\mathrm{st}}\nu_{-1}<_{\mathrm{st}}\nu_{0}<_{\mathrm{st}}\nu_{1}<_{\mathrm{st}}\nu_{2}$;
  2. (ii)
    For all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\nu_{-2},\nu_{2}])\to 0$ as $t\to\infty$;
  3. (iii)
    There are a decreasing connecting orbit from $\nu_{-1}$ to $\nu_{-2}$ and an increasing connecting orbit from $\nu_{-1}$ to $\nu_{0}$;
    there are a decreasing connecting orbit from $\nu_{1}$ to $\nu_{0}$ and an increasing connecting orbit from $\nu_{1}$ to $\nu_{2}$;
  4. (iv)
    There hold $\nu_{-2}\in B(\delta_{-2},r)$, $\nu_{0}\in B(\delta_{0},r)$, $\nu_{2}\in B(\delta_{2},r)$, and
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{-2})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{-2+a},r)$},$
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{0})\to 0\ \text{ as $t\to\infty$ for all $\mu\in B(\delta_{0},r)$},$
    and
    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{2+a},r)$}.$
Theorem 2.9 (Double-well with perturbation).
Consider the following one-dimensional McKean-Vlasov SDE,
$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+1)+f(X_{t})+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.10)
If the parameter $\theta$ and the Lipschitz functions $f\colon\mathbb{R}\to\mathbb{R}$, $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
$\theta\geq 0,\quad\left\lvert f(x)\right\rvert\leq L,\ \ 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\overline{\sigma}^{2}\ \text{ for all $x\in\mathbb{R}$},$
then the exactly same results as in Theorem 2.1 are valid.
3 Properties of Stochastic Order
In this section, some new properties of the stochastic order are examined. We first summarize some preliminary materials involving stochastic order in $\mathcal{P}_{2}(\mathbb{R})$. Then we give an equivalent characterization of order boundedness in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition 3.2), which plays a crucial role in the proof of our main results. As a consequence, we show any bounded subset of $\mathcal{P}_{p}(\mathbb{R})$, $p>2$, is order bounded in $\mathcal{P}_{2}(\mathbb{R})$ (see Corollary 3.3), and any order bounded subset of $\mathcal{P}_{2}(\mathbb{R})$ is relatively compact (see Proposition 3.4).
3.1 Preliminaries of stochastic order
Let $\mathcal{P}(\mathbb{R})$ be the set of probability measures on $\mathbb{R}$, and let $\mathcal{P}_{p}(\mathbb{R})$ be the set of those having finite $p$-th moments,
$\mathcal{P}_{p}(\mathbb{R}):=\left\{\mu\in\mathcal{P}(\mathbb{R}):\left\lVert\mu\right\rVert_{p}=\left(\int_{\mathbb{R}}\left\lvert x\right\rvert^{p}\mathop{}\!\mathrm{d}\mu(x)\right)^{1/p}<\infty\right\}.$
For $p\in[1,\infty)$, equip $\mathcal{P}_{p}(\mathbb{R})$ with the $p$-Wasserstein metric defined as below,
$\mathcal{W}_{p}(\mu,\nu):=\inf\left\{\left(\mathbf{E}[\left\lvert X-Y\right\rvert^{p}]\right)^{1/p}:\mathcal{L}(X)=\mu,\,\mathcal{L}(Y)=\nu\right\},$
where $\mathcal{L}(\cdot)$ represents the law of a random variable. Then $(\mathcal{P}_{p}(\mathbb{R}),\mathcal{W}_{p})$ is a complete metric space (see Villani [68, Theorem 6.18]).
Next, we review the so-called “stochastic order”, which is also known as stochastic domination. By Definition 2.1, for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$,
$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\text{ for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
Then “$\leq_{\mathrm{st}}$” is a partial order on $\mathcal{P}_{2}(\mathbb{R})$. Indeed, the reflexivity and transitivity are obvious, and the antisymmetry can be obtained by taking $f=\mathbf{1}_{(x,\infty)}$ for any $x\in\mathbb{R}$. Moreover, as shown in [48, Lemma 3.2], the stochastic order is a closed partial order relation in $\mathcal{P}_{2}(\mathbb{R})$, namely, the subset $\{(\mu,\nu):\mu\leq_{\mathrm{st}}\nu\}$ is closed in the product space $\mathcal{P}_{2}(\mathbb{R})\times\mathcal{P}_{2}(\mathbb{R})$. That is to say, $\mu\leq_{\mathrm{st}}\nu$ whenever $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ and $\mathcal{W}_{2}(\nu_{n},\nu)\to 0$ as $t\to\infty$ and $\mu_{n}\leq_{\mathrm{st}}\nu_{n}$ for all $n$.
If $\mu\leq_{\mathrm{st}}\nu$ and $\mu\neq\nu$, we write $\mu<_{\mathrm{st}}\nu$. The reversed signs $\geq_{\mathrm{st}}$, $>_{\mathrm{st}}$ are used in the usual way. For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$ with $\mu\leq_{\mathrm{st}}\nu$, the order interval is defined by $[\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\mu\leq_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$, $(\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\mu<_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$ and similarly we can define $[\mu,\nu)$ and $(\mu,\nu)$. Measures $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$ are said to be order related, if $\mu\leq_{\mathrm{st}}\nu$ or $\nu\leq_{\mathrm{st}}\mu$ holds. Otherwise, they are order unrelated. Given two subsets $A$ and $B$ of $\mathcal{P}_{2}(\mathbb{R})$, we write $A\leq_{\mathrm{st}}B$ ($A<_{\mathrm{st}}B$) when $\mu\leq_{\mathrm{st}}\nu$ ($\mu<_{\mathrm{st}}\nu$) holds for each choice of $\mu\in A$ and $\nu\in B$. Sets $A,B\subset\mathcal{P}_{2}(\mathbb{R})$ are said to be order related, if $A\leq_{\mathrm{st}}B$ or $B\leq_{\mathrm{st}}A$ holds. A totally ordered set (which is also referred as chain) $A$ in $\mathcal{P}_{2}(\mathbb{R})$ means that, any two measures $\mu$ and $\nu$ in $A$ are order related.
A measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$ is an upper bound (lower bound) of a set $A\subset\mathcal{P}_{2}(\mathbb{R})$ if $\mu\geq_{\mathrm{st}}A$ ($\mu\leq_{\mathrm{st}}A$). If the set $A$ has an upper (lower) bound, it is said to be bounded from above (below). A set $A$ which is bounded from both above and below is said to be order bounded. An upper bound $\mu_{0}$ of $A$ is said to be the supremum of $A$, denoted by $\mu_{0}=\sup_{\leq_{\mathrm{st}}}A$, if any other upper bound $\mu$ satisfies $\mu\geq_{\mathrm{st}}\mu_{0}$. Similarly, a lower bound $\mu_{0}$ of $A$ is said to be the infimum of $A$, denoted by $\mu_{0}=\inf_{\leq_{\mathrm{st}}}A$, if any other lower bound $\mu$ satisfies $\mu\leq_{\mathrm{st}}\mu_{0}$. By the antisymmetry of the stochastic order, if the supremum (infimum) exists, it must be unique.
There are some complementary material of the stochastic order that we will use later. Firstly, a useful equivalent characterization of the stochastic order is Strassen’s theorem (see e.g., Lindvall [47, equation (3)]), which says $\mu\leq_{\mathrm{st}}\nu$ if and only if there exist two random variables $X,Y$ such that
$\mathcal{L}(X)=\mu,\quad\mathcal{L}(Y)=\nu,\quad X\leq Y\ \mathrm{a.s.}$
Secondly, according to Hiai-Lawson-Lim [40, Proposition 3.11], test functions $f$ in Definition 2.1 can be restricted as bounded continuous increasing. If $\mu$, $\nu$ both have finite first moments, it is equivalent for test functions ranging over all $1$-Lipschitz increasing functions (see Fritz-Perrone [36, Theorem 4.2.1]), where a $1$-Lipschitz function $f$ means $\left\lvert f(x)-f(y)\right\rvert\leq\left\lvert x-y\right\rvert$ for all $x,y\in\mathbb{R}$. We may confine test functions in such different sets whenever convenient. Additionally, an important observation regarding the definition of $\mu\leq_{\mathrm{st}}\nu$, is that when we take the test function as $f=-\mathbf{1}_{(-\infty,x]}$ or $f=\mathbf{1}_{(x,\infty)}$, we have $\mu\big{(}(-\infty,x]\big{)}\geq\nu\big{(}(-\infty,x]\big{)}$ and $\mu\big{(}(x,\infty)\big{)}\leq\nu\big{(}(x,\infty)\big{)}$, for any $x\in\mathbb{R}$.
3.2 Equivalent characterization of order boundedness
The following lemma gives an equivalent characterization of the stochastic order by the comparison of cumulative distribution functions.
Lemma 3.1.
Let $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R})$ be two probability measures on $\mathbb{R}$, and let $F_{1},F_{2}:\mathbb{R}\to[0,1]$ be their cumulative distribution functions. Then $\mu_{1}\leq_{\mathrm{st}}\mu_{2}$ if and only if $F_{1}(x)\geq F_{2}(x)$ for all $x\in\mathbb{R}$.
Proof.
($\Rightarrow$) Note that $F_{i}(x)=\mu_{i}\big{(}(-\infty,x]\big{)}$, $i=1,2$. Take the test function $f=-\mathbf{1}_{(-\infty,x]}$ in the definition of $\mu_{1}\leq_{\mathrm{st}}\mu_{2}$, it follows that $F_{1}(x)\geq F_{2}(x)$, for any $x\in\mathbb{R}$.
($\Leftarrow$) By Hiai-Lawson-Lim [40, Proposition 3.11], it suffices to show $\int f\mathop{}\!\mathrm{d}\mu_{1}\leq\int f\mathop{}\!\mathrm{d}\mu_{2}$ for all bounded continuous increasing functions $f$. Let $f:\mathbb{R}\to\mathbb{R}$ be such a function, and we may assume $f$ is non-negative. Set
$f_{n}:=\sum_{k=1}^{n2^{n}}\frac{1}{2^{n}}\mathbf{1}_{\{f>k/2^{n}\}},$
and we have $f_{n}\uparrow f$ as $n\to\infty$. Note that the set $\{f>k/2^{n}\}$ is an open interval of the form $(x_{k},\infty)$, where (,):=assign($\infty$,$\infty$):$\emptyset$( ∞ , ∞ ) := ∅, since $f$ is increasing continuous. So we know that
$\int f_{n}\mathop{}\!\mathrm{d}\mu_{i}=\sum_{k=1}^{n2^{n}}\frac{1}{2^{n}}\mu_{i}\big{(}(x_{k},\infty)\big{)},\ \ i=1,2.$
By the assumption that $F_{1}(x)\geq F_{2}(x)$ for all $x\in\mathbb{R}$, we conclude that for any $x\in\mathbb{R}$,
$\mu_{1}\big{(}(x,\infty)\big{)}=1-F_{1}(x)\leq 1-F_{2}(x)=\mu_{2}\big{(}(x,\infty)\big{)}.$
Then
$\int_{\mathbb{R}}f_{n}\mathop{}\!\mathrm{d}\mu_{1}\leq\int_{\mathbb{R}}f_{n}\mathop{}\!\mathrm{d}\mu_{2}\quad\text{ for all $n\geq 1$}.$
The monotone convergence theorem gives $\int f\mathop{}\!\mathrm{d}\mu_{1}\leq\int f\mathop{}\!\mathrm{d}\mu_{2}$. ∎
Remark 3.1.
Lemma 3.1 implies the lattice property holds in $\mathcal{P}_{2}(\mathbb{R})$, that is to say, $\sup_{\leq_{\mathrm{st}}}\{\mu_{1},\mu_{2}\}$ exists for any two probability measures $\mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R})$. However, there is no lattice property in $\mathcal{P}_{2}(\mathbb{R}^{d})$ for $d\geq 2$. See counterexamples in Kamae-Krengel-O’Brien [44] or Müller-Scarsini [57].
We next give an equivalent characterization of order boundedness in $\mathcal{P}_{2}(\mathbb{R})$, which turns out to be very crucial for our approach.
Proposition 3.2.
Let $M$ be a subset of $\mathcal{P}_{2}(\mathbb{R})$. Then the following statements are true.
  1. (i)
    $M$ is bounded from above in $\mathcal{P}_{2}(\mathbb{R})$ with respect to the stochastic order if and only if
    $\int_{0}^{\infty}\bigg{(}\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u<\infty.$ (3.1)
  2. (ii)
    $M$ is bounded from below in $\mathcal{P}_{2}(\mathbb{R})$ with respect to the stochastic order if and only if
    $\int_{0}^{\infty}\bigg{(}\sup_{\mu\in M}\mu\big{(}(-\infty,-u]\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u<\infty.$ (3.2)
Proof.
We only prove (i), and the proof of (ii) is in a similar fashion.
($\Rightarrow$) Suppose $M\leq_{\mathrm{st}}\overline{\mu}$ for some $\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R})$. Then by the definition of stochastic order, we have for all $u\in\mathbb{R}$,
$\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\leq\overline{\mu}\big{(}(u,\infty)\big{)}.$
So, it follows
$\int_{0}^{\infty}\bigg{(}\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u\leq\int_{0}^{\infty}\overline{\mu}\big{(}(u,\infty)\big{)}u\mathop{}\!\mathrm{d}u=\int_{0}^{\infty}\int_{\mathbb{R}}\mathbf{1}_{\{x>u\}}\mathop{}\!\mathrm{d}\overline{\mu}(x)u\mathop{}\!\mathrm{d}u.$ (3.3)
By Tonelli’s theorem, we switch the integrals and obtain
$\int_{0}^{\infty}\int_{\mathbb{R}}\mathbf{1}_{\{x>u\}}\mathop{}\!\mathrm{d}\overline{\mu}(x)u\mathop{}\!\mathrm{d}u=\int_{\mathbb{R}}\mathbf{1}_{\{x>0\}}\int_{0}^{x}u\mathop{}\!\mathrm{d}u\mathop{}\!\mathrm{d}\overline{\mu}(x)=\frac{1}{2}\int_{\mathbb{R}}x^{2}\mathbf{1}_{\{x>0\}}\mathop{}\!\mathrm{d}\overline{\mu}(x)<\infty.$ (3.4)
Combining (3.3) and (3.4) finishes the “only if” part.
($\Leftarrow$) Note $\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}$ is decreasing in $u$, and then by condition (3.1), we have
$\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\to 0\ \text{ as $u\to\infty$}.$ (3.5)
Let $F_{\mu}$ be the cumulative distribution function of $\mu$, and define a function $F:\mathbb{R}\to[0,1]$ by
$F(x):=\inf_{\mu\in M}F_{\mu}(x).$
We want to show $F$ is also a cumulative distribution function. It is obvious that $F$ is increasing, and $\lim_{x\to-\infty}F(x)=0$. Also, (3.5) gives $\lim_{x\to\infty}F(x)=1$. Since $F_{\mu}$ is right continuous for every $\mu\in M$, we claim $F$ is also right continuous. In fact, by the monotonicity of $F$ and right continuity of $F_{\mu}$, we have
$F(x)\leq\lim_{x_{n}\downarrow x}F(x_{n})=\lim_{x_{n}\downarrow x}\inf_{\mu\in M}F_{\mu}(x_{n})\leq\inf_{\mu\in M}\lim_{x_{n}\downarrow x}F_{\mu}(x_{n})=\inf_{\mu\in M}F_{\mu}(x)=F(x).$
Thus, $F$ is a cumulative distribution function of some probability measure $\nu$. Then, we observe for all $x\in\mathbb{R}$,
$\nu\big{(}(x,\infty)\big{)}=1-F(x)=1-\inf_{\mu\in M}\mu\big{(}(-\infty,x]\big{)}=\sup_{\mu\in M}\big{(}1-\mu\big{(}(-\infty,x]\big{)}\big{)}=\sup_{\mu\in M}\mu\big{(}(x,\infty)\big{)}.$ (3.6)
Now we show $\nu\in\mathcal{P}_{2}(\mathbb{R})$. Split the second moment of $\nu$ into two parts,
$\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathop{}\!\mathrm{d}\nu(x)=\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathbf{1}_{\{x\leq 0\}}\mathop{}\!\mathrm{d}\nu(x)+\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathbf{1}_{\{x>0\}}\mathop{}\!\mathrm{d}\nu(x).$ (3.7)
On the one hand, using Tonelli’s theorem and (3.6), we see for some fixed $\mu\in M\subset\mathcal{P}_{2}(\mathbb{R})$,
$\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathbf{1}_{\{x\leq 0\}}\mathop{}\!\mathrm{d}\nu(x)$ $=2\int_{\mathbb{R}}\int_{\mathbb{R}}(-u)\mathbf{1}_{\{x\leq u\leq 0\}}\mathop{}\!\mathrm{d}u\cdot\mathbf{1}_{\{x\leq 0\}}\mathop{}\!\mathrm{d}\nu(x)$ (3.8)
$=2\int_{\mathbb{R}}\int_{\mathbb{R}}\mathbf{1}_{\{x\leq u\}}\mathop{}\!\mathrm{d}\nu(x)\cdot(-u)\mathbf{1}_{\{u\leq 0\}}\mathop{}\!\mathrm{d}u$
$=2\int_{\mathbb{R}}\nu\big{(}(-\infty,u]\big{)}\cdot(-u)\mathbf{1}_{\{u\leq 0\}}\mathop{}\!\mathrm{d}u$
$\leq 2\int_{\mathbb{R}}\mu\big{(}(-\infty,u]\big{)}\cdot(-u)\mathbf{1}_{\{u\leq 0\}}\mathop{}\!\mathrm{d}u$
$=\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathbf{1}_{\{x\leq 0\}}\mathop{}\!\mathrm{d}\mu(x)<\infty.$
On the other hand, use Tonelli’s theorem again,
$\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathbf{1}_{\{x>0\}}\mathop{}\!\mathrm{d}\nu(x)=2\int_{\mathbb{R}}\int_{\mathbb{R}}u\mathbf{1}_{\{0<u<x\}}\mathop{}\!\mathrm{d}u\cdot\mathbf{1}_{\{x>0\}}\mathop{}\!\mathrm{d}\nu(x)=2\int_{\mathbb{R}}\nu\big{(}(u,\infty)\big{)}\cdot u\mathbf{1}_{\{u>0\}}\mathop{}\!\mathrm{d}u.$ (3.9)
Combine (3.1), (3.6), (3.9), and we see
$\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathbf{1}_{\{x>0\}}\mathop{}\!\mathrm{d}\nu(x)=2\int_{0}^{\infty}\bigg{(}\sup_{\mu\in M}\mu\big{(}(x,\infty)\big{)}\bigg{)}\mathop{}\!\mathrm{d}u<\infty.$ (3.10)
It follows from (3.7), (3.8) and (3.10) that $\int_{\mathbb{R}}\left\lvert x\right\rvert^{2}\mathop{}\!\mathrm{d}\nu(x)<\infty$, so $\nu\in\mathcal{P}_{2}(\mathbb{R})$.
It remains to show that $\nu\geq_{\mathrm{st}}M$. Since $F_{\nu}=F=\inf_{\mu\in M}F_{\mu}$, we know that $F_{\nu}\leq F_{\mu}$ for all $\mu\in M$. Lemma 3.1 yields that $\nu\geq_{\mathrm{st}}\mu$ for all $\mu\in M$, which completes the proof. ∎
Remark 3.2.
Actaully, the condition (3.1) guarantees the existence of $\sup_{\leq_{\mathrm{st}}}M$ in $\mathcal{P}_{2}(\mathbb{R})$. Let $\nu\in\mathcal{P}_{2}(\mathbb{R})$ be the probability measure constructed in the proof of Proposition 3.2, and we can show $\nu=\sup_{\leq_{\mathrm{st}}}M$. Suppose $\tilde{\nu}\geq_{\mathrm{st}}M$, and we need to show $\tilde{\nu}\geq_{\mathrm{st}}\nu$. Using Lemma 3.1, we have $F_{\tilde{\nu}}\leq F_{\mu}$ for all $\mu\in M$. Hence, $F_{\tilde{\nu}}\leq\inf_{\mu\in M}F_{\mu}=F_{\nu}$, then we apply Lemma 3.1 again to obtain $\tilde{\nu}\geq_{\mathrm{st}}\nu$. This verifies $\nu=\sup_{\leq_{\mathrm{st}}}M$. Likewise, the condition (3.2) implies the existence of $\inf_{\leq_{\mathrm{st}}}M$ in $\mathcal{P}_{2}(\mathbb{R})$.
Corollary 3.3.
If $M$ is a bounded subset of $\mathcal{P}_{p}(\mathbb{R})$ for some $p>2$, then $M$ is order bounded in $\mathcal{P}_{2}(\mathbb{R})$.
Proof.
It is sufficient to check the conditions (3.1) and (3.2). We first notice
$\int_{0}^{\infty}\bigg{(}\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u=\int_{0}^{1}\bigg{(}\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u+\int_{1}^{\infty}\bigg{(}\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u.$ (3.11)
Then we estimate these two terms separately. Note that
$\int_{0}^{1}\bigg{(}\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u\leq\int_{0}^{1}u\mathop{}\!\mathrm{d}u=\frac{1}{2},$ (3.12)
and by Chebyshev’s inequality, for $u\geq 1$,
$\mu\big{(}(u,\infty)\big{)}\leq\frac{1}{u^{p}}\int_{\mathbb{R}}\left\lvert x\right\rvert^{p}\mathop{}\!\mathrm{d}\mu(x).$
It follows
$\int_{1}^{\infty}\bigg{(}\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u\leq\sup_{\mu\in M}\left\lVert\mu\right\rVert_{p}^{p}\cdot\int_{1}^{\infty}\frac{\mathop{}\!\mathrm{d}u}{u^{p-1}}<\infty.$ (3.13)
Combining (3.11)-(3.13) verifies (3.1). The condition (3.2) can be verified in a similar way. ∎
By the equivalent characterization of order bounded sets in $\mathcal{P}_{2}(\mathbb{R})$, we also have the following result.
Proposition 3.4.
If $M$ is an order bounded subset of $\mathcal{P}_{2}(\mathbb{R})$, then $M$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. In particular, any order interval $[\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$.
Proof.
According to Ambrosio-Gigli [2, Theorem 2.7], it suffices to show that $M$ is tight and $2$-uniformly integrable, i.e.,
$\lim_{N\to\infty}\sup_{\mu\in M}\int_{\{x:|x|\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)=0,$ (3.14)
Since $\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}$ is decreasing in $u$, and by Proposition 3.2 (i), we have
$\sup_{\mu\in M}\mu\big{(}(u,\infty)\big{)}\to 0\ \text{ as $u\to\infty$}.$ (3.15)
Similarly, we observe $\sup_{\mu\in M}\mu\big{(}(-\infty,-u]\big{)}$ is decreasing in $u$, so Proposition 3.2 (ii) entails
$\sup_{\mu\in M}\mu\big{(}(-\infty,-u]\big{)}\to 0\ \text{ as $u\to\infty$}.$ (3.16)
Then (3.15) and (3.16) together imply $M$ is tight.
It remains to show that $M$ is 2-uniformly integrable, i.e., (3.14) holds. Let $\underline{\mu},\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R})$ be a lower and an upper bounds of $M$ with respect to the stochastic order. Note that for any $\mu\in M$, by choosing $f^{+}_{N}(x)=|x|^{2}\mathbf{1}_{\{x:x\geq N\}}$ and $f^{-}_{N}(x)=-|x|^{2}\mathbf{1}_{\{x:x\leq-N\}}$ respectively, we deduce
$\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\overline{\mu}(x),\ \text{ and }\ \int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\underline{\mu}(x).$
Thus, for all $\mu\in M$,
$\int_{\{x:|x|\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\overline{\mu}(x)+\int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\underline{\mu}(x).$
Then we know that
$\lim_{N\to\infty}\sup_{\mu\in M}\int_{\{x:|x|\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\lim_{N\to\infty}\Bigg{(}\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\overline{\mu}(x)+\int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\underline{\mu}(x)\Bigg{)}=0.$
Hence, $M$ is $2$-uniformly integrable in $\mathcal{P}_{2}(\mathbb{R})$.
In particular, an order interval $[\mu,\nu]$ is closed due to the closedness of the stochastic order, so it is compact in $\mathcal{P}_{2}(\mathbb{R})$. ∎
Remark 3.3.
If $M$ is merely a relatively compact subset of $\mathcal{P}_{2}(\mathbb{R})$, then $M$ may not have upper or lower bounds in $\mathcal{P}_{2}(\mathbb{R})$ with respect to the stochastic order. In the following, we give an example of a relatively compact subset of $\mathcal{P}_{2}(\mathbb{R})$, but it fails to satisfy the equivalent conditions (3.1)(3.2) of order boundedness. Let $M=\{\mu_{n}:n\geq 3\}\subset\mathcal{P}_{2}(\mathbb{R})$ be given by
$\mu_{n}=\frac{1}{n^{2}\log n}\delta_{-n}+\left(1-\frac{2}{n^{2}\log n}\right)\delta_{0}+\frac{1}{n^{2}\log n}\delta_{n}.$
Since $\mathcal{W}_{2}(\mu_{n},\delta_{0})=(2/\log n)^{1/2}\to 0$ as $n\to\infty$, the set $\{\mu_{n}\}_{n=3}^{\infty}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. Note that
$\inf_{n\geq 3}F_{\mu_{n}}(x)=\left\{\begin{aligned} &0,&&x<0,\\ &1-\frac{1}{3^{2}\log 3},&&0\leq x<3,\\ &1-\frac{1}{k^{2}\log k},&&k-1\leq x<k,\ k\geq 4.\end{aligned}\right.$
It gives
$\int_{0}^{\infty}\bigg{(}\sup_{n\geq 3}\mu_{n}\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u=\int_{0}^{\infty}\big{[}1-\inf_{n\geq 3}F_{\mu_{n}}(u)\big{]}u\mathop{}\!\mathrm{d}u\geq\sum_{k=3}^{\infty}\frac{k}{(k+1)^{2}\log(k+1)}=\infty,$
so the condition (3.1) does not hold, and $M$ has no upper bound in $\mathcal{P}_{2}(\mathbb{R})$ with respect to the stochastic order. Likewise, we can check $M$ does not satisfy (3.2), and hence, it has no lower bound.
Corollary 3.5.
Let $[\mu,\nu]$ be an order interval and $\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, $\rho\in[\mu,\nu]$. Then the following convergences are equivalent.
  1. (i)
    $\mathcal{W}_{2}(\rho_{n},\rho)\to 0$ as $n\to\infty$;
  2. (ii)
    $\mathcal{W}_{p}(\rho_{n},\rho)\to 0$ as $n\to\infty$, for all $21\leq p<$ 2;
  3. (iii)
    $\rho_{n}\to\rho$ weakly as $n\to\infty$.
Proof.
It is obvious that (i)$\Rightarrow$(ii)$\Rightarrow$(iii), and we only need to show (iii)$\Rightarrow$(i).
By Proposition 3.4, we know that $[\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$. Then for any subsequence $\{\rho_{n_{k}}\}_{k\geq 1}\subset\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, there exists a subsequence $\{\rho_{n_{k_{j}}}\}_{j\geq 1}$ of $\{\rho_{n_{k}}\}_{k\geq 1}$ and a $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R})$ such that $\mathcal{W}_{2}(\rho_{n_{k_{j}}},\rho_{0})\to 0$ as $j\to\infty$. Thus, $\rho_{n_{k_{j}}}\to\rho_{0}$ weakly as $j\to\infty$. By (iii) and the uniqueness of the weak limit, we conclude $\rho_{0}=\rho$. Therefore, $\mathcal{W}_{2}(\rho_{n},\rho)\to 0$ as $n\to\infty$. ∎
4 Global Convergence to Order Interval
This section aims to prove Theorem 2.1 and Corollary 2.2. In Section 4.1, we target on the structure of global attractors of general semigroups on $\mathcal{P}_{2}(\mathbb{R})$ that preserve the stochastic order (namely, order-preserving semigroups). In Section 4.2, we prove a comparison theorem for extrinsic-law-dependent SDEs, and obtain the generation of order-preserving semigroup by the equation (1.1). The coming down from infinity property of hyper-dissipative McKean-Vlasov SDEs is shown in Section 4.3. At last, the proof of Theorem 2.1 and Corollary 2.2 is given in Section 4.4.
4.1 Global attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
In this subsection, we utilize the concepts of dissipative dynamical systems (see e.g., Hale [39]) to study the asymptotic behaviour of a semigroup on $\mathcal{P}_{2}(\mathbb{R})$ that preserves the stochastic order. With a Lyapunov condition, [14, 34] study the unique ergodicity of order-preserving Markov processes. Here, we focus on the structure of the global attractor of an order-preserving semigroup. Based on the existence of a globally attracting order bounded set, we prove that a global attractor of an order-preserving semigroup must be encompassed by two order related invariant measures, and these two invariant measures attract all solutions initiated from the regions above and below them with respect to the stochastic order, respectively (see Theorem 4.2).
Firstly, we give the definition of a semiflow on $\mathcal{P}_{2}(\mathbb{R})$.
Definition 4.1.
A map $P^{*}:\mathbb{R_{+}}\times\mathcal{P}_{2}(\mathbb{R})\rightarrow\mathcal{P}_{2}(\mathbb{R})$ is said to be a jointly continuous semigroup, or a semiflow on $\mathcal{P}_{2}(\mathbb{R})$, if
  1. (i)
    $P^{*}:\mathbb{R_{+}}\times\mathcal{P}_{2}(\mathbb{R})\rightarrow\mathcal{P}_{2}(\mathbb{R})$ is continuous;
  2. (ii)
    $P_{0}^{*}\mu=\mu$ for all $\mu\in\mathcal{P}_{2}(\mathbb{R})$;
  3. (iii)
    $P_{t}^{*}\circ P_{s}^{*}=P_{t+s}^{*}$ for all $t,s\in\mathbb{R_{+}}$.
Next, we introduce some necessary notations and definitions originated from the standard theory of dissipative dynamical systems. For any $\mu\in\mathcal{P}_{2}(\mathbb{R})$, the positive orbit $O_{+}(\mu)$ of $\mu$ is $\{P^{*}_{t}\mu:t\geq 0\}$. A measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$ is an invariant measure of $P^{*}$, if $P^{*}_{t}\mu=\mu$ for any $t\geq 0$. A subset $A\subset\mathcal{P}_{2}(\mathbb{R})$ is said to be positively invariant, if $P^{*}_{t}A\subset A$, for any $t\geq 0$, and invariant, if $P^{*}_{t}A=A$, for any $t\geq 0$. $A\subset\mathcal{P}_{2}(\mathbb{R})$ is said to be eventually positively invariant, if there exists $t_{0}\geq 0$ such that, $P^{*}_{t}A\subset A$, for any $t\geq t_{0}$. A compact invariant set $A\subset\mathcal{P}_{2}(\mathbb{R})$ is said to be the maximal compact invariant set if every compact invariant set is contained in $A$.
Definition 4.2.
A set $K\subset\mathcal{P}_{2}(\mathbb{R})$ is said to attract a set $B\subset\mathcal{P}_{2}(\mathbb{R})$ if $\mathcal{W}_{2}(P^{*}_{t}B,K)\to 0$, as $t\to\infty$.
Obviously, if a measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$ is attracted by a compact set $K\subset\mathcal{P}_{2}(\mathbb{R})$, then the positive orbit $O_{+}(\mu)$ of $\mu$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$.
Definition 4.3.
A compact invariant set $A\subset\mathcal{P}_{2}(\mathbb{R})$ is said to be the global attractor, if $A$ attracts each bounded subset of $\mathcal{P}_{2}(\mathbb{R})$.
Clearly, if the global attractor exists, it must be unique, maximal and all the invariant measures are contained in it.
Definition 4.4.
For any set $B\subset\mathcal{P}_{2}(\mathbb{R})$, the $\omega$-limit set of $B$ is defined by
$\omega(B):=\mathop{\bigcap}\limits_{s\geq 0}\overline{\bigcup\limits_{t\geq s}P^{*}_{t}B},$
where $\overline{D}$ means the closure of $D$ in $\mathcal{P}_{2}(\mathbb{R})$.
Now, we give some basic properties of $\omega$-limit sets and dissipative dynamics.
Lemma 4.1.
Given a semiflow $P^{*}$ on $\mathcal{P}_{2}(\mathbb{R})$, then
  1. (i)
    a measure $\nu\in\omega(B)$ if and only if there exist two sequences $t_{k}\to\infty$, and $\nu_{k}\in B$ such that $\mathcal{W}_{2}(P^{*}_{t_{k}}\nu_{k},\nu)\to 0$, as $k\to\infty$;
  2. (ii)
    if $B$ is connected in $\mathcal{P}_{2}(\mathbb{R})$, then $\omega(B)$ is connected in $\mathcal{P}_{2}(\mathbb{R})$;
  3. (iii)
    if there exists a compact set $K\subset\mathcal{P}_{2}(\mathbb{R})$ attracting a bounded set $B\subset\mathcal{P}_{2}(\mathbb{R})$, then $\omega(B)$ is a nonempty, compact, invariant subset of $K$ and $\omega(B)$ attracts $B$;
  4. (iv)
    if there exists a nonempty compact set $K\subset\mathcal{P}_{2}(\mathbb{R})$ that attracts each compact subset of $\mathcal{P}_{2}(\mathbb{R})$, then $\mathop{\bigcap}\limits_{t\geq 0}P^{*}_{t}K$ is the maximal compact invariant set.
Proof.
(i) is a direct consequence of the definition of $\omega$-limit sets. (ii)-(iv) are standard results for dissipative systems on complete metric spaces, for their proof one can see Hale [39, Lemma 3.1.1, Lemma 3.2.1 and Theorem 3.4.2 (ii)] respectively. ∎
Clearly, Lemma 4.1 (iii) implies that, if the global attractor $A$ exists, then $\omega(B)\subset A$ for any bounded set $B\subset\mathcal{P}_{2}(\mathbb{R})$. In the following, we give the definition of order-preserving semigroups.
Definition 4.5.
A semiflow $P^{*}$ on $\mathcal{P}_{2}(\mathbb{R})$ is called order-preserving (or, monotone), if $\mu\leq_{\mathrm{st}}\nu$ implies $P^{*}_{t}\mu\leq_{\mathrm{st}}P^{*}_{t}\nu$ for any $t>0$.
Now, we give our main theorem in this subsection concerning the structure of the global attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$.
Theorem 4.2.
(Global attractor). Let $P^{*}$ be an order-preserving semiflow on $\mathcal{P}_{2}(\mathbb{R})$. Assume that there exists an eventually positively invariant, order bounded, closed set $K\subset\mathcal{P}_{2}(\mathbb{R})$ such that $K$ attracts each bounded subset of $\mathcal{P}_{2}(\mathbb{R})$. Then,
  1. (i)
    $\omega(K)=\mathop{\bigcap}\limits_{t\geq 0}P^{*}_{t}K$ is the global attractor. Furthermore, if $K$ is connected in $\mathcal{P}_{2}(\mathbb{R})$, $\omega(K)$ is connected in $\mathcal{P}_{2}(\mathbb{R})$;
  1. (ii)
    $\omega(K)\subset[\underline{\nu},\overline{\nu}]$, where $\underline{\nu},\overline{\nu}\in\omega(K)$ are invariant measures;
  2. (iii)
    $\underline{\nu}$ is globally attracting from below and $\overline{\nu}$ is globally attracting from above i.e., for any $\rho\leq_{\mathrm{st}}\underline{\nu}$, we have $\mathcal{W}_{2}(P^{*}_{t}\rho,\underline{\nu})\to 0$, as $t\to\infty$; for any $\rho\geq_{\mathrm{st}}\overline{\nu}$, we have $\mathcal{W}_{2}(P^{*}_{t}\rho,\overline{\nu})\to 0$, as $t\to\infty$;
  3. (iv)
    If $\underline{\nu}$ attracts a set $U\subset\mathcal{P}_{2}(\mathbb{R})$, then $\underline{\nu}$ attracts each measure in $\bigcup\limits_{\mu\in U}\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\rho\leq_{\mathrm{st}}\mu\}$. Correspondingly, if $\overline{\nu}$ attracts a set $V\subset\mathcal{P}_{2}(\mathbb{R})$, then $\overline{\nu}$ attracts each measure in $\bigcup\limits_{\mu\in V}\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\rho\geq_{\mathrm{st}}\mu\}$;
  4. (v)
    If there exists $\rho_{1}<_{\mathrm{st}}\overline{\nu}$ (resp. $\rho_{2}>_{\mathrm{st}}\underline{\nu}$) such that $\mathcal{W}_{2}(P^{*}_{t}\rho_{1},\overline{\nu})\to 0$ (resp. $\mathcal{W}_{2}(P^{*}_{t}\rho_{2},\underline{\nu})\to 0$) as $t\to\infty$, then for any $\rho\geq_{\mathrm{st}}\rho_{1}$ (resp. $\rho\leq_{\mathrm{st}}\rho_{2}$), one has $\mathcal{W}_{2}(P^{*}_{t}\rho,\overline{\nu})\to 0$ (resp. $\mathcal{W}_{2}(P^{*}_{t}\rho,\underline{\nu})\to 0$), as $t\to\infty$.
  5. (vi)
    If $P^{*}$ has a unique invariant measure $\nu$, then the singleton set $\{\nu\}$ is the global attractor. That is to say, $\nu$ attracts all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
Proof.
(i). By Proposition 3.4, the order bounded, closed set $K$ is compact in $\mathcal{P}_{2}(\mathbb{R})$. Since $K$ is eventually positively invariant, by definition of $\omega(K)$ we have $\omega(K)\subset K$. Since $K$ attracts $K$, Lemma 4.1 (iii) implies that $\omega(K)$ is invariant. Then $\omega(K)=P^{*}_{t}\omega(K)\subset P^{*}_{t}K$ for any $t\geq 0$. Hence $\omega(K)\subset\mathop{\bigcap}\limits_{t\geq 0}P^{*}_{t}K.$ On the other hand, the definition of $\omega(K)$ implies that $\mathop{\bigcap}\limits_{t\geq 0}P^{*}_{t}K\subset\omega(K).$ Hence, $\omega(K)=\mathop{\bigcap}\limits_{t\geq 0}P^{*}_{t}K.$ By virtue of the fact that $K$ attracts every bounded subset of $\mathcal{P}_{2}(\mathbb{R})$, it follows from Lemma 4.1 (iv) that $\omega(K)$ is the maximal compact invariant set. To prove $\omega(K)$ is the global attractor, we only need to show that $\omega(K)$ attracts every bounded subset $B$ of $\mathcal{P}_{2}(\mathbb{R})$. In fact, by Lemma 4.1 (iii), $\omega(B)$ is an invariant subset of $K$ and $\omega(B)$ attracts $B$. Since $\omega(B)$ is an invariant subset of $K$, one has $\omega(B)\subset\omega(K)$. Therefore, $\omega(K)$ also attracts $B$. The connectedness of $\omega(K)$ follows directly from the connectedness of $K$ and Lemma 4.1 (ii). This proves (i).
(ii). Since $K$ is order-bounded, assume that there exists $\mu_{1}$, $\mu_{2}\in\mathcal{P}_{2}(\mathbb{R})$ such that $\mu_{1}\leq_{\mathrm{st}}K$ and $\mu_{2}\geq_{\mathrm{st}}K$. We only prove that $\sup_{\leq_{\mathrm{st}}}\omega(K)$ exists, $\sup_{\leq_{\mathrm{st}}}\omega(K)\in\omega(K)$ and $\sup_{\leq_{\mathrm{st}}}\omega(K)$ is an invariant measure. The proof for $\inf_{\leq_{\mathrm{st}}}\omega(K)$ is similar, we omit it here. Since $\mu_{2}\geq_{\mathrm{st}}K$ and $\omega(K)\subset K$, one has $\mu_{2}\geq_{\mathrm{st}}\omega(K)$. As is shown in (i), $\omega(K)$ is the global attractor. Thus, $\mathcal{W}_{2}(P^{*}_{t}\mu_{2},\omega(K))\to 0$ as $t\to\infty$. Since $\omega(K)$ is compact, one can choose $t_{k}\to\infty$ and $\overline{\nu}\in\omega(K)$, such that $\mathcal{W}_{2}(P^{*}_{t_{k}}\mu_{2},\overline{\nu})\to 0$ as $k\to\infty$. For any $k\geq 1$, the order-preservation implies $P^{*}_{t_{k}}\mu_{2}\geq_{\mathrm{st}}P^{*}_{t_{k}}\omega(K)=\omega(K)$. Therefore, the closedness of the stochastic order entails that $\overline{\nu}\geq_{\mathrm{st}}\omega(K)$. Together with the fact $\overline{\nu}\in\omega(K)$, we have $\overline{\nu}=\sup_{\leq_{\mathrm{st}}}\omega(K)$. Fix any $t\geq 0$. On the one hand, $P^{*}_{t}\overline{\nu}\geq_{\mathrm{st}}P^{*}_{t}\omega(K)=\omega(K)$ and $\overline{\nu}\in\omega(K)$ entail that $P^{*}_{t}\overline{\nu}\geq_{\mathrm{st}}\overline{\nu}$. On the other hand, $P^{*}_{t}\overline{\nu}\in P^{*}_{t}\omega(K)=\omega(K)$ and $\overline{\nu}=\sup_{\leq_{\mathrm{st}}}\omega(K)$ imply that $P^{*}_{t}\overline{\nu}\leq_{\mathrm{st}}\overline{\nu}$. Thus, by the antisymmetry of $\leq_{\mathrm{st}}$, $P^{*}_{t}\overline{\nu}=\overline{\nu}$, for any $t\geq 0$. That is to say, $\overline{\nu}$ is an invariant measure. Hence, we have proved (ii).
(iii). We only prove that $\overline{\nu}$ is globally attracting from above i.e., for any $\rho\geq_{\mathrm{st}}\overline{\nu}$, we have $\mathcal{W}_{2}(P^{*}_{t}\rho,\overline{\nu})\to 0$, as $t\to\infty$. The proof for $\underline{\nu}$ is similar, we omit it here. Note that $\{P_{t}^{*}\rho\}_{t\geq 0}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. For any $\pi\in\mathcal{P}_{2}(\mathbb{R})$ such that, $\mathcal{W}_{2}(P^{*}_{t_{k}}\rho,\pi)\to 0$ as $k\to\infty$ for some sequence $t_{k}\to\infty$, we only need to prove $\pi=\overline{\nu}$. In fact, since the global attractor $\omega(K)$ attracts $\rho$, one has $\pi\in\omega(K)$. For any $k\geq 1$, we have $P^{*}_{t_{k}}\rho\geq_{\mathrm{st}}P^{*}_{t_{k}}\overline{\nu}=\overline{\nu}$. Therefore, the closedness of the stochastic order entails that $\pi\geq_{\mathrm{st}}\overline{\nu}$. Together with the facts $\pi\in\omega(K)$ and $\overline{\nu}=\sup_{\leq_{\mathrm{st}}}\omega(K)$, we have $\pi=\overline{\nu}$. This proves (iii).
(iv). We only give the proof of the case of $\overline{\nu}$ (the case of $\underline{\nu}$ is similar). In fact, if $\rho\geq_{\mathrm{st}}\mu$ for some $\mu\in V$, by Lemma 4.1 (iii), it suffice to prove that $\omega(\rho)=\{\overline{\nu}\}$. Let $\pi\in\omega(\rho)$, Lemma 4.1 (i) entails that there exists a sequence $t_{k}\to\infty$ such that $\mathcal{W}_{2}(P^{*}_{t_{k}}\rho,\pi)\to 0$, as $k\to\infty$. Since $P^{*}$ is order-preserving and $\rho\geq_{\mathrm{st}}\mu$, we have $P^{*}_{t_{k}}\rho\geq_{\mathrm{st}}P^{*}_{t_{k}}\mu$ for any $k\geq 1$. Together with the closedness of the stochastic order, $\mathcal{W}_{2}(P^{*}_{t_{k}}\mu,\overline{\nu})\to 0$ as $k\to\infty$ implies that $\pi\geq_{\mathrm{st}}\overline{\nu}$. Since the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$, one has $\pi\in\omega(\rho)\subset\omega(K)\subset[\underline{\nu},\overline{\nu}]$. Therefore, $\pi=\overline{\nu}$. This proves (iv).
(v). We only give the proof of the case of $\overline{\nu}$ (the case of $\underline{\nu}$ is similar). In fact, for any $\rho\geq_{\mathrm{st}}\rho_{1}$, define $\pi:=\sup_{\leq_{\mathrm{st}}}\{\rho,\overline{\nu}\}\in\mathcal{P}_{2}(\mathbb{R})$ (the existence of $\pi$ follows from Remark 3.1). (iii) entails that $\mathcal{W}_{2}(P^{*}_{t}\pi,\overline{\nu})\to 0$, as $t\to\infty$. On the other hand, we also have $\mathcal{W}_{2}(P^{*}_{t}\rho_{1},\overline{\nu})\to 0$ as $t\to\infty$. Noticing that $\rho\in[\rho_{1},\pi]$, the order-preserving property of $P^{*}$ entails that $P^{*}_{t}\rho_{1}\leq_{\mathrm{st}}P^{*}_{t}\rho\leq_{\mathrm{st}}P^{*}_{t}\pi$ for any $t\geq 0$. Thus, the relative compactness of $\{P_{t}^{*}\rho\}_{t\geq 0}$, the closedness and antisymmetry of the stochastic order entail that $\mathcal{W}_{2}(P^{*}_{t}\rho,\overline{\nu})\to 0$, as $t\to\infty$. Now, we obtain (v).
(vi). By virtue of (ii), one has $\nu\leq_{\mathrm{st}}\omega(K)\leq_{\mathrm{st}}\nu$. Thus, the antisymmetry of the stochastic order entails that $\omega(K)=\{\nu\}$. Hence, $\{\nu\}$ is the global attractor. This proves (vi). ∎
Remark 4.1.
(i). If we consider a semiflow preserving the stochastic order acting only on a closed subspace $B$ of $\mathcal{P}_{2}(\mathbb{R})$ (rather than the entire space $\mathcal{P}_{2}(\mathbb{R})$), by repeating the same proof, Theorem 4.2 still holds, where the order interval within $B$ is the restriction of the order interval to $B$. In fact, for semiflow preserving some closed partial order relation on a general complete metric space, by assuming the atttracting set $K$ is compact and order bounded, Theorem 4.2 still holds.
(ii). For semiflows preserving some closed partial order relation on a strongly ordered space, Hirsch [41, Theorem 3.3] implies the unique equilibrium must be globally attracting (Theorem 4.2 (vi)). However, Proposition A.1 shows that the strong ordering relation generated by $\leq_{\mathrm{st}}$ is empty. So, [41, Theorem 3.3] is not applicable here.
4.2 Comparison principle for extrinsic-law-dependent SDEs
Now let us go back to the McKean-Vlasov equation (1.1). The driving noise $\{B_{t}\}_{t\in\mathbb{R}}$ is a two-sided one-dimensional standard Brownian motion defined on a probability space $(\Omega,\mathcal{F},\mathbf{P})$, and let $\{\mathcal{F}_{t}\}_{t\in\mathbb{R}}$ be the natural filtration generated by $B$, that is, $\mathcal{F}_{t}=\sigma\{B_{u}-B_{v}:u,v\leq t\}$. Under Assumption 1, for any $s\in\mathbb{R}$ and $\xi\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$, it is well-known that (1.1) has a unique solution $\{X_{t}^{s,\xi}\}_{t\geq s}$ starting at time $s$ with initial data $\xi$, i.e., $\{X_{t}^{s,\xi}\}_{t\geq s}$ solves the following equation (see e.g., Wang [69])
$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mathcal{L}(X_{t}))(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},\ \ X_{s}=\xi,\ \ t\geq s.$
When the starting time $s=0$, we denote the solution $\{X_{t}^{0,\xi}\}_{t\geq 0}$ by $\{X_{t}^{\xi}\}_{t\geq 0}$.
Next, we present a comparison theorem for extrinsic-law-dependent SDEs with respect to the stochastic order. Let $C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ be the space of continuous functions from $[s,\infty)$ to $\mathcal{P}_{2}(\mathbb{R})$. For any $\mu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $\xi\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$, we denote $\{X_{t}^{s,\mu,\xi}\}_{t\geq s}$ by the solution of the following extrinsic-law-dependent SDE
$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mu_{t})(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},\ \ X_{s}=\xi,\ \ t\geq s.$ (4.1)
Similarly, we denote the solution $\{X_{t}^{0,\mu,\xi}\}_{t\geq 0}$ by $\{X_{t}^{\mu,\xi}\}_{t\geq 0}$.
Then we have the following comparison principle.
Theorem 4.3.
Suppose Assumption 1 holds. Let $\mu,\nu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $T>0$. Assume
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{t}(x)\leq\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{t}(x),\ \text{ for all }\ t\in[s,s+T].$ (4.2)
If $\xi,\eta\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ and $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, then $\mathcal{L}(X_{t}^{s,\mu,\xi})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{s,\nu,\eta})$ for all $t\in[s,s+T]$.
Proof.
Since $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, Strassen’s theorem (see Lindvall [47, equation (3)]) yields that there exist $\tilde{\xi},\tilde{\eta}\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ such that
[…]
It suffices to show that $\mathcal{L}(X_{t}^{s,\mu,\tilde{\xi}})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{s,\nu,\tilde{\eta}})$ for all $t\in[s,s+T]$.
Note that
$\mathop{}\!\mathrm{d}X^{s,\mu,\tilde{\xi}}_{t}$ $=\Big{[}-V^{\prime}(X^{s,\mu,\tilde{\xi}}_{t})-\theta X^{s,\mu,\tilde{\xi}}_{t}+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{t}(x)\Big{]}\mathop{}\!\mathrm{d}t+\sigma(X^{s,\mu,\tilde{\xi}}_{t})\mathop{}\!\mathrm{d}B_{t};$ (4.4)
$\mathop{}\!\mathrm{d}Y^{s,\nu,\tilde{\eta}}_{t}$ $=\Big{[}-V^{\prime}(Y^{s,\nu,\tilde{\eta}}_{t})-\theta Y^{s,\nu,\tilde{\eta}}_{t}+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{t}(x)\Big{]}\mathop{}\!\mathrm{d}t+\sigma(Y^{s,\nu,\tilde{\eta}}_{t})\mathop{}\!\mathrm{d}B_{t}.$
Since $\theta>0$, according to the comparison theorem for usual SDEs [37, Theorem 1.2], we know that
[…]
The desired result is proved. ∎
Remark 4.2.
Theorem 4.3 is not only used in the proof of the order-preservation for semigroups of McKean-Vlasov SDEs, but also crucial in the proof of our main result Theorem 2.3.
Associated to the McKean-Vlasov SDE (1.1), there is a semigroup $P_{t}^{*}$ acting on $\mathcal{P}_{2}(\mathbb{R})$ by
$P_{t}^{*}\mu:=\mathcal{L}(X_{t}^{\xi})\ \text{ with }\ \mathcal{L}(\xi)=\mu.$ (4.5)
In the following, we show that $P_{t}^{*}$ is an order-preserving semiflow.
Theorem 4.4.
Under Assumption 1, the semigroup $P^{*}:[0,\infty)\times\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ is an order-preserving semiflow.
Proof.
We first prove $P^{*}$ is order-preserving. It suffices to show that $P^{*}_{t}\rho_{1}\leq_{\mathrm{st}}P^{*}_{t}\rho_{2}$ for any $\rho_{1}<_{\mathrm{st}}\rho_{2}$ and all $t\geq 0$. By Strassen’s theorem, there exist $\xi_{1},\xi_{2}\in L^{2}(\Omega,\mathcal{F}_{0},\mathbf{P})$ such that
$\mathcal{L}(\xi_{1})=\rho_{1},\ \ \mathcal{L}(\xi_{2})=\rho_{2},\ \ \text{ and}\ \ \xi_{1}\leq\xi_{2},\ \mathbf{P}-a.s.$
Since $\rho_{1}\neq\rho_{2}$, we have $\mathbf{P}\{\xi_{1}<\xi_{2}\}>0$ and hence
$\mathbf{E}[\xi_{1}]<\mathbf{E}[\xi_{2}].$ (4.6)
Let
$T:=\sup\{t>0:\mathbf{E}[X_{s}^{\xi_{1}}]<\mathbf{E}[X_{s}^{\xi_{2}}],\ \text{ for all }s\in[0,t]\}.$
Note that $\mathbf{E}[X_{t}^{\xi_{i}}]$ is continuous in $t$ for $i=1,2$, then it follows that $T>0$. The proof is divided into the following two cases:
$\textbf{Case 1}.\ \ T=\infty,\ \ \textbf{Case 2}.\ \ T<\infty.$
In case 1, let $\mu^{i}_{t}:=\mathcal{L}(X_{t}^{\xi_{i}}),t\in[0,\infty)$, $i=1,2$. Then we know that
$X_{t}^{\mu^{i},\xi_{i}}=X_{t}^{\xi_{i}},\ \text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{t}(x)\leq\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{t}(x),\ \text{ for all }\ t\geq 0.$
Then we conclude from Theorem 4.3 that
$P^{*}_{t}\rho_{1}=\mathcal{L}(X_{t}^{\mu^{1},\xi_{1}})\leq_{\mathrm{st}}\mathcal{L}(X_{t}^{\mu^{2},\xi_{2}})=P^{*}_{t}\rho_{2},\ \text{ for all }t\geq 0.$
In case 2, we have $0<T<\infty$. Let $\mu^{i}$ be defined as in case 1, we have $X_{t}^{\mu^{i},\xi_{i}}=X_{t}^{\xi_{i}}$ for all $t\geq 0$ and
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{t}(x)<\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{t}(x),\ t\in[0,T)\ \text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{T}(x)=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{T}(x).$
Using Theorem 4.3 again, we deduce
$P^{*}_{t}\rho_{1}\leq_{\mathrm{st}}P^{*}_{t}\rho_{2},\ \text{ for all }t\in[0,T].$
Note that $\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{T}(x)=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{T}(x)$, by the proof of (4.6), we conclude that $P^{*}_{T}\rho_{1}=P^{*}_{T}\rho_{2}$. Then we have $P^{*}_{t}\rho_{1}=P^{*}_{t}\rho_{2}$ for all $t\geq T$.
So the order-preservation of $P^{*}$ is proved. It remains to show that $P^{*}$ is jointly continuous, which follows from [48, Proposition 2.3] directly. ∎
4.3 Coming down from infinity
In this subsection, we prove hyper-dissipative McKean-Vlasov SDEs have the coming down from infinity property. It means, there exists a compact subset of $\mathcal{P}_{2}(\mathbb{R})$ attracting all bounded sets. As a byproduct, we show $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is a compact map.
Suppose Assumption 1 (ii)-(iv) hold. Then for any $x\in\mathbb{R}$, $\mu\in\mathcal{P}_{2}(\mathbb{R})$, we have
$\begin{split}-xV^{\prime}(x)-x(W^{\prime}*\mu)(x)&=-xV^{\prime}(x)-\theta x^{2}+\theta x\int_{\mathbb{R}}y\mathop{}\!\mathrm{d}\mu(y)\\ &\leq-\alpha|x|^{2+\delta}-\theta|x|^{2}+\theta|x|\|\mu\|_{2}+L|x|+\beta.\end{split}$ (4.7)
The following lemma provides $p$-th moment estimation of the solution $\{X_{t}^{\xi}\}_{t\geq 0}$.
Lemma 4.5.
Suppose Assumption 1 holds. For any $p\geq 2$ and $\xi\in L^{p}(\Omega,\mathcal{F}_{0},\mathbf{P})$, there exists $C(p)=C(p,\alpha,\beta,\delta,\overline{\sigma},L)>0$ such that the unique solution $\{X_{t}^{\xi}\}_{t\geq 0}$ of (1.1) satisfies
$\mathbf{E}[|X_{t}^{\xi}|^{p}]\leq e^{-pt}\mathbf{E}[|\xi|^{p}]+C(p),\ \text{ for all }\ t\geq 0.$ (4.8)
Proof.
Applying Itô’s formula to $|X_{t}^{\xi}|^{2}$ together with (4.7), we have
$\begin{split}\mathop{}\!\mathrm{d}|X_{t}^{\xi}|^{2}&=\big{(}-2X_{t}^{\xi}V^{\prime}(X_{t}^{\xi})-2\theta|X_{t}^{\xi}|^{2}+2\theta X_{t}^{\xi}\mathbf{E}[X_{t}^{\xi}]+\sigma^{2}(X_{t}^{\xi})\big{)}\mathop{}\!\mathrm{d}t+2X_{t}^{\xi}\sigma(X_{t}^{\xi})\mathop{}\!\mathrm{d}B_{t}\\ &\leq\big{(}-2\alpha|X_{t}^{\xi}|^{2+\delta}-2\theta|X_{t}^{\xi}|^{2}+2\theta|X_{t}^{\xi}|(\mathbf{E}[|X_{t}^{\xi}|^{2}])^{\frac{1}{2}}+2L|X_{t}^{\xi}|+2\beta+\overline{\sigma}^{2}\big{)}\mathop{}\!\mathrm{d}t\\ &\ \ \ \ \qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+2X_{t}^{\xi}\sigma(X_{t}^{\xi})\mathop{}\!\mathrm{d}B_{t}.\end{split}$ (4.9)
Then Young’s inequality gives
$\mathop{}\!\mathrm{d}|X_{t}^{\xi}|^{2}\leq\big{(}-(2+\theta)|X_{t}^{\xi}|^{2}+\theta\mathbf{E}[|X_{t}^{\xi}|^{2}]+3^{2\delta^{-1}+1}(2\alpha)^{-2\delta^{-1}}+L^{2}+2\beta+\overline{\sigma}^{2}\big{)}\mathop{}\!\mathrm{d}t+2X_{t}^{\xi}\sigma(X_{t}^{\xi})\mathop{}\!\mathrm{d}B_{t}.$
Thus, applying Itô’s formula to $e^{2t}|X_{t}^{\xi}|^{2}$ and taking expectations yield
$e^{2t}\mathbf{E}[|X_{t}^{\xi}|^{2}]\leq\mathbf{E}[|\xi|^{2}]+2^{-1}\big{(}3^{2\delta^{-1}+1}(2\alpha)^{-2\delta^{-1}}+L^{2}+2\beta+\overline{\sigma}^{2}\big{)}e^{2t},$
which implies (4.8) for $p=2$.
For $p>2$, applying Itô’s formula to $|X_{t}^{\xi}|^{p}$ together with (4.9) and Young’s inequality gives
$\begin{split}\mathop{}\!\mathrm{d}|X_{t}^{\xi}|^{p}&=\frac{p}{2}|X_{t}^{\xi}|^{p-2}\mathop{}\!\mathrm{d}|X_{t}^{\xi}|^{2}+\frac{1}{2}\frac{p}{2}\big{(}\frac{p}{2}-1\big{)}|X_{t}^{\xi}|^{p-4}\mathop{}\!\mathrm{d}\langle|X^{\xi}|^{2}\rangle_{t}\\ &=\big{(}-p|X_{t}^{\xi}|^{p-2}X_{t}^{\xi}V^{\prime}(X_{t}^{\xi})-p\theta|X_{t}^{\xi}|^{p}+p\theta|X_{t}^{\xi}|^{p-2}X_{t}^{\xi}\mathbf{E}[X_{t}^{\xi}]+\frac{p(p-1)}{2}|X_{t}^{\xi}|^{p-2}\sigma^{2}(X_{t}^{\xi})\big{)}\mathop{}\!\mathrm{d}t\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+p|X_{t}^{\xi}|^{p-2}X_{t}^{\xi}\sigma(X_{t}^{\xi})\mathop{}\!\mathrm{d}B_{t}\\ &\leq\big{(}-p\alpha|X_{t}^{\xi}|^{p+\delta}-\theta|X_{t}^{\xi}|^{p}+\theta\mathbf{E}[|X_{t}^{\xi}|^{p}]+C(p)(|X_{t}^{\xi}|^{p-2}+|X_{t}^{\xi}|^{p-1})\big{)}\mathop{}\!\mathrm{d}t\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+p|X_{t}^{\xi}|^{p-2}X_{t}^{\xi}\sigma(X_{t}^{\xi})\mathop{}\!\mathrm{d}B_{t}\\ &\leq\big{(}-(p+\theta)|X_{t}^{\xi}|^{p}+\theta\mathbf{E}[|X_{t}^{\xi}|^{p}]+C(p)\big{)}\mathop{}\!\mathrm{d}t+p|X_{t}^{\xi}|^{p-2}X_{t}^{\xi}\sigma(X_{t}^{\xi})\mathop{}\!\mathrm{d}B_{t},\end{split}$
where $C(p)=C(p,\alpha,\beta,\delta,\overline{\sigma},L)$ is a constant, depending only on $(p,\alpha,\beta,\delta,\overline{\sigma},L)$, may be different line by line. Similarly, applying Itô’s formula to $e^{pt}|X_{t}^{\xi}|^{p}$ and taking expectations, we derive (4.8). ∎
Indeed, we want to show the solutions are $p$-integrable for all $p\geq 2$ if the initial conditions are just square integrable. For this purpose, we need the following extrinsic-law case.
Lemma 4.6.
Suppose Assumption 1 holds. If $\mu\in C([0,\infty);\mathcal{P}_{2}(\mathbb{R}))$, and if $\xi\in L^{2}(\Omega,\mathcal{F}_{0},\mathbf{P})$, then for any $p\geq 2$, there exists a constant $C(p)=C(p,\alpha,\beta,\delta,\overline{\sigma},\kappa,\theta,L)>0$ such that the unique solution $\{X_{t}^{\mu,\xi}\}_{t\geq 0}$ of (4.1) satisfies
$\big{(}\mathbf{E}\big{[}|X_{t}^{\mu,\xi}|^{p}\big{]}\big{)}^{1/p}\leq C(p)\Big{(}t^{-\frac{1}{\delta}}+t^{\frac{\kappa}{2}}+\sup_{s\in[0,t]}\|\mu_{s}\|_{2}^{\frac{2}{2+\delta}}\Big{)}\ \text{ for all $t>0$}.$
Proof.
For $t\geq 0$, set
$Z_{t}:=\int_{0}^{t}\sigma(X_{s}^{\mu,\xi})\mathop{}\!\mathrm{d}B_{s},\ \ \ Y_{t}:=X_{t}^{\mu,\xi}-Z_{t}.$
Then (4.1) implies that $Y_{t}$ solves the following random differential equation,
$Y_{t}=\xi+\int_{0}^{t}\Big{(}-V^{\prime}(Y_{s}+Z_{s})-\theta(Y_{s}+Z_{s})+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{s}(x)\Big{)}\mathop{}\!\mathrm{d}s.$
Hence, $|Y_{t}|^{2}$ is differentiable in $t$. Then by Assumption 1 (ii)-(iv), we have
$\begin{split}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}(|Y_{t}|^{2})&=2Y_{t}\Big{(}-V^{\prime}(Y_{t}+Z_{t})-\theta(Y_{t}+Z_{t})+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{t}(x)\Big{)}\\ &\leq-2\alpha|Y_{t}|^{2+\delta}+2\beta-2\theta|Y_{t}|^{2}+2\theta|Y_{t}|\big{(}|Z_{t}|+\theta^{-1}|V^{\prime}(Z_{t})|\big{)}+2\theta|Y_{t}|\|\mu_{t}\|_{2}\\ &\leq-2\alpha|Y_{t}|^{2+\delta}+\theta\|\mu_{t}\|_{2}^{2}+2\beta+\gamma+\gamma|Z_{t}|^{2\kappa},\end{split}$
where $\gamma=3\theta(1+\theta^{-2}L^{2})$. For $t\geq 0$, let
$y_{t}:=|Y_{t}|^{-\delta},\ \ \ \mu_{t}^{*}:=\sup_{s\in[0,t]}\|\mu_{s}\|_{2},\ \ \ Z_{t}^{*}:=\sup_{s\in[0,t]}|Z_{s}|.$
Then for any $t\geq 0$ with $y_{t}<\infty$, i.e., $Y_{t}\neq 0$, we have
$\begin{split}\frac{\mathop{}\!\mathrm{d}y_{t}}{\mathop{}\!\mathrm{d}t}&=-\frac{\delta}{2}|Y_{t}|^{-(2+\delta)}\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}(|Y_{t}|^{2})\geq\alpha\delta-\frac{\alpha\delta}{2}\Big{(}\frac{\theta}{\alpha}(\mu_{t}^{*})^{2}+\frac{2\beta+\gamma}{\alpha}+\frac{\gamma}{\alpha}(Z_{t}^{*})^{2\kappa}\Big{)}y_{t}^{\frac{2+\delta}{\delta}}.\end{split}$
Hence, for any $s\in[0,t]$,
$\frac{\mathop{}\!\mathrm{d}y_{s}}{\mathop{}\!\mathrm{d}s}\geq\frac{\alpha\delta}{2}>0,\ \text{ if }\ y_{s}\leq\Big{(}\frac{\theta}{\alpha}(\mu_{t}^{*})^{2}+\frac{2\beta+\gamma}{\alpha}+\frac{\gamma}{\alpha}(Z_{t}^{*})^{2\kappa}\Big{)}^{-\frac{\delta}{2+\delta}}.$ (4.10)
Then we have the following claim.
Claim: We conclude that
$y_{t}\geq\min\bigg{\{}\frac{\alpha\delta}{2}t,\ \Big{(}\frac{\theta}{\alpha}(\mu_{t}^{*})^{2}+\frac{2\beta+\gamma}{\alpha}+\frac{\gamma}{\alpha}(Z_{t}^{*})^{2\kappa}\Big{)}^{-\frac{\delta}{2+\delta}}\bigg{\}},\ \text{ for any }\ t\geq 0.$ (4.11)
Proof of Claim: For any fixed $t\geq 0$, set
$A_{t}:=\Big{(}\frac{\theta}{\alpha}(\mu_{t}^{*})^{2}+\frac{2\beta+\gamma}{\alpha}+\frac{\gamma}{\alpha}(Z_{t}^{*})^{2\kappa}\Big{)}^{-\frac{\delta}{2+\delta}}.$
The proof of the claim is divided into the following two cases.
$\begin{split}&\textbf{Case 1:}\ y_{t}\geq\frac{\alpha\delta}{2}t,\ \text{ if }\ y_{s}\leq A_{t},\ \text{ for all }\ s\in[0,t];\\ &\textbf{Case 2:}\ y_{t}\geq A_{t},\ \text{ if there exists $s\in[0,t]$ such that }\ y_{s}\geq A_{t}.\end{split}$
Case 1 is a direct consequence of (4.10). To prove Case 2, we assume otherwise that $y_{t}<A_{t}$. Note that $y_{\cdot}$ is continuous in $[s,t]$ and $y_{s}\geq A_{t}$, let
$r:=\inf\{u\in[s,t]:y_{v}\leq A_{t},\text{ for all }v\in[u,t]\},$
we conclude that $s\leq r<t$ and
$y_{u}\leq A_{t},\text{ for all }u\in[r,t],\ \text{ and }\ A_{t}=y_{r}>y_{t}.$
Then the mean value theorem indicates
$\frac{\mathop{}\!\mathrm{d}y_{u}}{\mathop{}\!\mathrm{d}u}\Big{|}_{u=u_{0}}=\frac{y_{t}-y_{r}}{t-r}<0,\ \text{ for some }\ u_{0}\in[r,t],$
which gives a contradiction to (4.10) and the fact that $y_{u_{0}}\leq A_{t}$. We complete the proof of the claim.
By (4.11), we know that, for any $p\geq 1$ and $t>0$,
$\begin{split}|X_{t}^{\mu,\xi}|^{p}&\leq 2^{p-1}(|Y_{t}|^{p}+|Z_{t}|^{p})\\ &\leq 2^{p-1}(y_{t}^{-\frac{p}{\delta}}+(Z_{t}^{*})^{p})\\ &\leq 2^{p-1}\max\bigg{\{}\Big{(}\frac{\alpha\delta}{2}t\Big{)}^{-\frac{p}{\delta}},\Big{(}\frac{\theta}{\alpha}(\mu_{t}^{*})^{2}+\frac{2\beta+\gamma}{\alpha}+\frac{\gamma}{\alpha}(Z_{t}^{*})^{2\kappa}\Big{)}^{\frac{p}{2+\delta}}\bigg{\}}+2^{p-1}(Z_{t}^{*})^{p}\\ &\leq C(p)\Big{(}t^{-\frac{p}{\delta}}+(\mu_{t}^{*})^{\frac{2p}{2+\delta}}+1+(Z_{t}^{*})^{\kappa p}\Big{)}.\end{split}$
Assumption 1 (v) and the Burkholder-Davis-Gundy (BDG) inequality imply that there exists $C(\kappa,p)>0$ such that
$\mathbf{E}\big{[}(Z_{t}^{*})^{\kappa p}\big{]}\leq C(\kappa,p)\mathbf{E}\bigg{[}\Big{(}\int_{0}^{t}|\sigma(X_{s}^{\mu,\xi},\mu_{s})|^{2}\mathop{}\!\mathrm{d}s\Big{)}^{\frac{\kappa p}{2}}\bigg{]}\leq C(\kappa,p)\cdot(\overline{\sigma}t)^{\frac{\kappa p}{2}}.$
Thus, for any $p\geq 2$, there is some constant $C(p)>0$ satisfying, for all $t>0$,
$\begin{split}\big{(}\mathbf{E}\big{[}|X_{t}^{\mu,\xi}|^{p}\big{]}\big{)}^{1/p}\leq C(p)\Big{(}t^{-\frac{1}{\delta}}+t^{\frac{\kappa}{2}}+\sup_{s\in[0,t]}\|\mu_{s}\|_{2}^{\frac{2}{2+\delta}}\Big{)}.\end{split}$
Next, we give the $p$-th moment estimates for the solutions of (1.1) when the initial conditions are square integrable.
Lemma 4.7.
Suppose Assumption 1 holds. If $\xi\in L^{2}(\Omega,\mathcal{F}_{0},\mathbf{P})$, then $X_{t}^{\xi}\in L^{p}(\Omega,\mathcal{F}_{t},\mathbf{P})$ for all $p\geq 2$, $t>0$. Moreover, for any $p\geq 2$, there exists a constant $C(p)\geq 1$ such that, for all $t\geq 1$,
$\big{(}\mathbf{E}\big{[}|X_{t}^{\xi}|^{p}\big{]}\big{)}^{1/p}\leq\mathrm{e}^{-(t-1)}\big{(}\mathbf{E}[|\xi|^{2}]\big{)}^{1/2}+C(p).$
Proof.
Denote $\mu_{t}:=\mathcal{L}(X_{t}^{\xi})$ for $t\geq 0$, and by Lemma 4.5 and Lemma 4.6, we know that for any $t>0$,
$\big{(}\mathbf{E}\big{[}|X_{t}^{\xi}|^{p}\big{]}\big{)}^{1/p}\leq C(p)\Big{(}t^{-\frac{1}{\delta}}+t^{\frac{\kappa}{2}}+\big{(}\mathbf{E}[|\xi|^{2}]\big{)}^{\frac{1}{2+\delta}}\Big{)}.$ (4.12)
Then for any $t\geq 1$, Lemma 4.5 and (4.12) yield
$\big{(}\mathbf{E}\big{[}|X_{t}^{\xi}|^{p}\big{]}\big{)}^{1/p}\leq\mathrm{e}^{-(t-1)}\big{(}\mathbf{E}\big{[}|X_{1}^{\xi}|^{p}\big{]}\big{)}^{1/p}+C(p)\leq e^{-(t-1)}C(p)\big{(}\mathbf{E}[|\xi|^{2}]\big{)}^{\frac{1}{2+\delta}}+C(p).$
Apply again Young’s inequality by noticing $\frac{1}{(2+\delta)/2}+\frac{1}{(2+\delta)/\delta}=1$, we complete the proof. ∎
Now we are ready to prove the coming down from infinity under the hyper-dissipative condition.
Theorem 4.8.
Suppose Assumption 1 holds. Let $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ be the semigroup of the McKean-Vlasov SDE (1.1). Then for any $t>0$, $P_{t}^{*}$ is a compact map. Moreover, there exists a nonempty set $K\subset\mathcal{P}_{\infty}(\mathbb{R})$ such that
  1. (i)
    $K$ is a connected compact set in $\mathcal{P}_{2}(\mathbb{R})$;
  2. (ii)
    $P_{t}^{*}K\subset K$ for all $t\geq 2$;
  3. (iii)
    $K$ is order bounded in $\mathcal{P}_{2}(\mathbb{R})$, i.e., there exist $\underline{\mu},\overline{\mu}\in\mathcal{P}_{2}(\mathbb{R})$ such that $\underline{\mu}\leq_{\mathrm{st}}K\leq_{\mathrm{st}}\overline{\mu}$;
  4. (iv)
    $K$ absorbs all bounded sets in $\mathcal{P}_{2}(\mathbb{R})$, i.e., for any bounded set $B\subset\mathcal{P}_{2}(\mathbb{R})$, there exists $T>0$ such that $P_{t}^{*}B\subset K$ for all $t\geq T$.
Proof.
We first show the compactness of $P_{t}^{*}$ for any $t>0$. For any bounded set $B$ in $\mathcal{P}_{2}(\mathbb{R})$, it follows from (4.12) in Lemma 4.7 that $P_{t}^{*}B$ is a bounded subset of $\mathcal{P}_{p}(\mathbb{R})$ for all $p>2$. By Corollary 3.3, $P_{t}^{*}B$ is order bounded in $\mathcal{P}_{2}(\mathbb{R})$ with respect to the stochastic order. Then Proposition 3.4 implies $P_{t}^{*}B$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$, so $P_{t}^{*}$ is compact.
Next, we prove (i)-(iv). Set
$K_{p}:=\Big{\{}\mu\in\mathcal{P}_{p}(\mathbb{R}):\|\mu\|_{p}\leq\frac{e}{e-1}C(p)\Big{\}},\ \text{ and }\ K:=\bigcap_{p\geq 2}K_{p},$ (4.13)
where the constant $C(p)$ come from Lemma 4.7. Clearly, $\delta_{0}\in K$, so $K$ is nonempty. We now show (i)-connectedness. Because $K$ is convex and the operation of convex combinations is continuous in $\mathcal{P}_{2}(\mathbb{R})$, we obtain $K$ is connected. Then we want to show (iii). Since $K$ is bounded in $\mathcal{P}_{p}(\mathbb{R})$ for all $p>2$, it follows from Corollary 3.3 that $K$ is order bounded. By Proposition 3.4, to prove (i)-compactness, we only need to show $K$ is closed. Suppose a $\mathcal{P}_{2}$-Cauchy sequence $\{\mu_{n}\}_{n\geq 1}\subset K$ converges to a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$, and it remains to check $\mu\in K$, i.e., $\mu\in K_{p}$ for all $p\geq 2$. By [68, Theorem 6.9], $\mu_{n}$ weakly converges to $\mu$. For any $p\geq 2$, since $\{\mu_{n}\}_{n\geq 1}\subset K_{p}$, then
$\begin{split}\int_{\mathbb{R}}|x|^{p}\mathop{}\!\mathrm{d}\mu(x)&=\lim_{N\to\infty}\int_{\mathbb{R}}(|x|^{p}\wedge N)\mathop{}\!\mathrm{d}\mu(x)\\ &=\lim_{N\to\infty}\lim_{n\to\infty}\int_{\mathbb{R}}(|x|^{p}\wedge N)\mathop{}\!\mathrm{d}\mu_{n}(x)\\ &\leq\sup_{n\geq 1}\int_{\mathbb{R}}|x|^{p}\mathop{}\!\mathrm{d}\mu_{n}(x)\\ &\leq\frac{e}{e-1}C(p).\end{split}$
This gives $\mu\in K_{p}$. Then $\mu\in K$ as $p\geq 2$ is arbitrary.
The statements (ii)(iv) can be easily checked by Lemma 4.7 and the construction of $K$ in (4.13). The proof is completed. ∎
4.4 Proof of Theorem 2.1 and Corollary 2.2
We first give the following lemma, which is a direct consequence of Feng-Qu-Zhao [33, Theorem 4.8]. Let $C_{b}(\mathbb{R};\mathcal{P}_{2}(\mathbb{R}))$ be the collection of all bounded continuous paths $\{\nu_{t}\}_{t\in\mathbb{R}}$ in $\mathcal{P}_{2}(\mathbb{R})$.
Lemma 4.9.
Suppose Assumption 1 holds, and let $\{\nu_{t}\}_{t\in\mathbb{R}}\in C_{b}(\mathbb{R};\mathcal{P}_{2}(\mathbb{R}))$. Consider the following time-inhomogeneous SDE
$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\nu_{t})(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (4.14)
Let $X_{t}^{s,\mu}$ be the unique solution of equation (4.14) at time $t$ with starting time $s$ and initial distribution $\mu$ and $P^{*}(t,s)$ be the dual semigroup of equation (4.14), i.e., $P^{*}(t,s)\mu:=\mathcal{L}(X_{t}^{s,\mu})$. Then the equation (4.14) has a unique entrance measure $\{\rho^{\nu}_{t}\}_{t\in\mathbb{R}}$ in $C_{b}(\mathbb{R};\mathcal{P}_{2}(\mathbb{R}))$, i.e., $P^{*}(t,s)\rho^{\nu}_{s}=\rho^{\nu}_{t}$ for all $t\geq s$ and there exist $C,\lambda>0$ such that for all $\mu\in\mathcal{P}_{2}(\mathbb{R})$, $t\geq s$, we have
$d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\rho^{\nu}_{t}\big{)}\leq C(1+\|\mu\|_{2}^{2})e^{-\lambda(t-s)},$ (4.15)
where $d_{2}(\cdot,\cdot)$ is defined by
$d_{2}(\mu_{1},\mu_{2}):=\int_{\mathbb{R}}(1+|x|^{2})\mathop{}\!\mathrm{d}|\mu_{1}-\mu_{2}|(x),\ \text{ for all }\ \mu_{1},\mu_{2}\in\mathcal{P}_{2}(\mathbb{R}).$
In particular,
  1. (i)
    if $\{\nu_{t}\}_{t\in\mathbb{R}}$ is an entrance measure of (1.1), i.e., $P^{*}_{t}\nu_{s}=\nu_{t+s}$ for all $t\geq 0,s\in\mathbb{R}$, then $\rho^{\nu}_{t}=\nu_{t}$ for all $t\in\mathbb{R}$;
  2. (ii)
    if $\{\nu_{t}\}_{t\in\mathbb{R}}$ is a singleton $\{\nu\}\subset\mathcal{P}_{2}(\mathbb{R})$, then the unique entrance measure $\{\rho^{\nu}_{t}\}_{t\in\mathbb{R}}$ is also a singleton $\{\rho\}\subset\mathcal{P}_{2}(\mathbb{R})$ such that $\rho$ is the unique invariant measure of (4.14) in $\mathcal{P}_{2}(\mathbb{R})$.
The next lemma unveils the totally ordered structure of the set of invariant measures.
Lemma 4.10.
Suppose Assumption 1 holds, and let $\mu_{1},\mu_{2}$ be two invariant measures of the equation (1.1). If $\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{1}(x)\leq\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{2}(x)$, then $\mu_{1}\leq_{\mathrm{st}}\mu_{2}$.
Proof.
Let $m_{i}=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{i}(x),i=1,2$. Now we consider the following SDEs
$\mathop{}\!\mathrm{d}X_{t}^{i}=\big{(}-V^{\prime}(X_{t}^{i})-\theta X_{t}^{i}+\theta m_{i}\big{)}\mathop{}\!\mathrm{d}t+\sigma(X_{t}^{i})\mathop{}\!\mathrm{d}B_{t},\quad i=1,2.$ (4.16)
Obviously, $\mu_{i}$ is an invariant measures of (4.16). Denote by $X_{t}^{i,x}$ the solution of (4.16) with initial value $x$ and starting time 00, $i=1,2$. Since $m_{1}\leq m_{2}$, it follows from the comparison theorem (Theorem 4.3) that
$\mathcal{L}(X_{t}^{1,0})\leq_{\mathrm{st}}\mathcal{L}(X_{t}^{2,0})\ \text{ forall $t\geq 0$}.$ (4.17)
By Lemma 4.9 (ii), $\mu_{i}$ is the unique invariant measure of (4.16) and
$d_{2}(\mathcal{L}(X_{t}^{i,0}),\mu_{i})\to 0\ \text{ as $t\to\infty$, for $i=1,2$},$ (4.18)
Next, we want to show the convergence in $\mathcal{P}_{2}(\mathbb{R})$. Observe that, for any bounded continuous function $f:\mathbb{R}\to\mathbb{R}$, with $|f|_{\infty}:=\sup_{x\in\mathbb{R}}|f(x)|$,
$\Big{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{1}(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{2}(x)\Big{|}\leq|f|_{\infty}d_{2}\big{(}\nu_{1},\nu_{2}\big{)},$ (4.19)
and
$\Big{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1}(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{2}(x)\Big{|}\leq d_{2}\big{(}\nu_{1},\nu_{2}\big{)}.$ (4.20)
Combine (4.18)-(4.20), and it follows from the equivalent characterization of $\mathcal{W}_{2}$-convergence (see Villani [68, Theorem 6.9]) that
$\mathcal{W}_{2}(\mathcal{L}(X_{t}^{i,0}),\mu_{i})\to 0\ \text{ as $t\to\infty$, for $i=1,2$}.$ (4.21)
Together with (4.17)(4.21), the closedness of stochastic order in $\mathcal{P}_{2}(\mathbb{R})$ implies that $\mu_{1}\leq_{\mathrm{st}}\mu_{2}$. ∎
Remark 4.3.
In general, we can consider two SDEs as below,
$\mathop{}\!\mathrm{d}X_{t}^{1}=\big{(}-V^{\prime}(X_{t}^{1})-\theta X_{t}^{1}+\theta m_{1}\big{)}\mathop{}\!\mathrm{d}t+\sigma(X_{t}^{1})\mathop{}\!\mathrm{d}B_{t},$ (4.22)
$\mathop{}\!\mathrm{d}X_{t}^{2}=\big{(}-V^{\prime}(X_{t}^{2})-\theta X_{t}^{2}+\theta m_{2}\big{)}\mathop{}\!\mathrm{d}t+\sigma(X_{t}^{2})\mathop{}\!\mathrm{d}B_{t}.$ (4.23)
By the proof of Lemma 4.10, whenever $m_{1}\leq m_{2}$, the invariant measures $\mu_{1}$, $\mu_{2}$ of the equations (4.22), (4.23) are order related, precisely $\mu_{1}\leq_{\mathrm{st}}\mu_{2}$.
Now we are in a position to prove Theorem 2.1 and Corollary 2.2.
Proof of Theorem 2.1.
(i). By Lemma 4.10, we see $\mathcal{M}$ is totally ordered with respect to the stochastic order. Take $K\subset\mathcal{P}_{\infty}(\mathbb{R})$ as the set in (4.13), and Theorem 4.8 (iv) gives $\mathcal{M}\subset K\subset\mathcal{P}_{\infty}(\mathbb{R})$. The finiteness of $\mathcal{M}$ is postponed to the end of the proof.
(ii). By Theorem 4.4, the semigroup $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is an order-preserving semiflow. Theorem 4.8 gives a eventually positively invariant, order bounded, closed set $K\subset\mathcal{P}_{2}(\mathbb{R})$, and this set $K$ attracts all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. Now, apply Theorem 4.2, whose statements (i)(ii) say that, there are two invariant measures $\underline{\nu},\overline{\nu}\in\mathcal{P}_{2}(\mathbb{R})$ such that the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$. By the definition of global attractor, one has $\mathcal{M}\subset\omega(K)$. Obviously, $\underline{\nu}=\inf_{\leq_{\mathrm{st}}}\mathcal{M}$ and $\overline{\nu}=\sup_{\leq_{\mathrm{st}}}\mathcal{M}$. Next, we need to show the global convergence to the order interval $[\underline{\nu},\overline{\nu}]$. Since $\omega(K)$ is the global attractor, the inclusion $\omega(K)\subset[\underline{\nu},\overline{\nu}]$ tells us, for all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\underline{\nu},\overline{\nu}])\to 0$ as $t\to\infty$.
(iii). This is a consequence of Theorem 4.2 (iii).
(i)-finiteness. Now we prove that $\mathcal{M}$ is a finite set. Let $\{\mu_{m}\}_{m\in\mathbb{R}}$ be probability measures given in (1.4), i.e.,
$\frac{\mathop{}\!\mathrm{d}\mu_{m}}{\mathop{}\!\mathrm{d}x}\propto\frac{1}{\sigma^{2}(x)}\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}.$
By a simple calculation, $\mu_{m}$ solves the following Fokker-Planck equation
$\mathcal{L}^{*}_{m}\mu_{m}=0,$
where $\mathcal{L}_{m}^{*}$ is the dual operator of the following $\mathcal{L}_{m}$:
$\mathcal{L}_{m}\varphi(x)=\frac{1}{2}\sigma^{2}(x)\partial^{2}_{x}\varphi(x)+\big{(}-V^{\prime}(x)-\theta x+\theta m\big{)}\partial_{x}\varphi(x).$
Hence, $\mu_{m}$ is an invariant measure of the following SDE
$\mathop{}\!\mathrm{d}X_{t}=\big{(}-V^{\prime}(X_{t})-\theta X_{t}+\theta m\big{)}\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},$ (4.24)
and it is unique (see Meyn-Tweedie [54, 55] or Lemma 4.9). Thus, a measure $\mu$ is an invariant measure of (1.1) if and only if
$\mu=\mu_{m},\ \text{ and }\ \int_{\mathbb{R}}(x-m)\mathop{}\!\mathrm{d}\mu(x)=0\ \text{ for some $m\in\mathbb{R}$}.$ (4.25)
It is equivalent to
$\mu=\mu_{m},\ \text{ and }\ \int_{\mathbb{R}}\frac{x-m}{\sigma^{2}(x)}\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x=0\ \text{ for some $m\in\mathbb{R}$}.$
Consider the following self-consistency function,
$F(m):=\int_{\mathbb{R}}\frac{x-m}{\sigma^{2}(x)}\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x.$ (4.26)
Then the number of zeros of $F$ is exactly the number of invariant measures of the equation (1.1), and each zero of $F$ is the mean value of the corresponding invariant measure. We assume there are infinitely many invariant measures, and argue by contradiction. By the statement (ii), the set of invariant measures $\mathcal{M}$ is contained in an order interval $[\underline{\nu},\overline{\nu}]$. So the boundedness of $[\underline{\nu},\overline{\nu}]$ in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition 3.4) entails boundedness of the zero set of $F$ in $\mathbb{R}$. Since invariant measures are infinitely many, the zero set of $F$ has an accumulation point. By Assumption 1 (iii)(v), we have $F$ is an analytic function, and this implies $F$ is identically zero. This is impossible, and we complete the proof. ∎
Remark 4.4.
In Remark 2.2, we show that the set of invariant measures cannot attract all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$. Here, by Theorem 4.2 (i)(ii) and Theorem 4.8 (i), there exists a connected compact invariant set (the global attractor) within $[\underline{\nu},\overline{\nu}]$ which attracts all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
Proof of Corollary 2.2.
The uniqueness of invariant measure enatils that $\underline{\nu}=\nu=\overline{\nu}$. Then the antisymmetry of stochastic order implies that $[\underline{\nu},\overline{\nu}]=\{\nu\}$. Hence, this corollary is a direct consequence of Theorem 2.1 (ii). ∎
5 Unbounded Open Neighbourhood Attracted by Supremum (Infimum) Invariant Measure
This section attempts to prove Theorem 2.3 and Corollary 2.4. In Section 5.1, we find the number of invariant measures of the equation (1.1) is bounded by the zero-crossing number of the function $V^{\prime}$. In Section 5.2, we show the existence of positively invariant neighbourhoods of invariant measures, for both the semigroup $P_{t}^{*}$ and the measure-iterating map $\Psi$. After providing a structure theorem of local attractors of order-preserving semigroups in Section 5.3, we proceed to prove Theorem 2.3 and Corollary 2.4 by tracing solutions of (1.1) backward along connecting orbits.
5.1 Number of invariant measures of granular media equations
In this subsection, we provide the prior information on the number of invariant measures. Assuming $V^{\prime}$ has $(2n-1)$ simple zeros, Alecio [1] shows that the equation (1.1) has only one invariant measure if $\left\lvert\sigma\right\rvert$ is above a certain upper critical threshold, and has $(2n-1)$ invariant measures if $\left\lvert\sigma\right\rvert$ is below another lower critical threshold. There is a gap between these two critical thresholds, and the information is missing in the intermediate stage. Here, for any non-degenerate $\sigma$, we prove that the number of invariant measures of the equation (1.1) does not exceed the zero-crossing number of $V^{\prime}$ in Propoaition 5.3. This facilitates the quantitative results in our main theorems.
Let $f:\mathbb{R}\to\mathbb{R}$ be a function. Recall that its zero-crossing number $Z(f)$ is the maximal integer $N$ such that, there exists an increasing sequence $x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}$ satisfying $f(x_{n})f(x_{n+1})<0$ for all $n=1,2,\dots,N$. It is easy to check that, if $V^{\prime}$ has $(2n-1)$ simple zeros, then $Z(V^{\prime})=2n-1$. Next, we will see the zero-crossing number of the function $V^{\prime}$ is essential to determine the number of invariant measures. We begin with a simple case where the zero-crossing number is one.
Lemma 5.1.
Suppose Assumption 1 holds. If the function $f:\mathbb{R}\to\mathbb{R}$ is continuous with polynomial growth and has zero-crossing number $Z(f)=1$, then the function $F:\mathbb{R}\to\mathbb{R}$ defined by
$F(m):=\int_{\mathbb{R}}f(x)\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x$
has exactly one zero.
Proof.
Since $f$ has polynomial growth, $F$ is well-defined. Given $Z(f)=1$, either case of the following two is true.
  • There is some $x_{0}\in\mathbb{R}$ such that $f(x)\leq 0$ on $(-\infty,x_{0})$ and $f(x)\geq 0$ on $(x_{0},\infty)$;
  • There is some $x_{0}\in\mathbb{R}$ such that $f(x)\geq 0$ on $(-\infty,x_{0})$ and $f(x)\leq 0$ on $(x_{0},\infty)$.
It suffices to consider the first case, and by definition of the zero-crossing number, there are $x_{-}<x_{0}<x_{+}$ such that
$f(x_{-})<0,\ \ \ f(x_{+})>0.$
Now define another function $\tilde{F}:\mathbb{R}\to\mathbb{R}$ by
$\tilde{F}(m):=\int_{\mathbb{R}}f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x,$
and we first show $\tilde{F}$ has only one zero. Denote $\Sigma(x):=\int_{x_{0}}^{x}\frac{1}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y$, and by Assumption 1 (v), we observe $\Sigma(x)<0$ on $(-\infty,x_{0})$, $\Sigma(x)>0$ on $(x_{0},\infty)$. A straightforward differentiation gives
$\tilde{F}^{\prime}(m)$ $=\int_{\mathbb{R}}2\theta\Sigma(x)f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x$
$=\int_{\mathbb{R}}2\theta\Sigma(x)f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\exp\big{\{}2\theta m\Sigma(x)\big{\}}\mathop{}\!\mathrm{d}x.$
For $m\geq 0$, since $\Sigma(x)f(x)\geq 0$ for all $x\in\mathbb{R}$, we have
$\tilde{F}^{\prime}(m)$ $\geq\int_{x_{0}}^{\infty}2\theta\Sigma(x)f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\exp\big{\{}2\theta m\Sigma(x)\big{\}}\mathop{}\!\mathrm{d}x$ (5.1)
$\geq\int_{x_{0}}^{\infty}2\theta\Sigma(x)f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x=:\lambda_{+},$
where $\lambda_{+}$ is a constant independent of $m$. Noticing $x_{+}>x_{0}$, $f(x_{+})>0$, $\Sigma(x_{+})>0$ and $\Sigma f$ is continuous, it follows $\lambda_{+}>0$. And in the same manner, for $m<0$, we see
$\tilde{F}^{\prime}(m)\geq\int_{-\infty}^{x_{0}}2\theta\Sigma(x)f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x=:\lambda_{-}>0.$ (5.2)
Combining (5.1) and (5.2) gives, for all $m\in\mathbb{R}$,
$\tilde{F}^{\prime}(m)\geq\min\{\lambda_{+},\lambda_{-}\}>0,$
so the function $\tilde{F}$ has only one zero. To show $F$ has exactly one zero, we just simply note that $\tilde{F}(m)=F(m)G(m)$ for a positive function $G$ given by
$G(m)=\exp\bigg{\{}2\int_{0}^{x_{0}}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}.$
Hence, the function $F$ has only one zero as well. ∎
The following lemma extends to the general case of finite zero-crossings.
Lemma 5.2.
Suppose Assumption 1 holds. If the function $f:\mathbb{R}\to\mathbb{R}$ is continuous with polynomial growth and has zero-crossing number $Z(f)=n$, then the function $F:\mathbb{R}\to\mathbb{R}$ defined by
$F(m):=\int_{\mathbb{R}}f(x)\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x$
has at most $n$ zeros.
Proof.
By Lemma 5.1, the conclusion is true for $n=1$, and we prove by induction. Assume the conclusion is true for $n$. If the function $f$ is continuous and $Z(f)=n+1$, then we may assume there is an increasing sequence $-\infty=x_{-1}<x_{0}<\dots<x_{n}<x_{n+1}=\infty$ such that, $f$ changes its sign on any two adjacent intervals $(x_{k-1},x_{k})$, $k=0,1,\dots,n+1$. Define $\tilde{F}:\mathbb{R}\to\mathbb{R}$ by
$\tilde{F}(m):=\int_{\mathbb{R}}f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x.$
Taking the derivative of $\tilde{F}$ yields
$\tilde{F}^{\prime}(m)$ $=\int_{\mathbb{R}}2\theta\Sigma(x)f(x)\exp\bigg{\{}-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x$
$=G(m)\cdot\int_{\mathbb{R}}2\theta\Sigma(x)f(x)\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x,$
where
$\Sigma(x)=\int_{x_{0}}^{x}\frac{1}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y,\quad G(m)=\exp\bigg{\{}2\int_{0}^{x_{0}}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}.$
Note the zero crossing number of $\Sigma f$ is $n$ and $G$ is positive, so $\tilde{F}^{\prime}$ has at most $n$ zeros by the induction. This implies $\tilde{F}$ has at most $(n+1)$ zeros, and since $F=G\tilde{F}$, the function $F$ has at most $(n+1)$ zeros. We complete the proof. ∎
Now a bound of the number of invariant measures is given as below.
Proposition 5.3.
Suppose Assumption 1 holds. If the function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ has zero-crossing number $Z(V^{\prime})=2n-1$, then the equation (1.1) has at most $(2n-1)$ invariant measures.
Proof.
Consider the self-consistency function defined in (4.26),
$F(m):=\int_{\mathbb{R}}\frac{x-m}{\sigma^{2}(x)}\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x.$
Then the number of zeros of $F$ is exactly the number of invariant measures of the equation (1.1). Use integration by parts, and we observe
$F(m)=-\int_{\mathbb{R}}\frac{V^{\prime}(x)}{\theta\sigma^{2}(x)}\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}\mathop{}\!\mathrm{d}x.$
Assumption 1 (iv)(v) tell us $\theta>0$ and $\sigma^{2}(x)>0$ for all $x\in\mathbb{R}$, so $Z(V^{\prime}/\theta\sigma^{2})=Z(V^{\prime})=2n-1$. Lemma 5.2 implies that $F$ has at most $(2n-1)$ zeros, and thus, the equation (1.1) has at most $(2n-1)$ invariant measures. ∎
5.2 Positively invariant neighbourhoods of invariant measures
In this subsection, we prove the existence of positively invariant $\mathcal{P}_{2}(\mathbb{R})$-neighbourhoods of invariant measures under $P^{*}$ when assuming the locally dissipative condition. The existence of positively invariant $\mathcal{P}_{2}(\mathbb{R})$-neighbourhoods is also obtained for the measure-iterating map $\Psi$ defined in (5.13).
Theorem 5.4.
Suppose that Assumption 1 holds. If the locally dissipative condition holds at $a\in\mathbb{R}$ with configuration $(r_{a},g_{a})$, then the equation (1.1) has an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. Moreover, $B(\delta_{a},r_{a})$ is positively invariant sets, i.e.,
$P_{t}^{*}B(\delta_{a},r_{a})\subset B(\delta_{a},r_{a}),\ \text{ for all }\ t\geq 0.$ (5.3)
Proof.
By (ii) in the definition of local dissipation (Definition 2.3), we know that there exists $0<\tilde{r}_{a}<r_{a}$ such that
$g_{a}(z,r_{a}^{2})>0,\ \text{ for all }z\geq\tilde{r}_{a}^{2}.$ (5.4)
To prove (5.3), it suffices to prove that $\overline{B(\delta_{a},r)}$ is positively invariant for all $\tilde{r}_{a}\leq r<r_{a}$, which implies that $B(\delta_{a},r)$ is positively invariant for all $\tilde{r}_{a}<r\leq r_{a}$.
For any fixed $\tilde{r}_{a}\leq r<r_{a}$ and $\mu\in\overline{B(\delta_{a},r)}$, let
$\mathcal{T}_{\mu}:=\sup\big{\{}t\geq 0:P_{s}^{*}\mu\in\overline{B(\delta_{a},r)}\ \text{ for all }\ 0\leq s\leq t\big{\}}.$
To prove that $\overline{B(\delta_{a},r)}$ is positively invariant, it suffices to show that $\mathcal{T}_{\mu}=\infty$. By Theorem 4.4, we know that $P^{*}:\mathbb{R}\times\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ is jointly continuous, so $\alpha_{t}:=\mathcal{W}_{2}^{2}(P^{*}_{t}\mu,\delta_{a})$ is continuous in $t$. Then it is equivalent to show that
$\mathcal{T}_{\mu}=\sup\big{\{}t\geq 0:\alpha_{s}\leq r^{2}\ \text{ for all }\ 0\leq s\leq t\big{\}}=\infty.$
Otherwise, if $\mathcal{T}_{\mu}<\infty$, then it follows from the continuity of $\alpha$ and definition of $\mathcal{T}_{\mu}$ that there exists $T_{0}\geq\mathcal{T}_{\mu}$ and $\epsilon_{0}>0$ such that
$\alpha_{T_{0}}=r^{2},\ \text{ and }\ r^{2}<\alpha_{t}\leq r_{a}^{2}\ \text{ for all }\ T_{0}<t\leq T_{0}+\epsilon_{0}.$ (5.5)
Let $X_{t}^{\xi}$ be the solution of (1.1) with intial distribution $\mathcal{L}(\xi)=\mu$ and starting time 00. Then $P_{t}^{*}\mu=\mathcal{L}(X_{t}^{\xi})$ for all $t\geq 0$. For any $a\in\mathbb{R}$, denote $\tau^{a}:\mathbb{R}\to\mathbb{R}$ by the translation $x\mapsto x+a$ and $\tau^{a}_{\#}:\mathcal{P}(\mathbb{R})\to\mathcal{P}(\mathbb{R})$ by $\tau^{a}_{\#}\mu=\mu\circ(\tau^{a})^{-1}$. Applying Itô’s formula to $|X_{t}^{\xi}-a|^{2}$ on $[T_{0},T_{0}+\epsilon_{0}]$, we have for any $T_{0}<t\leq T_{0}+\epsilon_{0}$,
$\begin{split}|X_{t}^{\xi}-a|^{2}&=|X_{T_{0}}^{\xi}-a|^{2}+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s}\\ &\ \ \ \ +\int_{T_{0}}^{t}\big{(}-2(X_{s}^{\xi}-a)V^{\prime}(X_{s}^{\xi})-2(X_{s}^{\xi}-a)(W^{\prime}*P^{*}_{s}\mu)(X_{s}^{\xi})+\sigma^{2}(X_{s}^{\xi})\big{)}\mathop{}\!\mathrm{d}s\\ &=|X_{T_{0}}^{\xi}-a|^{2}+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s}\\ &\ \ \ \ +\int_{T_{0}}^{t}\big{(}-2(X_{s}^{\xi}-a)V^{\prime}(X_{s}^{\xi})-2(X_{s}^{\xi}-a)(W^{\prime}*\tau^{-a}_{\#}P^{*}_{s}\mu)(X_{s}^{\xi}-a)+\sigma^{2}(X_{s}^{\xi})\big{)}\mathop{}\!\mathrm{d}s\\ &\leq|X_{T_{0}}^{\xi}-a|^{2}-\int_{T_{0}}^{t}g_{a}\big{(}|X_{s}^{\xi}-a|^{2},r_{a}^{2}\big{)}\mathop{}\!\mathrm{d}s+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s},\end{split}$ (5.6)
where we have used the fact
$\|\tau^{-a}_{\#}P^{*}_{s}\mu\|_{2}^{2}=\int_{\mathbb{R}}|x-a|^{2}\mathop{}\!\mathrm{d}P^{*}_{s}\mu(x)=\mathcal{W}_{2}^{2}(P^{*}_{s}\mu,\delta_{a})=\alpha_{s}\in[r^{2},r_{a}^{2}]\ \text{ for all $s\in[T_{0},T_{0}+\varepsilon]$}.$
Taking expectation on both sides and by the convexity of $g_{a}(\cdot,r_{a}^{2})$, we get
$\alpha_{T_{0}+\epsilon_{0}}\leq r^{2}-\int_{T_{0}}^{T_{0}+\epsilon_{0}}g_{a}(\alpha_{t},r_{a}^{2})\mathop{}\!\mathrm{d}t.$ (5.7)
Then by (5.4), (5.5) and (5.7), we conclude that $\alpha_{T_{0}+\epsilon_{0}}<r^{2}$ which contradicts (5.5). Hence,
$P_{t}^{*}\overline{B(\delta_{a},r)}\subset\overline{B(\delta_{a},r)},\ \text{ for all }\ \tilde{r}\leq r<r_{a},$ (5.8)
and (5.3) follows.
It remains to prove there exists an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. We first give the following two claims.
Claim 1: Let $S\subset\mathcal{P}_{2}(\mathbb{R})$ be a bounded closed convex positively invariant set under $P^{*}$. If $P_{t}^{*}$ is compact for some $t>0$, then there exists a fixed point $\mu_{t}$ of $P_{t}^{*}$ in $S$, i.e., $P^{*}_{t}\mu_{t}=\mu_{t}$.
Claim 2: Let $S\subset\mathcal{P}_{2}(\mathbb{R})$ be a bounded closed convex positively invariant set under $P^{*}$, and $P_{t}^{*}$ is compact for all $t>0$, then there exists an invariant measure $\mu$ in $S$, i.e., $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
Proof of Claim 1: Let $S_{t}=P_{t}^{*}S$ and $\overline{\mathrm{Conv}(S_{t})}$ be the closed convex hull of $S_{t}$. Since $P_{t}^{*}$ is compact, it follows $S_{t}$ is relatively compact. Noting also that $S$ is closed convex positively invariant, we have $\overline{\mathrm{Conv}(S_{t})}\subset S$ is compact in $\mathcal{P}_{2}(\mathbb{R})$, and
$P_{t}^{*}\overline{\mathrm{Conv}(S_{t})}\subset P_{t}^{*}S=S_{t}\subset\overline{\mathrm{Conv}(S_{t})}.$
Note that $\overline{\mathrm{Conv}(S_{t})}$ being compact convex in $\mathcal{P}_{2}(\mathbb{R})$ entails that it is compact convex in $\mathcal{P}_{1}(\mathbb{R})$. Let us consider the space $\mathcal{M}_{1}(\mathbb{R})$ of finite signed Borel measures on $\mathbb{R}$ with finite first moments,
$\mathcal{M}_{1}(\mathbb{R}):=\left\{\mu\in\mathcal{M}(\mathbb{R}):\int_{\mathbb{R}}\left\lvert x\right\rvert\mathop{}\!\mathrm{d}\left\lvert\mu\right\rvert(x)<\infty\right\}$
with the following norm:
[…]
where a $1$-Lipschitz function $f$ means $|f(x)-f(y)|\leq|x-y|$ for all $x,y\in\mathbb{R}$. By [48, Lemma 3.4 and Remark 3.2], we have the following properties of $\mathcal{M}_{1}(\mathbb{R})$:
$(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$ is a normed vector space; (5.10)
$\mathcal{P}_{1}(\mathbb{R})$ is a closed convex subset of $(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$. (5.11)
Hence, $\overline{\mathrm{Conv}(S_{t})}$ is a compact convex subset of a normed vector space. Then the Tychonoff fixed-point theorem yields that there exists $\mu_{t}\in\overline{\mathrm{Conv}(S_{t})}\subset S$ such that $P^{*}_{t}\mu_{t}=\mu_{t}$.
Proof of Claim 2: Fix a positive sequence $\{t_{n}\}_{n\geq 1}$ such that $t_{n}\downarrow 0$ as $n\to\infty$. For any $n\geq 1$, consider the continuous compact map $P^{*}_{t_{n}}:S\to S$. Then Claim 1 yields that there exists $\mu_{n}\in S$ such that $P^{*}_{t_{n}}\mu_{n}=\mu_{n}$. Let $k_{n}$ be an integer such that $k_{n}t_{n}>t_{1}$ for all $n\geq 1$. Then
$\mu_{n}=P_{k_{n}t_{n}}^{*}\mu_{n}=P_{t_{1}}^{*}(P_{k_{n}t_{n}-t_{1}}^{*}\mu_{n})\in P_{t_{1}}^{*}(P_{k_{n}t_{n}-t_{1}}^{*}S)\subset P_{t_{1}}^{*}S\subset S\ \text{ for all $n\geq 1$}.$
Since $P_{t_{1}}^{*}S\subset S$ is relatively compact and $S$ is closed, there exists a subsequence of $\{\mu_{n}\}_{n\geq 1}$, which still denotes by $\{\mu_{n}\}_{n\geq 1}$, and a $\mu\in S$ such that $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ as $n\to\infty$. Then we show that $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
In fact, for any $t>0$, we have $t=k_{n,t}t_{n}+\tau_{n,t}$, $k_{n,t}\in\mathbb{N}$ and $\tau_{n,t}\in[0,t_{n})$. Then,
$\mathcal{W}_{2}(P^{*}_{t}\mu,\mu)$ $\leq\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{t}\mu_{n},\mu)$
$=\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}(P^{*}_{k_{n,t}t_{n}}\mu_{n}),\mu)$
$=\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}\mu_{n},\mu).$
Since $\tau_{n,t}\to 0$ and $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ as $n\to\infty$, by virtue of the joint continuity of the $P^{*}$, we have $\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})\to 0$ and $\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}\mu_{n},\mu)\to 0$ as $n\to\infty$. Hence, we obtain that $P^{*}_{t}\mu=\mu$ and the claim is proved.
Now we are in a position to show that there exists an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. Since $P^{*}_{t}$ is compact for all $t>0$ (see Theorem 4.8) and $\overline{B(\delta_{a},\tilde{r}_{a})}$ is bounded closed convex positively invariant under $P^{*}$, by Claim 2, there exists an invariant measure $\mu\in\overline{B(\delta_{a},\tilde{r}_{a})}\subset B(\delta_{a},r_{a})$. Moreover, Lemma 4.7 shows that $\mu\in\mathcal{P}_{\infty}(\mathbb{R})$. Now, we conclude that the invariant measure $\mu\in B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. ∎
Now for any fixed $\mu\in\mathcal{P}_{2}(\mathbb{R})$, we consider the following SDE:
$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mu)(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (5.12)
Under Assumption 1, we know that SDE (5.12) has a unique invariant measure $\rho^{\mu}\in\mathcal{P}_{2}(\mathbb{R})$ (see e.g., Meyn-Tweedie [54, 55] or Lemma 4.9 (ii)). Define a measure-iterating map $\Psi:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by
[…]
Then we have the following theorem.
Theorem 5.5.
Suppose that Assumption 1 holds. If the equation (1.1) is locally dissipative at $a\in\mathbb{R}$ with configuration $(r_{a},g_{a})$, then we have $\Psi\big{(}\overline{B(\delta_{a},r_{a})}\big{)}\subset B(\delta_{a},r_{a})$.
Proof.
Note that there is some $\tilde{r}_{a}\in(0,r_{a})$ such that (5.4) holds. Now fix any $\mu\in\overline{B(\delta_{a},r_{a})}$. Let $\{X_{t}^{\mu,x}\}_{t\geq 0}$ be the solution of (5.12) with initial value $x\in\mathbb{R}$ and starting time 0. Similar to (5.6), applying Itô’s formula to $|X_{t}^{\mu,a}-a|^{2}$ yields
$|X_{t}^{\mu,a}-a|^{2}\leq-\int_{0}^{t}g_{a}\big{(}|X_{s}^{\mu,a}-a|^{2},r_{a}^{2}\big{)}\mathop{}\!\mathrm{d}s+2\int_{0}^{t}(X_{s}^{\mu,a}-a)\sigma(X_{s}^{\mu,a})\mathop{}\!\mathrm{d}B_{s}.$
Taking expectation on both sides and by the convexity of $g_{a}(\cdot,r_{a}^{2})$, we have that for any $t>0$,
$g_{a}\bigg{(}\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}\mathop{}\!\mathrm{d}s,r_{a}^{2}\bigg{)}\leq\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}g_{a}\big{(}|X_{s}^{\mu,a}-a|^{2},r_{a}^{2}\big{)}\big{]}\mathop{}\!\mathrm{d}s\leq-\frac{1}{t}\mathbf{E}[|X_{t}^{\mu,a}-a|^{2}]\leq 0.$ (5.14)
According to Lemma 4.9 (ii), we know that there exist $C_{\mu}>0,\lambda_{\mu}>0$ such that for any $t\geq 0$,
$d_{2}\big{(}\mathcal{L}(X_{t}^{\mu,a}),\Psi(\mu)\big{)}\leq C_{\mu}(1+|a|^{2})e^{-\lambda_{\mu}t},$
where
$d_{2}(\nu_{1},\nu_{2})=\int_{\mathbb{R}}(1+|x|^{2})\mathop{}\!\mathrm{d}|\nu_{1}-\nu_{2}|(x),$ (5.15)
Then
$\big{|}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}-\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}\big{|}=\Big{|}\int_{\mathbb{R}}|x-a|^{2}\mathop{}\!\mathrm{d}\big{(}\mathcal{L}(X_{s}^{\mu,a})-\Psi(\mu)(x)\big{)}\Big{|}\leq 2C_{\mu}(1+|a|^{2})^{2}e^{-\lambda_{\mu}s}.$
Thus
$\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}\mathop{}\!\mathrm{d}s=\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}.$
Then the continuity of $g_{a}(\cdot,r_{a}^{2})$ and (5.14) yield $g_{a}(\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2},r_{a}^{2})\leq 0$. Hence, by (5.4), $\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}\leq\tilde{r}_{a}^{2}$, i.e., $\Psi(\mu)\in\overline{B(\delta_{a},\tilde{r}_{a})}$. Hence,
$\Psi\big{(}\overline{B(\delta_{a},r_{a})}\big{)}\subset\overline{B(\delta_{a},\tilde{r}_{a})}\subset B(\delta_{a},r_{a}).$
The proof is completed. ∎
Remark 5.1.
Note that $\mu$ is an invariant measure of the equation (1.1) if and only if $\Psi(\mu)=\mu$, and then Theorem 5.5 indicates that there is no invariant measure on $\partial\overline{B(\delta_{a},r_{a})}:=\overline{B(\delta_{a},r_{a})}\setminus B(\delta_{a},r_{a})$.
5.3 Local attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
To investigate the local dynamics near the invariant measures of McKean-Vlasov SDEs, we establish a connecting orbit theorem and a structure theorem on local attractors of order-preserving semigroups.
Firstly, the following proposition concerns about connecting orbits on an order interval in $\mathcal{P}_{2}(\mathbb{R})$. It is a $\mathcal{P}_{2}$-version of the general connecting orbit theorem for an abstract monotone semiflow on a convex compact subset of a Hausdorff locally convex topological vector space (see Liu-Qu-Yao-Zhi [48, Theorem 4.2]).
Proposition 5.6.
Let $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ be an order-preserving semiflow, and suppose $\mu_{1}<_{\mathrm{st}}\mu_{2}$ are two order related invariant measures in $\mathcal{P}_{2}(\mathbb{R})$. Then at least one of the following holds:
  1. (a)
    $P_{t}^{*}$ has a third invariant measure $\nu\in\mathcal{P}_{2}(\mathbb{R})$ with $\mu_{1}<_{\mathrm{st}}\nu<_{\mathrm{st}}\mu_{2}$;
  2. (b)
    there exists an increasing connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\mu_{1}$ to $\mu_{2}$;
  3. (c)
    there exists a decreasing connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\mu_{2}$ to $\mu_{1}$.
Proof.
To prove the proposition, we only need to verify the assumptions (M1)-(M5) in [48, Theorem 4.2].
  1. (M1)
    Recall $(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$ is a normed space, see (5.10). A cone
    $C:=\bigg{\{}\mu\in\mathcal{M}_{1}(\mathbb{R}):\int_{\mathbb{R}^{d}}f\mathop{}\!\mathrm{d}\mu\geq 0\text{ for all nonnegative 1-Lipschitz increasing functions $f$}\bigg{\}}$
    in $\mathcal{M}_{1}(\mathbb{R})$ is given by [48, Lemma 3.5].
  2. (M2)
    By Proposition 3.4, we see that $S:=[\mu_{1},\mu_{2}]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$, and hence, $S$ is compact in $\mathcal{P}_{1}(\mathbb{R})$. Recall $\mathcal{P}_{1}(\mathbb{R})$ is closed convex in $\mathcal{M}_{1}(\mathbb{R})$, see (5.11). Then $S$ is compact in $\mathcal{M}_{1}(\mathbb{R})$. The convexity and non-singleton of $S$ are obvious. By Corollary 3.5, (5.9) and the dual representation of $\mathcal{W}_{1}$, the 2-Wasserstein metric on $[\mu,\nu]$ induces the relative topology in $\mathcal{M}_{1}(\mathbb{R})$.
  3. (M3)
    By [48, Lemma 3.5], the partial order induced by the cone $C$ in the normed space $\mathcal{M}_{1}(\mathbb{R})$ coincides with the stochastic order when restricted on $\mathcal{P}_{2}(\mathbb{R})$. The infimum and the supremum of $S$ in $\mathcal{M}_{1}(\mathbb{R})$ are clearly endpoints $\mu_{1}$, $\mu_{2}$, of the order interval.
  4. (M4)
    The semiflow $P_{t}^{*}$ is order-preserving.
  5. (M5)
    $\mu_{1}$ and $\mu_{2}$ are invariant measures of $P^{\ast}_{t}$.
Based on the existence of connecting orbits, we will present a theorem for order-preserving semigroups that ensures the local attraction of the forward limit of a connecting orbit.
Theorem 5.7.
(Local attractor). Let $P^{*}$ be an order-preserving semiflow on $\mathcal{P}_{2}(\mathbb{R})$. Assume that there exist a positively invariant bounded open set $B\subset\mathcal{P}_{2}(\mathbb{R})$ and an order-bounded closed set $K\subset\mathcal{P}_{2}(\mathbb{R})$ such that $K$ attracts $B$. Assume further there is only one invariant measure $\mu_{0}$ in $\overline{B}$. Then the following hold.
  1. (i)
    If $\mu_{0}\geq_{\mathrm{st}}\omega(B)$ and there exists another invariant measure $\mu_{-1}<_{\mathrm{st}}\mu_{0}$ with an increasing connecting orbit $\{\nu^{-1,0}_{t}\}_{t\in\mathbb{R}}$ from $\mu_{-1}$ to $\mu_{0}$ such that $\nu^{-1,0}_{t_{-1}}\leq_{\mathrm{st}}\omega(B)$ for some $t_{-1}\in\mathbb{R}$, then $\mu_{0}$ attracts $B$.
  2. (ii)
    If $\mu_{0}\leq_{\mathrm{st}}\omega(B)$ and there exists another invariant measure $\mu_{1}>_{\mathrm{st}}\mu_{0}$ with a decreasing connecting orbit $\{\nu^{1,0}_{t}\}_{t\in\mathbb{R}}$ from $\mu_{1}$ to $\mu_{0}$ such that $\nu^{1,0}_{t_{1}}\geq_{\mathrm{st}}\omega(B)$ for some $t_{1}\in\mathbb{R}$, then $\mu_{0}$ attracts $B$.
  3. (iii)
    If there exist another two invariant measures $\mu_{-1}<_{\mathrm{st}}\mu_{0}$ and $\mu_{1}>_{\mathrm{st}}\mu_{0}$ with an increasing connecting orbit $\{\nu^{-1,0}_{t}\}_{t\in\mathbb{R}}$ from $\mu_{-1}$ to $\mu_{0}$ and a decreasing connecting orbit $\{\nu^{1,0}_{t}\}_{t\in\mathbb{R}}$ from $\mu_{1}$ to $\mu_{0}$ such that, $\nu^{-1,0}_{t_{-1}}\leq_{\mathrm{st}}\omega(B)$ for some $t_{-1}\in\mathbb{R}$ and $\nu^{1,0}_{t_{1}}\geq_{\mathrm{st}}\omega(B)$ for some $t_{1}\in\mathbb{R}$, then $\mu_{0}$ attracts $B$.
Proof.
(i). By Proposition 3.4, $K$ is compact in $\mathcal{P}_{2}(\mathbb{R})$. Since the compact set $K$ attracts $B$, Lemma 4.1 (iii) entails that $\omega(B)$ is nonempty, compact, invariant. Considering the closed set
$S:=\{\nu^{-1,0}_{t}:\ t\geq t_{-1}\}\cup\{\mu_{0}\}\cup\omega(B)\subset\mathcal{P}_{2}(\mathbb{R}),$
it is obvious that $S$ is positively invariant. Since $B$ is positively invariant, we have $\omega(B)\subset\overline{B}$. Hence, $\mu_{0}$ is the unique invariant measure in $S$. Observing that $S$ has a lower bound $\nu^{-1,0}_{t_{-1}}\in S$ and an upper bound $\mu_{0}\in S$, Theorem 4.2 (i) (vi) and Remark 4.1 (i) entail that $\omega(S)=\{\mu_{0}\}$ is the global attractor of $P^{*}|_{S}$, where $P^{*}|_{S}$ means the semiflow $P^{*}$ restricted to $S$. The invariance of $\omega(B)$ and $\omega(B)\subset S$ entail that
$\omega(B)=\omega(\omega(B))\subset\omega(S).$
Thus, the nonemptiness of $\omega(B)$ implies that $\omega(B)=\{\mu_{0}\}$. Therefore, by Lemma 4.1 (iii), $\mu_{0}$ attracts $B$. This proves (i).
(ii). Considering the semiflow $P^{*}$ restricted to the positively invariant closed set
$S:=\{\nu^{1,0}_{t}:\ t\geq t_{1}\}\cup\{\mu_{0}\}\cup\omega(B)\subset\mathcal{P}_{2}(\mathbb{R})$
with lower bound $\mu_{0}\in S$ and upper bound $\nu^{1,0}_{t_{1}}\in S$, the proof of (ii) is analogous to (i).
(iii). Considering the semiflow $P^{*}$ restricted to the positively invairant closed set
$S:=\{\nu^{-1,0}_{t}:\ t\geq t_{-1}\}\cup\{\mu_{0}\}\cup\omega(B)\cup\{\nu^{1,0}_{t}:\ t\geq t_{1}\}\subset\mathcal{P}_{2}(\mathbb{R})$
with lower bound $\nu^{-1,0}_{t_{-1}}\in S$ and upper bound $\nu^{1,0}_{t_{1}}\in S$, the proof of (iii) is analogous to (i). ∎
5.4 Proof of Theorem 2.3 and Corollary 2.4
We first prove Theorem 2.3 (i)-(iii). Then we use the structure theorems of global attractors (Theorem 4.2) and local attractors (Theorem 5.7) to show local convergence and attracting unbounded open sets (Theorem 2.3 (iv)(v)). To this end, we prepare two crucial propositions by tracing solutions backward along connecting orbits (see Proposition 5.8-5.9).
Proof of (i)-(iii) in Theorem 2.3.
(ii). By Theorem 5.4, we know that the open ball $B(\delta_{a_{k}},r_{k})$ is positively invariant under $P_{t}^{*}$, and there is some invariant measure $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$. Next, we show that the open balls $\{B(\delta_{a_{k}},r_{k})\}_{k=1}^{n}$ are pairwise disjoint. Note that (2.1) implies
$r_{a_{i}}^{2}+r_{a_{j}}^{2}\leq 2^{-1}|a_{i}-a_{j}|^{2},\ \text{ for all }\ 1\leq i\neq j\leq n.$ (5.16)
Suppose on the contrary that there exists $\nu\in B(\delta_{a_{i}},r_{a_{i}})\cap B(\delta_{a_{j}},r_{a_{j}})$. Then,
$\mathcal{W}_{2}^{2}(\nu,\delta_{a_{i}})=\int_{\mathbb{R}}|x-a_{i}|^{2}\mathop{}\!\mathrm{d}\nu(x)\geq 2^{-1}|a_{i}-a_{j}|^{2}-\int_{\mathbb{R}}|x-a_{j}|^{2}\mathop{}\!\mathrm{d}\nu(x)=2^{-1}|a_{i}-a_{j}|^{2}-\mathcal{W}_{2}^{2}(\nu,\delta_{a_{j}}).$
Hence, $2^{-1}|a_{i}-a_{j}|^{2}\leq\mathcal{W}_{2}^{2}(\nu,\delta_{a_{i}})+\mathcal{W}_{2}^{2}(\nu,\delta_{a_{j}})<r_{a_{i}}^{2}+r_{a_{j}}^{2}$, which gives a contradiction to (5.16). This proves (ii) except (2.3), and (2.3) will be obtained later.
(i). For each $k=1,2,\dots,n-1$, notice $\frac{1}{2}(a_{k+1}-a_{k})^{2}\geq r_{k}^{2}+r_{k+1}^{2}\geq\frac{1}{2}(r_{k}+r_{k+1})^{2}$, so we have $r_{k}+r_{k+1}\leq a_{k+1}-a_{k}$. Together with $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$ and $\nu_{2k+1}\in B(\delta_{a_{k+1}},r_{k+1})$, we have
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k-1}(x)<a_{k}+r_{k}\leq a_{k+1}-r_{k+1}<\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k+1}(x).$
Then Lemma 4.10 entails that $\nu_{1}<_{\mathrm{st}}\nu_{3}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$. In the order interval $[\nu_{2k-1},\nu_{2k+1}]$, by Proposition 5.6, at least one of the three situations happens:
  • there is another invariant measure $\nu_{2k}$ with $\nu_{2k-1}<_{\mathrm{st}}\nu_{2k}<_{\mathrm{st}}\nu_{2k+1}$;
  • there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$;
  • there is a decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k-1}$.
We want to exclude the existence of connecting orbits, and we argue by contradiction. Assume there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$. Since $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$ and $\nu_{2k+1}\in B(\delta_{a_{k+1}},r_{k+1})$, we have $P_{t_{0}}^{*}\mu\in B(\delta_{a_{k+1}},r_{k+1})$ for some $\mu\in B(\delta_{a_{k}},r_{k})$ and some $t_{0}>0$. But Theorem 5.4 and disjointness in (ii) tell us $P_{t}^{*}\mu\in B(\delta_{a_{k}},r_{k})$ for all $t\geq 0$, and $B(\delta_{a_{k}},r_{k})\cap B(\delta_{a_{k+1}},r_{k+1})=\emptyset$. This is a contradiction, so there is no increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$. In the same manner, we can show there is no decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k-1}$. Hence, we obtain $(2n-1)$ order related invariant measures, say $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$. Proposition 5.3 shows that there are no other invariant measures.
(ii)-(2.3) We only need to show that $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$, and $\nu_{2k}\notin\overline{B(\delta_{a_{k+1}},r_{k+1})}$ can be proved in the same way. We first show that $\nu_{2k}\notin B(\delta_{a_{k}},r_{k})$. Assume otherwise, $\nu_{2k}\in B(\delta_{a_{k}},r_{k})$, and apply Proposition 5.6 to the order interval $[\nu_{2k},\nu_{2k+1}]$. Then at least one of the following statements holds:
  • there is another invariant measure $\mu$ with $\nu_{2k}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu_{2k+1}$;
  • there is an increasing connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$;
  • there is a decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k}$.
The statement (i) excludes the existence of $\mu$. Repeating the argument in the proof of (i), we can invoke Theorem 5.4 and disjointness in (ii) to show there are no connecting orbits, regardless increasing or decreasing. This is impossible, and thus, we must have $\nu_{2k}\notin B(\delta_{a_{k}},r_{k})$. By Remark 5.1, $\nu_{2k}\notin\partial\overline{B(\delta_{a_{k}},r_{k})}$, and hence, we have $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$.
(iii). For each $k=1,2,\dots,n-1$, we prove there is a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$, and the existence of an increasing connecting orbit from $\nu_{2k}$ to $\mu_{2k+1}$ can be shown similarly. In the order interval $[\nu_{2k-1},\nu_{2k}]$, we use again Proposition 5.6. Then at least one of the following three is true:
  • there is another invariant measure $\mu$ with $\nu_{2k-1}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu_{2k}$;
  • there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k}$;
  • there is a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$.
The existence of $\mu$ is excluded by the statement (i). The statement (ii) suggests that $\nu_{2k-1}$ has an positively invariant neighbourhood $B(\delta_{a_{k}},r_{k})$. By (ii)-(2.3), $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$. By repeating the proof in (i), an increasing connecting orbit does not exist. The only option left is a decreasing connecting orbit $\{\nu_{t}^{k,\downarrow}\}_{t\in\mathbb{R}}$ from $\nu_{2k}$ to $\nu_{2k-1}$. ∎
Before the proof of (iv)(v) in Theorem 2.3, we need to verify a core condition in Theorem 5.7 that, $\omega(B(\delta_{a_{k}},r_{k}))$ has upper/lower bounds in connecting orbits. In Proposition 5.8-5.9, we will fix the connecting orbit as the laws of the solution flow of (1.1), which gives a time-inhomogeneous SDE. Noting that the solution of this SDE can be traced backward along the connecting orbit, and by the comparison theorem, we can find a measure in the connecting orbit at a point of time sufficiently close to $\infty$- ∞, which is a lower bound or upper bound of $\omega(B(\delta_{a_{k}},r_{k}))$ with respect to stochastic order.
Proposition 5.8.
Suppose Assumption 1 holds. Assume that a bounded subset $B$ in $\mathcal{P}_{2}(\mathbb{R})$ is positively invariant under $P^{*}$ and there exists a connecting orbit $\{\tilde{\nu}_{t}\}_{t\in\mathbb{R}}$ from an invariant measure $\nu$ to another invariant measure $\pi$. If
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu(x)<\inf_{\mu\in B}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x),\ \ \Big{(}resp.\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu(x)>\sup_{\mu\in B}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x)\Big{)},$ (5.17)
then there exists $t_{0}\in\mathbb{R}$ such that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ ((((resp. $\tilde{\nu}_{t}\geq_{\mathrm{st}}\omega(B)$)))) for all $t\leq t_{0}$.
Proof.
We first prove the “$<$” case in (5.17), and the case “$>$” can be proved similarly. Now consider the following time-inhomogeneous SDE
$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\tilde{\nu}_{t})(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (5.18)
Let $X_{t}^{s,\mu}$ be the unique solution of equation (5.18) at time $t$ with starting time $s$ and initial distribution $\mu$ and $P^{*}(t,s)$ be the dual semigroup of equation (5.18), i.e., $P^{*}(t,s)\mu:=\mathcal{L}(X_{t}^{s,\mu})$. Since $\{\tilde{\nu}_{t}\}_{t\in\mathbb{R}}$ is a connecting orbit, then Lemma 4.9 (i) yields that it is an entrance measure of equation (5.18), i.e., $P^{*}(t,s)\tilde{\nu}_{s}=\tilde{\nu}_{t}$ for all $t\geq s$. Note that $\sup_{t\in\mathbb{R}}\|\tilde{\nu}_{t}\|_{2}<\infty$ (by definition, the connecting orbit is bounded in $\mathcal{P}_{2}(\mathbb{R})$), by Lemma 4.9, we know that there exist $C>0,\lambda>0$ such that for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$ and $t\geq s$,
$d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big{)}\leq C(1+\|\mu\|_{2}^{2})e^{-\lambda(t-s)},$ (5.19)
where $d_{2}$ is given as in (5.15). Note that
$\Big{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\mathcal{L}(X_{t}^{s,\mu})(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)\Big{|}\leq|f|_{\infty}d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big{)},\ \text{ for all }\ f\in C_{b}(\mathbb{R}),$ (5.20)
and
$\Big{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\mathcal{L}(X_{t}^{s,\mu})(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)\Big{|}\leq d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big{)}.$ (5.21)
By (5.19) and (5.20), we conclude that $\mathcal{L}(X_{t}^{s,\mu_{s}})\to\tilde{\nu}_{t}$ weakly as $s\to-\infty$ for all bounded sets $\{\mu_{s}\}_{s\leq t}$ in $\mathcal{P}_{2}(\mathbb{R})$. Then (5.19), (5.21) together with [68, Theorem 6.9] entail that, for all $t\in\mathbb{R}$,
$\lim_{s\to-\infty}\mathcal{W}_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu_{s}}),\tilde{\nu}_{t}\big{)}=0,\ \text{ for any bounded set $\{\mu_{s}\}_{s\leq t}$ in $\mathcal{P}_{2}(\mathbb{R})$}.$ (5.22)
Note that $\{\tilde{\nu}_{t}\}_{t\in\mathbb{R}}$ is a connecting orbit from $\nu$ to $\pi$ in $\mathcal{P}_{2}(\mathbb{R})$, then
$\limsup_{t\to-\infty}\Big{|}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)-\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu(x)\Big{|}\leq\limsup_{t\to-\infty}\mathcal{W}_{1}(\tilde{\nu}_{t},\nu)\leq\lim_{t\to-\infty}\mathcal{W}_{2}(\tilde{\nu}_{t},\nu)=0.$
Hence, by (5.17), there exists $t_{0}\in\mathbb{R}$ such that
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)<\inf_{\mu\in B}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x),\ \text{ for all }\ t\leq t_{0}.$ (5.23)
Next, we prove that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ for all $t\leq t_{0}$. By Theorem 2.1 (ii), the order interval $[\underline{\nu},\overline{\nu}]$ attracts the bounded set $B$. Then Proposition 3.4 and Lemma 4.1 (iii) tell us $\omega(B)$ is nonempty. For any $\rho\in\omega(B)$, it follows from Lemma 4.1 (i) that there exist $\{(t_{k},\rho_{k})\}_{k\geq 1}\subset\mathbb{R}^{+}\times B$ such that $t_{k}\to\infty$ as $k\to\infty$ and
$\lim_{k\to\infty}\mathcal{W}_{2}(P^{*}_{t_{k}}\rho_{k},\rho)=0.$ (5.24)
Fix $t\leq t_{0}$, and we consider the following two SDEs starting from $t-t_{k}$:
$\mathop{}\!\mathrm{d}X_{s}$ $=-V^{\prime}(X_{s})\mathop{}\!\mathrm{d}s-(W^{\prime}*\nu_{s})(X_{s})\mathop{}\!\mathrm{d}s+\sigma(X_{s})\mathop{}\!\mathrm{d}B_{s},$ $\mathcal{L}(X_{t-t_{k}})$ $=\rho_{k},$
$\mathop{}\!\mathrm{d}Y_{s}$ $=-V^{\prime}(Y_{s})\mathop{}\!\mathrm{d}t-(W^{\prime}*P^{*}_{s-t+t_{k}}\rho_{k})(Y_{s})\mathop{}\!\mathrm{d}s+\sigma(Y_{s})\mathop{}\!\mathrm{d}B_{s},$ $\mathcal{L}(Y_{t-t_{k}})$ $=\rho_{k}.$
Note that $\rho_{k}\in B$ and $B$ is positively invariant, then Assumption 1 (iv), (5.23) and Theorem 4.3 yield
$\mathcal{L}(X_{t}^{t-t_{k},\rho_{k}})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{t-t_{k},\rho_{k}})=P^{*}_{t_{k}}\rho_{k}.$
On the other hand, (5.22) and boundedness of $B$ imply $\mathcal{W}_{2}(\mathcal{L}(X_{t}^{t-t_{k},\rho_{k}}),\tilde{\nu}_{t})\to 0$ as $k\to\infty$. Thus, (5.24) and the closedness of the stochastic order give $\tilde{\nu}_{t}\leq_{\mathrm{st}}\rho$. Hence, $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ as $\rho\in\omega(B)$ is arbitrary. ∎
Proposition 5.9.
Assume all assumptions in Theorem 2.3 hold. Let $\{\nu^{k,\uparrow}_{t}\}_{t\in\mathbb{R}}$ ((((resp. $\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$)))) be the increasing ((((resp. decreasing)))) connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$ ((((resp. $\nu_{2k-1}$)))) in Theorem 2.3 (iii). Then there exists $t_{k}\in\mathbb{R}$ such that, for all $t\leq t_{k}$,
$\nu^{k,\uparrow}_{t}\leq_{\mathrm{st}}\omega\big{(}B(\delta_{a_{k+1}},r_{k+1})\big{)},\ \text{ {$\big{(}$}resp. $\nu^{k,\downarrow}_{t}\geq_{\mathrm{st}}\omega\big{(}B(\delta_{a_{k}},r_{k})\big{)}${$\big{)}$}}.$
Proof.
We only prove there is some $t_{k}\in\mathbb{R}$ such that $\nu^{k,\uparrow}_{t}\leq_{\mathrm{st}}\omega\big{(}B(\delta_{a_{k+1}},r_{k+1})\big{)}$ for all $t\leq t_{k}$, and the proof of the other case is similar. Note that $B(\delta_{a_{k+1}},r_{k+1})$ is positively invariant under $P^{*}$, by virtue of Proposition 5.8, it suffices to show that
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k}(x)<\inf\Big{\{}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x):\mu\in B(\delta_{a_{k+1}},r_{k+1})\Big{\}}.$
Let $\{\mu_{m}\}_{m\in\mathbb{R}}$ be probability measures given in (1.4). According to (4.25), a measure $\mu$ is an invariant measure of (1.1) if and only if $\mu=\mu_{m}$ and $\int x\mathop{}\!\mathrm{d}\mu_{m}(x)=m$ for some $m\in\mathbb{R}$. By (i) of Theorem 2.3, there exist $m_{1}<m_{2}<\cdots<m_{2n-1}$ such that
$\nu_{i}=\mu_{m_{i}},\ \text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{i}(x)=m_{i},\ \text{ for all }\ i=1,2,\cdots,2n-1.$ (5.25)
We have the following claim.
Claim: For all $k=1,2,\cdots,n-1$, $a_{k}+r_{k}<m_{2k}<a_{k+1}-r_{k+1}$.
Proof of Claim: We only prove $m_{2k}<a_{k+1}-r_{k+1}$, and the other one can be obtained similarly.
We first show that $m_{2k}\notin[a_{k+1}-r_{k+1},a_{k+1}+r_{k+1}]$. Otherwise, since
$\bigg{\{}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x):\mu\in\overline{B(\delta_{a_{k+1}},r_{k+1})}\bigg{\}}=[a_{k+1}-r_{k+1},a_{k+1}+r_{k+1}],$
there exists $\mu\in\overline{B(\delta_{a_{k+1}},r_{k+1})}$ such that $\int x\mathop{}\!\mathrm{d}\mu(x)=m_{2k}$. By the definition of $\Psi$ (see (5.13)) and Theorem 5.5, it follows that $\Psi(\mu)=\mu_{m_{2k}}=\nu_{2k}\in\Psi(\overline{B(\delta_{a_{k+1}},r_{k+1})})\subset B(\delta_{a_{k+1}},r_{k+1})$. This gives a contradiction to (2.3). Hence $m_{2k}\notin[a_{k+1}-r_{k+1},a_{k+1}+r_{k+1}]$.
Note also, by Theorem 2.3 (i)(ii) that $\nu_{2k}<_{\mathrm{st}}\nu_{2k+1}$ and $\nu_{2k+1}\in B(\delta_{a_{k+1}},r_{k+1})$, we conclude that
$m_{2k}=\int x\mathop{}\!\mathrm{d}\nu_{2k}(x)\leq\int x\mathop{}\!\mathrm{d}\nu_{2k+1}(x)\leq a_{k+1}+r_{k+1}.$
This together with $m_{2k}\notin[a_{k+1}-r_{k+1},a_{k+1}+r_{k+1}]$ yields $m_{2k}<a_{k+1}-r_{k+1}$. The Claim is proved.
By Claim, it follows
$m_{2k}=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k}(x)<\inf\Big{\{}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x):\mu\in\overline{B(\delta_{a_{k+1}},r_{k+1})}\Big{\}}=a_{k+1}-r_{k+1}.$
This finishes the proof. ∎
Proof of (iv)(v) in Theorem 2.3.
(iv). For $k=1$, according to the statement (ii), the open ball $B(\delta_{a_{1}},r_{1})$ is positively invariant under $P^{*}$. By Theorem 2.3 (i)(ii) and Remark 5.1, there is only one invariant measure $\nu_{1}$ in $\overline{B(\delta_{a_{1}},r_{1})}$. Theorem 2.1 (ii) and Theorem 2.3 (i) entail that $[\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{1}},r_{1})$ and hence $\omega(B(\delta_{a_{1}},r_{1}))\subset[\nu_{1},\nu_{2n-1}]$. Then $\nu_{1}\leq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$. Let $\{\nu_{t}^{1,\downarrow}\}_{t\in\mathbb{R}}$ be a decreasing connecting orbit from $\nu_{2}$ to $\nu_{1}$ given in Theorem 2.3 (iii). Proposition 5.9 shows that there exists $t_{1}\in\mathbb{R}$ such that $\nu_{t_{1}}^{1,\downarrow}\geq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$. Then Theorem 5.7 (ii) implies that $\nu_{1}$ attracts $B(\delta_{a_{1}},r_{1})$ under $P^{*}$, which means $\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$ as $t\to\infty$.
For $k=2,3,\dots,n-1$, let $\{\nu_{t}^{k,\downarrow}\}_{t\in\mathbb{R}}$ be a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$ and $\{\nu_{t}^{k,\uparrow}\}_{t\in\mathbb{R}}$ be an increasing connecting orbit from $\nu_{2k-2}$ to $\nu_{2k-1}$ given in Theorem 2.3 (iii). It follows from the statement (ii) that $B(\delta_{a_{k}},r_{k})$ is positively invariant under $P^{*}$. By Theorem 2.3 (i)(ii) and Remark 5.1, there is only one invariant measure $\nu_{2k-1}$ in $\overline{B(\delta_{a_{k}},r_{k})}$. We also see the order interval $[\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{k}},r_{k})$ by Theorem 2.1 (ii) and Theorem 2.3 (i). By Proposition 5.9, we can find some $t_{k}\in\mathbb{R}$ such that $\nu_{t_{k}}^{k,\uparrow}\leq_{\mathrm{st}}\omega(B(\delta_{a_{k}},r_{k}))\leq_{\mathrm{st}}\nu_{t_{k}}^{k,\downarrow}$. Theorem 5.7 (iii) suggests that $\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$.
The last case is $k=n$. We just note that $\nu_{2n-1}\geq_{\mathrm{st}}\omega(B(\delta_{a_{n}},r_{n}))$ by Theorem 2.1 (ii) and Theorem 2.3 (i). Then Theorem 5.7 (i) shows the result valid as in the proof of case $k=1$.
(v). Noticing the conditions in Theorem 4.2 are verified in the proof of Theorem 2.1, we apply Theorem 4.2 (iv) to finish the proof. If $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r_{1})$, then from Theorem 4.2 (iv), we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$. If $\mu\geq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{n}},r_{n})$, the convergence result follows as well.
Now, it remains to show the basin of attraction of $\nu_{1}$ (resp. $\nu_{2n-1}$) contains $\cup_{a\leq 0}B(\delta_{a_{1}+a},r_{1})$ (resp. $\cup_{a\geq 0}B(\delta_{a_{n}+a},r_{n})$). We only need to show for any $\mu\in\cup_{a\leq 0}B(\delta_{a_{1}+a},r_{1})$, there exists $\nu\in B(\delta_{a_{1}},r_{1})$ such that $\mu\leq_{\mathrm{st}}\nu$. And the other case of $\nu_{2n-1}$ can be proved similarly. For $a\in\mathbb{R}$, let $\tau^{a}$ be the translation map by $a$ on $\mathbb{R}$, say $\tau^{a}(x):=x+a$. Let $\tau_{\#}^{a}$ be the pushforward of $\tau^{a}$ acting on probability measures, that is, $\mathop{}\!\mathrm{d}(\tau_{\#}^{a}\nu)(x):=\mathop{}\!\mathrm{d}\nu(x-a)$. Then it follows $\tau_{\#}^{a}\nu\leq_{\mathrm{st}}\nu$ for $a\leq 0$, and $\tau_{\#}^{a}\nu\geq_{\mathrm{st}}\nu$ for $a\geq 0$. Note $\tau_{\#}^{a}(B(\delta_{a_{1}},r_{1}))=B(\delta_{a_{1}+a},r_{1})$, and hence, for any $\mu\in B(\delta_{a_{1}+a},r_{1})$ with $a\leq 0$, there exists $\nu\in B(\delta_{a_{1}},r_{1})$ such that $\mu=\tau_{\#}^{a}\nu\leq_{\mathrm{st}}\nu$. ∎
Proof of Corollary 2.4.
(i). This follows from (5.25).
(ii). Note from (4.24) that for any $m\in\mathbb{R}$, $\mu_{m}$ is the unique invariant measure of the following SDE:
$\mathop{}\!\mathrm{d}X^{m}_{t}=\big{(}-V^{\prime}(X^{m}_{t})-\theta X^{m}_{t}+\theta m\big{)}\mathop{}\!\mathrm{d}t+\sigma(X^{m}_{t})\mathop{}\!\mathrm{d}B_{t}.$ (5.26)
We prove (ii) in the following three cases.
Case $k=1$: Let $\{\nu^{1,\downarrow}_{t}\}_{t\in\mathbb{R}}$ be the decreasing connecting orbit from $\nu_{2}=\mu_{m_{2}}$ to $\nu_{1}$ in Theorem 2.3 (iii). Then $\mathcal{W}_{2}(\nu_{t}^{1,\downarrow},\mu_{m_{2}})\to 0$ as $t\to-\infty$. Fixing any $m\in(-\infty,m_{2})$, we conclude that there exists $t_{0}\in\mathbb{R}$ such that
$\alpha^{1,\downarrow}_{t}:=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu^{1,\downarrow}_{t}(x)\geq m,\ \text{ for all }\ t\leq t_{0}.$
Similar to (5.18), consider the following time-inhomogeneous SDE:
$\mathop{}\!\mathrm{d}Y_{t}=\big{(}-V^{\prime}(Y_{t})-\theta Y_{t}+\theta\alpha^{1,\downarrow}_{t}\big{)}\mathop{}\!\mathrm{d}t+\sigma(Y_{t})\mathop{}\!\mathrm{d}B_{t}.$ (5.27)
Then it is easy to check that $\{\nu^{1,\downarrow}_{t}\}_{t\in\mathbb{R}}$ is an entrance measure of (5.27). Let $X_{t}^{m;s,0}$ and $Y^{s,0}_{t}$ be the unique solution at time $t$ start from $s\leq t$ with initial data 0 of equation (5.26) and (5.27) respectively. Then Lemma 4.9 entails that there exist $C>0,\lambda>0$ such that for all $t\geq s$,
$d_{2}\big{(}\mathcal{L}(X_{t}^{m;s,0}),\mu_{m}\big{)}\leq Ce^{-\lambda(t-s)},\ \text{ and }\ d_{2}\big{(}\mathcal{L}(Y^{s,0}_{t}),\nu_{t}^{1,\downarrow}\big{)}\leq Ce^{-\lambda(t-s)}.$
Similar to the proof of (5.22), we conclude that
$\lim_{s\to-\infty}\mathcal{W}_{2}\big{(}\mathcal{L}(X_{t_{0}}^{m;s,0}),\mu_{m}\big{)}=0,\ \text{ and }\ \lim_{s\to-\infty}\mathcal{W}_{2}\big{(}\mathcal{L}(Y^{s,0}_{t_{0}}),\nu^{1,\downarrow}_{t_{0}}\big{)}=0.$ (5.28)
On the other hand, Theorem 4.3 yields
$\mathcal{L}(X_{t_{0}}^{m;s,0})\leq_{\mathrm{st}}\mathcal{L}(Y^{s,0}_{t_{0}}),\ \text{ for all }\ s\leq t_{0}.$
Then the closedness of the stochastic order gives $\mu_{m}\leq_{\mathrm{st}}\nu^{1,\downarrow}_{t_{0}}$. Note that
$\mathcal{W}_{2}(P_{t}^{*}\nu^{1,\downarrow}_{t_{0}},\nu_{1})=\mathcal{W}_{2}(\nu^{1,\downarrow}_{t_{0}+t},\nu_{1})\to 0,\ \text{ as }\ t\to\infty,$
then Theorem 4.2 (v) shows that $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{1})\to 0$ as $t\to\infty$.
Case $1<k<n$: Let $\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$ and $\{\nu^{k-1,\uparrow}_{t}\}_{t\in\mathbb{R}}$ be the corresponding decreasing and increasing connecting orbits from $\nu_{2k}=\mu_{m_{2k}}$ to $\nu_{2k-1}$, and $\nu_{2k-2}=\mu_{m_{2k-2}}$ to $\nu_{2k-1}$ respectively in Theorem 2.3 (iii). Fix any $m\in(m_{2k-2},m_{2k})$, the connecting orbits indicate that there exists $t_{0}\in\mathbb{R}$ such that
$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu^{k,\downarrow}_{t}(x)\geq m,\text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu^{k-1,\uparrow}_{t}(x)\leq m\ \text{ for all }\ t\leq t_{0}.$
Similar to that in Case $k=1$, we have $\nu^{k-1,\uparrow}_{t_{0}}\leq_{\mathrm{st}}\mu_{m}\leq_{\mathrm{st}}\nu^{k,\downarrow}_{t_{0}}$. Note that
$\mathcal{W}_{2}(P^{*}_{t}\nu^{k-1,\uparrow}_{t_{0}},\nu_{2k-1})\to 0,\ \text{ and }\ \mathcal{W}_{2}(P^{*}_{t}\nu^{k,\downarrow}_{t_{0}},\nu_{2k-1})\to 0,\ \text{ as }\ t\to\infty,$
and the order-preserving property of $P^{*}$ entails that
$P^{*}_{t}\nu^{k-1,\uparrow}_{t_{0}}=\nu^{k-1,\uparrow}_{t_{0}+t}\leq_{\mathrm{st}}P^{*}_{t}\mu_{m}\leq_{\mathrm{st}}\nu^{k,\downarrow}_{t_{0}+t}=P^{*}_{t}\nu^{k,\downarrow}_{t_{0}},\ \text{ for all }\ t\geq 0.$
Thus, the relative compactness of $\{P_{t}^{*}\mu_{m}\}_{t\geq 0}$, the closedness and antisymmetry of the stochastic order entail that $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{2k-1})\to 0$ as $t\to\infty$.
Case $k=n$: Let $\{\nu^{n-1,\uparrow}_{t}\}_{t\in\mathbb{R}}$ be the increasing connecting orbit from $\nu_{2n-2}=\mu_{m_{2n-2}}$ to $\nu_{2n-1}$ and fix any $m\in(m_{2n-2},\infty)$. Similar to Case $k=1$, there exists $t_{0}\in\mathbb{R}$ such that $\mu_{m}\geq_{\mathrm{st}}\nu^{n-1,\uparrow}_{t_{0}}$. Then applying Theorem 4.2 (v) again, we conclude $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{2n-1})\to 0$ as $t\to\infty$. ∎
6 Applications to Classical Examples: Proof of Theorem 2.5-2.9
Proof of Theorem 2.5.
According to Theorem 2.3, it suffices to show that equation (2.4) are locally dissipative at $-1,1$ with configurations $(r,g_{-1}),(r,g_{1})$ for $r=\frac{9-\sqrt{17}}{8}$ and for some $g_{\pm 1}$.
Note that
$V^{\prime}(x)=x^{3}-x,\ W(x)=\frac{\theta}{2}x^{2}.$
Then we construct $g_{\pm 1}$ as follows.
Construction of $g_{1}$: For any $x\in\mathbb{R},\mu\in\mathcal{P}_{2}(\mathbb{R})$, by a simple calculation, we have
$-2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big{(}2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}\big{)},$
where
$\overline{\sigma}:=\sup_{x\in\mathbb{R}}|\sigma(x)|.$
Then we choose $g_{1}$ as follows: for any $z,w\geq 0$,
$g_{1}(z,w)=2z^{2}-6z^{\frac{3}{2}}+(4+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-|\overline{\sigma}|^{2}.$ (6.1)
Obviously, $g_{1}$ is continuous and $g_{1}(z,\cdot)$ is decreasing for any $z\geq 0$ and hence
$\inf_{0\leq w\leq r^{2}}g_{1}(z,w)=g_{1}(z,r^{2}),\ \text{ for any }\ r>0.$
Note that
$\frac{\partial g_{1}(z,w)}{\partial z}=4z-9z^{\frac{1}{2}}+4+2\theta-\theta w^{\frac{1}{2}}z^{-\frac{1}{2}},$
and
$\frac{\partial^{2}g_{1}(z,w)}{\partial z^{2}}=4-\frac{9}{2}z^{-\frac{1}{2}}+\frac{1}{2}\theta w^{\frac{1}{2}}z^{-\frac{3}{2}}.$
Then by calculation, we know that for any $w>0$, $g_{1}(\cdot,w)$ is convex if and only if
$\theta\geq\frac{27}{16w^{\frac{1}{2}}}.$ (6.2)
Note that for any $10<r<$ 1,
$g_{1}(r^{2},r^{2})=2r^{2}(r-1)(r-2)-\overline{\sigma}^{2},$
and
$\frac{\partial g_{1}}{\partial z}(r^{2},r^{2})=4r^{2}-9r+4+\theta>\theta-1.$
Now for any fixed $10<r<$ 1, if we choose
$\theta\geq\frac{27}{16r}>1,\ \text{ and }\ \overline{\sigma}^{2}<2r^{2}(r-1)(r-2),$
then
$g_{1}(r^{2},r^{2})>0,\ \text{ and }\ \frac{\partial g_{1}}{\partial z}(r^{2},r^{2})>0.$
By (6.2), we know that $g_{1}(\cdot,r^{2})$ is convex. Hence,
$\frac{\partial g_{1}}{\partial z}(z,r^{2})\geq\frac{\partial g_{1}}{\partial z}(r^{2},r^{2})>0\ \text{ for all }\ z\geq r^{2},$
and thus
$g_{1}(z,r^{2})\geq g_{1}(r^{2},r^{2})>0\ \text{ for all }\ z\geq r^{2}.$
Hence, for any $10<r<$ 1 and choosing $g_{1}$ as in (6.1), equation (2.4) is locally dissipative at $1$ with configuration $(r,g_{1})$ whenever
$\theta\geq\frac{27}{16r},\ \text{ and }\ \overline{\sigma}^{2}<2r^{2}(r-1)(r-2).$ (6.3)
Note that
$\max_{0<r<1}2r^{2}(r-1)(r-2)=2r^{2}(r-1)(r-2)\Big{|}_{r=\frac{9-\sqrt{17}}{8}}=\frac{51\sqrt{17}-107}{256}.$ (6.4)
Combining (6.3)(6.4), under the condition (2.5) in Theorem 2.5, equation (2.4) is locally dissipative at $1$ with configuration $(r,g_{1})$ where $r=\frac{9-\sqrt{17}}{8}$ and $g_{1}$ is given as in (6.1).
Construction of $g_{-1}$: Similarly, let $g_{-1}=g_{1}$ be given as in (6.1), we can show that equation (2.4) is also locally dissipative at $-1$ with configuration $\big{(}\frac{9-\sqrt{17}}{8},g_{-1}\big{)}$. ∎
Proof of Theorem 2.6.
By a simple calculation, we know that
$-2xV^{\prime}(x+1)\leq-2x^{2}(x+1)(x+2)+\frac{1}{4}.$
Then we have
$-2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big{(}2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}-\frac{1}{4}\big{)}.$
Choose $g_{1}$ as follows: for any $z,w\geq 0$,
$g_{1}(z,w)=2z^{2}-6z^{\frac{3}{2}}+(4+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-|\overline{\sigma}|^{2}-\frac{1}{4}.$ (6.5)
Repeating the proof of that in Theorem 2.5, we know that the equation (2.6) is locally dissipative at $1$ with configuration $(r,g_{1})$ with $r=\frac{9-\sqrt{17}}{8}$ whenever
$\theta\geq\frac{27(9+\sqrt{17})}{128},\ \text{ and }\ \overline{\sigma}^{2}+\frac{1}{4}<\frac{51\sqrt{17}-107}{256},$
i.e., (2.7) holds. Similarly, choose $g_{-1}=g_{1}$ as in (6.5), the equation (2.6) is locally dissipative at $-1$ with configuration $\big{(}\frac{9-\sqrt{17}}{8},g_{-1}\big{)}$. Then the desired results follow from Theorem 2.3 directly. ∎
Proof of Theorem 2.7.
Take
$g_{1}(z,w)$ $=2z^{2}-8z^{\frac{3}{2}}+(6+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{1}$ $=\frac{6-3\sqrt{3}}{2},$
$g_{-2}(z,w)$ $=2z^{2}-10z^{\frac{3}{2}}+(12+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{-2}$ $=\frac{(\sqrt{33}-15)^{2}}{64}.$
Through calculations similar in Theorem 2.5, local dissipative conditions hold at $1,-2$ for the equation (2.8). We apply Theorem 2.3 to get the desired results. ∎
Proof of Theorem 2.8.
Take
$g_{0}(z,w)$ $=2z^{3}-10z^{2}+(8+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{0}$ $=\frac{5-\sqrt{13}}{3},$
$g_{2}(z,w)$ $=2z^{3}-20z^{\frac{5}{2}}+70z^{2}-100z^{\frac{3}{2}}+(48+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{2}$ $=r_{0},$
$g_{-2}(z,w)$ $=g_{2}(z,w),$ $r_{-2}$ $=r_{0}.$
Then the equation (2.9) is locally dissipative at $0,\pm 2$. The proof is finished by applying Theorem 2.3. ∎
Proof of Theorem 2.9.
The conclusions directly follow from Theorem 2.1. ∎
Acknowledgement
We acknowledge the financial supports of EPSRC grant ref (ref. EP/S005293/2), Royal Society through the Award of Newton International Fellowship (ref. NIF/R1/221003) and National Natural Science Foundation of China (No. 12171280).
Appendix A Nonexistence of Strong Ordered Pairs for Stochastic Order
In addition to closedness of the stochastic order mentioned in Section 3.1, almost all classical results in the theory of systems that preserve a certain partial order (monotone dynamical systems) are established on the fundamental assumption that the state space $X$ is not only endowed with a partial order relation $\leq$”, but is also a “strongly ordered space” (see e.g., Hirsch [41, 42]). A strongly ordered space implies that the interior of the partial order relation $\leq$” (referred to as the strong ordering $\ll$”) is a nonempty relation on $X$, and its closure is $\leq$”. The strong ordering $\ll$” is a relation for which two points “$x\ll y$” means that there are open neighbourhoods $U,V$ in $X$ of $x,y$ respectively such that $U\leq V$, where the term “$U\leq V$” refers to $u\leq v$ for any $u\in U$ and $v\in V$. If $x\ll y$, we call $(x,y)$ is a strong ordered pair. Based on the tool of order topology, especially the topology generated by the interiors of order intervals enclosed by strong ordered pairs, the theory of monotone dynamical systems and its applications to cooperative irreducible ODEs (coordinate order in $\mathbb{R}^{d}$), delay equations and parabolic equations (pointwise order in functional spaces) have undergone extensively investigation (see e.g., [60, 43, 71]).
Unfortunately, in the appendix, we will show that the strong ordering relation generated by the stochastic order “$\leq_{\mathrm{st}}$” in $\mathcal{P}_{2}(\mathbb{R})$ is empty, meaning that NO strong ordered pairs exist. We actually prove a stronger conclusion – for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$, every open set $B\subset\mathcal{P}_{2}(\mathbb{R})$ has an element that is order unrelated with $\mu$ (see Proposition A.1). This implies that classical theory of monotone dynamical systems are inapplicable to order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$. Consequently, exploration of properties of the stochastic order and establishment of the theory for systems preserving the stochastic order, as discussed in Section 3.2, 4.1 and 5.3, are indispensable.
Proposition A.1.
Let $\mu\in\mathcal{P}_{2}(\mathbb{R})$ and $U$ be an open set in $\mathcal{P}_{2}(\mathbb{R})$. Then there exists $\nu\in U$ such that $\mu,\nu$ are order unrelated. In particular, any order interval in $\mathcal{P}_{2}(\mathbb{R})$ has no interior point.
Proof.
Since $U$ is open, there exist $\nu_{0}\in U$ and $\delta>0$ such that
$B(\nu_{0},2\delta):=\{\nu\in\mathcal{P}_{2}(\mathbb{R}):\mathcal{W}_{2}(\nu,\nu_{0})<2\delta\}\subset U.$
Claim. There exist $c>0$ and $\nu_{1}\in B(\nu_{0},\delta)$ such that
$\nu_{1}((-\infty,-r])>0,\ \ \nu_{1}([r,\infty))>0,\ \text{ for all }\ r\geq c.$ (A.1)
Proof of Claim: Note that there exists $c>0$ such that $\nu_{0}((-c,c))>0$. Let
$\rho(x)=\frac{e^{-|x|}}{\int_{\mathbb{R}}|x|^{2}e^{-|x|}\mathop{}\!\mathrm{d}x},\ \ \rho_{\epsilon}(x):=\epsilon^{-1}\rho(\epsilon^{-1}x),\ \epsilon\in(0,1).$
For any $\epsilon\in(0,1)$, we have
$\begin{split}\alpha_{\epsilon}&:=\int_{(-\infty,-c]\cup[c,\infty)}\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x\\ &\leq\frac{1}{c^{2}}\int_{(-\infty,-c]\cup[c,\infty)}|x|^{2}\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x\\ &\leq\frac{1}{c^{2}}\int_{\mathbb{R}}|x|^{2}\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x\\ &=\frac{\epsilon^{2}}{c^{2}}\int_{\mathbb{R}}|x|^{2}\rho(x)\mathop{}\!\mathrm{d}x\\ &=\frac{\epsilon^{2}}{c^{2}}.\end{split}$
Choose $\epsilon_{0}\in(0,1)$ such that $\epsilon_{0}^{2}<c^{2}\nu_{0}((-c,c))$ and we define, for any $0<\epsilon<\epsilon_{0}$,
$\mathop{}\!\mathrm{d}\nu_{1,\epsilon}(x)=\begin{cases}\frac{\nu_{0}((-c,c))-\alpha_{\epsilon}}{\nu_{0}((-c,c))}\mathop{}\!\mathrm{d}\nu_{0}(x),&\text{on }(-c,c),\\ \mathop{}\!\mathrm{d}\nu_{0}(x)+\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x,&\text{on }(-\infty,-c]\cup[c,\infty).\end{cases}$
Since $\nu_{0}\in\mathcal{P}_{2}(\mathbb{R})$, it is easy to check that $\nu_{1,\epsilon}\in\mathcal{P}_{2}(\mathbb{R})$ satisfying (A.1) for all $0<\epsilon<\epsilon_{0}$. Note that for any $f\in C_{b}(\mathbb{R})$, where $C_{b}(\mathbb{R})$ is the set of all bounded continuous functions from $\mathbb{R}$ to $\mathbb{R}$,
$\begin{split}&\ \ \ \ \bigg{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{1,\epsilon}(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{0}(x)\bigg{|}\\ &\leq\frac{\alpha_{\epsilon}}{\nu_{0}((-c,c))}\int_{(-c,c)}|f(x)|\mathop{}\!\mathrm{d}\nu_{0}(x)+\int_{(-\infty,-c]\cup[c,\infty)}|f(x)|\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x\\ &\leq 2|f|_{\infty}\alpha_{\epsilon}\\ &\leq\frac{2|f|_{\infty}}{c^{2}}\epsilon^{2}\to 0\ \text{ as }\ \epsilon\to 0,\end{split}$ (A.2)
[…]
$\begin{split}&\ \ \ \ \lim_{\epsilon\to 0}\bigg{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1,\epsilon}(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{0}(x)\bigg{|}\\ &\leq\lim_{\epsilon\to 0}\Big{(}c^{2}\alpha_{\epsilon}+\int_{(-\infty,-c]\cup[c,\infty)}|x|^{2}\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x\Big{)}\\ &\leq\lim_{\epsilon\to 0}2\epsilon^{2}=0.\end{split}$ (A.3)
[…]
If $\mu$ and $\nu_{1}$ are order unrelated, the proof is completed by choosing $\nu=\nu_{1}$. Otherwise, without loss of generality, we assume $\mu\leq_{\mathrm{st}}\nu_{1}$. Then we know that
$\mu((-\infty,-r])\geq\nu_{1}((-\infty,-r]),\ \ \mu([r,\infty))\leq\nu_{1}([r,\infty)),\ \text{ for all }\ r\geq c.$ (A.4)
We will finish the proof by considering the following two cases.
Case 1. $\mu([r_{0},\infty))<\nu_{1}([r_{0},\infty))$ for some $r_{0}\geq c$. By (A.1) and (A.4), one has
$\mu((-\infty,-r])\geq\nu_{1}((-\infty,-r])>0,\ \text{ for all }\ r\geq c,$ (A.5)
Note that $\lim_{r\to\infty}\mu((-\infty,-r])=0$, then for any positive integer $n\geq c$, there exists $r_{n}>n+1$ such that
$0<2\mu((-\infty,-r_{n}])<\nu_{1}((-\infty,-n]).$
For any $n\geq c$, define
$\mathop{}\!\mathrm{d}\nu_{2,n}(x)=\begin{cases}\mathop{}\!\mathrm{d}\nu_{1}(x),&\text{on }(-n,\infty),\\ \big{(}\nu_{1}((-\infty,-n])-2\mu((-\infty,-r_{n}])\big{)}\mathop{}\!\mathrm{d}x,&\text{on }(-(n+1),-n],\\ 0,&\text{on }(-r_{n},-(n+1)],\\ 2\mathop{}\!\mathrm{d}\mu(x),&\text{on }(-\infty,-r_{n}].\end{cases}$ (A.6)
[…]
$\begin{split}&\ \ \ \ \bigg{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{2,n}(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{1}(x)\bigg{|}\\ &\leq\int_{(-\infty,-n]}|f(x)|\mathop{}\!\mathrm{d}\nu_{2,n}(x)+\int_{(-\infty,-n]}|f(x)|\mathop{}\!\mathrm{d}\nu_{1}(x)\\ &\leq 2|f|_{\infty}\nu_{1}((-\infty,-n])\to 0\ \text{ as }\ n\to\infty.\end{split}$
Note also that $\mu,\nu_{1}\in\mathcal{P}_{2}(\mathbb{R})$, then for any $n\geq c$,
$\begin{split}&\ \ \ \ \bigg{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{2,n}(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1}(x)\bigg{|}\\ &\leq\int_{(-\infty,-n]}|x|^{2}\mathop{}\!\mathrm{d}\nu_{2,n}(x)+\int_{(-\infty,-n]}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1}(x)\\ &\leq(n+1)^{2}\nu_{1}((-\infty,-n])+2\int_{(-\infty,-r_{n}]}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)+\int_{(-\infty,-n]}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1}(x)\\ &\leq 2\int_{(-\infty,-r_{n}]}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)+5\int_{(-\infty,-n]}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1}(x)\to 0\ \text{ as }\ n\to\infty.\end{split}$
Hence, it follows from Theorem 6.9 in [68] again that $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{2,n},\nu_{1})=0.$ Thus, there exists $n_{0}\geq c$ such that $\nu:=\nu_{2,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$. By (A.5), the construction of $\nu_{2,n_{0}}$ in (A.6) and the choice of $r_{0}$ in Case 1, we know that
$\begin{split}0<\mu&((-\infty,-r])<2\mu((-\infty,-r])=\nu((-\infty,-r]),\text{ for all }\ r\geq r_{n_{0}},\\ &\mu([r_{0},\infty))<\nu_{1}([r_{0},\infty))=\nu([r_{0},\infty)).\end{split}$
This, together with Lemma 3.1, indicates that $\mu,\nu$ are order unrelated.
Case 2. $\mu([r,\infty))=\nu_{1}([r,\infty))$ for all $r\geq c$. By (A.1), one has
$\mu([r,\infty))=\nu_{1}([r,\infty))>0\ \text{ for all }\ r\geq c.$ (A.7)
Note that
$\lim_{r\to\infty}\mu((-\infty,-r])=\lim_{r\to\infty}\mu([r,\infty))=0,$
then it follows from (A.5) and (A.7) that, for any positive integer $n\geq c$, there exists $r_{n}>n+1$ such that
$0<2\mu((-\infty,-r_{n}])<\nu_{1}((-\infty,-n]),\ \text{ and }\ 0<2\mu([r_{n},\infty))<\nu_{1}([n,\infty)).$
For any $n\geq c$, define
$\mathop{}\!\mathrm{d}\nu_{3,n}(x)=\begin{cases}\mathop{}\!\mathrm{d}\nu_{1}(x),&\text{on }(-n,n),\\ \big{(}\nu_{1}([n,\infty))-2\mu([r_{n},\infty))\big{)}\mathop{}\!\mathrm{d}x,&\text{on }[n,n+1),\\ \big{(}\nu_{1}((-\infty,-n])-2\mu((-\infty,-r_{n}])\big{)}\mathop{}\!\mathrm{d}x,&\text{on }(-(n+1),-n],\\ 0,&\text{on }(-r_{n},-(n+1)]\cup[n+1,r_{n}),\\ 2\mathop{}\!\mathrm{d}\mu(x),&\text{on }(-\infty,-r_{n}]\cup[r_{n},\infty).\end{cases}$
Similarly, we can conclude that $\nu_{3,n}\in\mathcal{P}_{2}(\mathbb{R})$ and $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{3,n},\nu_{1})=0.$ Choose $n_{0}\geq c$ large enough such that $\nu:=\nu_{3,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$, then we know that for all $r\geq r_{n_{0}}$,
$\begin{split}0<&\mu((-\infty,-r])<2\mu((-\infty,-r])=\nu((-\infty,-r]),\ \text{ and }\\ 0<&\mu([r,\infty))<2\mu([r,\infty))=\nu([r,\infty)).\end{split}$
Hence, Lemma 3.1 gives $\mu,\nu$ are order unrelated. ∎

### v2

@@
-Global Dynamics of McKean-Vlasov SDEs via Stochastic Order
+Global Dynamics of Granular Media Equations via Stochastic Order
 Abstract
-This paper studies the rich dynamics of general one-dimensional McKean-Vlasov stochastic differential equations based on the preservation of stochastic order.
-The existing work is limited to the case of additive noises, while our framework allows multiplicative noises.
-When the equation has multiple invariant measures, we show basins of attraction of certain invariant measures contain unbounded open sets in the $2$-Wasserstein space, which is vacant in previous research even for additive noises.
-Also, our main results target on total orderedness and finiteness of invariant measures, global convergence to order interval enclosed by two order-related invariant measures, alternating arrangement of invariant measures in terms of stability (locally attracting) and instability (connecting orbits).
-Our theorems cover a wide range of classical granular media equations, such as symmetric, flat-bottom and asymmetric double-well confinement potentials with quadratic interaction, multi-well landscapes, and double-well landscapes with perturbations.
+This paper studies the rich dynamics of one-dimensional granular media equations with attractive quadratic interactions.
+Building on the monotone dynamical systems framework developed in an earlier work, we allow for multiplicative noise, in contrast to most existing results restricted to additive noise.
+Within this framework, we show that, in the one-dimensional setting, invariant measures are totally ordered with respect to the stochastic order.
+The basins of attraction of the minimal and maximal invariant measures contain unbounded open sets in the $2$-Wasserstein space, which is vacant in previous research even for additive noises.
+Also, our main results address the global convergence to the order interval enclosed by the minimal and maximal invariant measures, and an alternating arrangement of invariant measures in terms of stability (locally attracting) and instability (as the backward limit of a connecting orbit).
+Our theorems cover a wide range of classical granular media equations, such as double-well and multi-well landscapes.
 Specific values for the parameter ranges, explicit descriptions of attracting sets and phase diagrams are provided.
 MSC2020 subject classifications: Primary 60H10, 60B10; secondary 37C65, 60E15
-Keywords: McKean-Vlasov SDEs, order-preserving semigroups, invariant measures, connecting orbits, basins of attraction, stochastic order
+Keywords: Granular media equations, long-time behavior, invariant measures, basins of attraction, order-preserving semiflow, stochastic order
 1 Introduction
-This paper focuses on one-dimensional McKean-Vlasov processes governed by the following stochastic differential equation (SDE),
-$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mathcal{L}(X_{t}))(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},$ (1.1)
-where $\{B_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Brownian motion on $\mathbb{R}$, $\mathcal{L}(X_{t})$ is the law of the random variable $X_{t}$, $V:\mathbb{R}\to\mathbb{R}$ is a potential function, $W:\mathbb{R}\to\mathbb{R}$ is a local interaction, and $\sigma:\mathbb{R}\to\mathbb{R}$ is a diffusion term.
 The category of McKean-Vlasov SDEs refers to a class of SDEs whose coefficients depend on laws of solutions.
 McKean [38] introduced it as a generalization of Itô’s diffusion processes by considering a nonlinear version of Fokker-Planck equations.
-For (1.1), the corresponding Fokker-Planck partial differential equation (PDE) is given by
+Among McKean–Vlasov SDEs, one of the most important and widely studied classes with concrete applications is given by granular media models:
+$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mathcal{L}(X_{t}))(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},$ (1.1)
+where $\{B_{t}\}_{t\in\mathbb{R}}$ is a standard two-sided Brownian motion on $\mathbb{R}$, $\mathcal{L}(X_{t})$ is the law of the random variable $X_{t}$, $V:\mathbb{R}\to\mathbb{R}$ is a confining potential function, $W:\mathbb{R}\to\mathbb{R}$ is an interaction potential, $\sigma:\mathbb{R}\to\mathbb{R}$ is a diffusion term and
+$(W^{\prime}*\mu)(x):=\int_{\mathbb{R}}W^{\prime}(x-y)\mathop{}\!\mathrm{d}\mu(y),\ \text{ for all probability measure $\mu$ on $\mathbb{R}$.}$
+The law $\rho_{t}=\mathcal{L}(X_{t})$ of the solution to (1.1) formally solves the associated nonlinear Fokker-Planck equation
 $\partial_{t}\rho_{t}=\frac{1}{2}\partial_{xx}\big(\sigma^{2}\rho_{t}\big)+\partial_{x}\big[(V^{\prime}+W^{\prime}*\rho_{t})\rho_{t}\big],$ (1.2)
-where $\rho_{t}$ is the probability density function of $X_{t}$.
-After that, the aspect of particle approximation property was examined for McKean-Vlasov SDEs.
-The equation (1.1) can be deduced as the mean-filed limit (or hydrodynamic limit) of the following interacting particle system,
+which describes the evolution of the particle density.
+From another point of view, The equation (1.1) can be deduced as the mean-field limit (or hydrodynamic limit) of the following interacting particle system,
 $\mathop{}\!\mathrm{d}X_{t}^{i,N}=-V^{\prime}\big(X_{t}^{i,N}\big)\mathop{}\!\mathrm{d}t-\frac{1}{N}\sum_{j=1}^{N}W^{\prime}\big(X_{t}^{i,N}-X_{t}^{j,N}\big)\mathop{}\!\mathrm{d}t+\sigma\big(X_{t}^{i,N}\big)\mathop{}\!\mathrm{d}B_{t}^{i},\quad i=1,2,\dots,N,$ (1.3)
@@
 This is called “propagation of chaos”, and we refer to [10, 11, 39, 48] on this topic.
-Due to the origin from infinite-particle limits, McKean-Vlasov SDEs are natural mathematical models for large-scale phenomena.
-See [9] for galactic dynamics, [10] for plasma physics, and [59] for molecular and fluid dynamics.
-[11, 13, 45] use this kind of SDEs to model collective behaviour of different species.
-Another application on mathematical biology is [26, 28, 35, 49] to neuroscience.
-For optimal controls and differential games of McKean-Vlasov type, as well as study on machine learning and neural networks, see [15, 24, 70].
-The equation (1.1) is a model for the time evolution of granular materials (see Benedetto-Caglioti-Pulvirenti [8]), so it is also called a granular media equation.
-A typical setting for the local interaction is attractive quadratic, namely
-$W(x)=\frac{\theta}{2}x^{2},\quad\theta>0.$
-As for the potential function, it has more choices ranging from double-well to multi-well shapes.
-Depending on the noise intensity and the number of critical points of the potential, the number of invariant measures of the equation (1.1) may vary.
-We refer to Dawson [27], Tugaut [66] for phase transitions on the number of invariant measures in double-well landscapes, and Alecio [1] for multi-well cases.
-Roughly speaking, if $V^{\prime}$ has $(2n-1)$ simple zeros, the McKean-Vlasov SDE (1.1) admits a unique invariant measure when $\sigma>\overline{\sigma}_{c}$ above some upper-critical value, and it admits $(2n-1)$ invariant measures when $\sigma<\underline{\sigma}_{c}$ below a lower-critical value.
-For more related work on phase transitions, see also [16, 29, 30].
-Besides phase transitions on the number of invariant measures, convergence to invariant measures has been an active research direction.
-The study for the dynamics of (1.1) can be divided into two different cases according to uniqueness and non-uniqueness of invariant measures.
-If there is a unique invariant measure for (1.1), since the interacting particle system (1.3) also has a unique invariant measure, uniform-in-time propagation of chaos can give the convergence result.
-Malrieu [51] goes through this way for uniformly convex interaction terms, and obtains exponential convergence rate via a log-Sobolev inequality.
-Cattiaux-Guillin-Malrieu [20] relaxes to the non-uniformly convex case, and instead uses the method of synchronous couplings.
-We refer to [32, 46] and literature therein for more on uniform propagation of chaos, and [38] for uniform log-Sobolev inequalities of mean-filed particle systems.
-Investigation of the associated PDE (1.2) is an alternative path to prove convergence to the unique invariant measure for (1.1).
-From this perspective, [5, 6] shows convergence to the unique invariant measure with vanishing potential ($V=0$) by Bakry-Émery’s $\Gamma$-calculus, while Benedetto-Caglioti-Carrillo-Pulvirenti [7] utilizes free energy as a Lyapunov functional when $V$ is convex.
-Actually, when $\sigma$ is a constant, the PDE (1.2) is a Wasserstein gradient flow of the free energy (see Ambrosio-Gigli-Savare [3]), which is a sum of entropy, potential energy and interaction energy,
-$\mathcal{F}(\rho):=\frac{\sigma^{2}}{2}\int_{\mathbb{R}}\rho(x)\log\rho(x)\mathop{}\!\mathrm{d}x+\int_{\mathbb{R}}V(x)\rho(x)\mathop{}\!\mathrm{d}x+\frac{1}{2}\int_{\mathbb{R}\times\mathbb{R}}W(x-y)\rho(x)\rho(y)\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y.$
-By the “Riemannian structure” on Wasserstein spaces, also known as Otto’s calculus (see Otto [58]), free energy attains its global minimum at the unique invariant measure.
-Then nonlinear log-Sobolev inequalities, that link free energy and its dissipation, guarantee the exponential convergence.
-Carrillo-McCann-Villani [18] achieves this by checking Bakry-Émery criterion, and Carrillo-McCann-Villani [19] develops an abstract theory of gradient flows in length spaces to handle degenerate convexity.
-For other relevant work, see [4, 31] and refenrence therein for Bakry-Émery analysis, [23, 50] for log-Sobolev inequalties, and [17, 64] for entropy dissipation.
-In the case of non-unique invariant measures, the dynamics is not as intuitive as the case of unique invariant measures.
-When the equation (1.1) has three invariant measures with a double-well potential, one of the predictions is local convergence.
-Now, propagation of chaos is hard to resort to, as the particle system (1.3) still has a unique invariant measure.
-Nonetheless, the gradient flow (1.2) remains workable, where free energy has two local minima and one local maximum, attained at three invariant measures respectively.
-Assuming qualitative smallness of the noise intensity $\sigma$, Tugaut [67] uses WJ-inequality to show the local convergence.
-See Bolley-Gentil-Guillin [12] for the derivation of WJ-inequalities and consequent spectral gaps in Wasserstein distance.
-Meanwhile, Monmarché-Reygner [56] executes a local version of log-Sobolev argument to obtain the local convergence in a sharp range of $\sigma$ for double-well landscapes.
-We refer to Tamura [62, 63] for early works using PDE techniques, and Cormier [25] by Lions derivatives.
-As for other directions than local convergence for the double-well case, the convergence to one of invariant measures is carried out by Tugaut [65], and on the other hand, Liu-Qu-Yao-Zhi [48] indicates that the central invariant measure is unstable and has connecting orbits to the other two.
-The present paper is a continuation of [48], which establishes a general framework to study unstable dynamics in multi-dimensional cooperative McKean-Vlasov SDEs.
-In this paper, we investigate the richer dynamical behaviour of the one-dimensional equation (1.1) with non-unique invariant measures, from the stochastic-order-preserving perspective.
-For general case, Theorem 2.1 and Corollary 2.2 show the set of invariant measures is totally ordered and finite, and its infimum and supremum enclose a $\mathcal{P}_{2}$-compact order interval $[\underline{\nu},\overline{\nu}]$, which attracts any bounded set in the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$.
-The infimum $\underline{\nu}$ attracts all probability measures smaller than $\underline{\nu}$, and the supremum $\overline{\nu}$ attracts those greater than $\overline{\nu}$.
-Given the zero-crossing number of the derivative of the potential function, Theorem 2.3 and Corollary 2.4 unveil an alternating arrangement of invariant measures in terms of stability and instability, with stable ones $\underline{\nu}$, $\overline{\nu}$ at the beginning and the end.
-A stable invariant measure attracts an open neighbourhood, and an unstable invariant measure means there exists a connecting orbit away from it.
-We can further prove the infimum $\underline{\nu}$ and the supremum $\overline{\nu}$ attract unbounded open neighbourhoods in $\mathcal{P}_{2}(\mathbb{R})$.
-Additionally, Theorem 2.5-2.9 provide some applications of the main results to classical examples, and their phase diagrams are presented in Figure 1, 2.
-Our approach is based on order-preservation of the semigroup generated by the equation (1.1) on the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$.
-To be precise, take an attractive quadratic interaction, and let $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ be the semigroup of (1.1).
-Consider the stochastic order “$\leq_{\mathrm{st}}$” defined by
-$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\ \ \text{for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
-Then $P_{t}^{*}$ is order-preserving on $\mathcal{P}_{2}(\mathbb{R})$ (see Theorem 4.4), that is
-$\mu\leq_{\mathrm{st}}\nu\ \text{ implies }\ P_{t}^{*}\mu\leq_{\mathrm{st}}P_{t}^{*}\nu\ \text{ for $t\geq 0$}.$
-Generally speaking, our method is parallel to the gradient flow approach.
-When the invariant measure is unique, to show global convergence, the gradient flow approach relies on the decreasing free energy along each single orbit, and a functional inequality involving free energy and its dissipation.
-By contrast, we develop the global attractor structures of order-preserving semigroups based on the characterization of order-boundedness and the coming down from infinity property.
-For the case of multiple invariant measures, the gradient flow approach needs a local functional inequality to show local convergence, while we establish the local attractor structures by tracing solutions backward along connecting orbits.
-To the best of our knowledge, the gradient flow approach concerning the basin of attraction of a single invariant measure is limited to the case of additive noises and local convergence, when invariant measures are non-unique.
-The main advantage of the order-preserving method is that we step further to give unbounded basins of attraction, even for multiplicative noises.
-To illustrate that the basins of attraction of $\underline{\nu}$ (infimum invariant measure), $\overline{\nu}$ (supremum invariant measure) in Theorem 2.1 are unbounded, we exhibit inside a class of measures with explicit densities.
-Precisely, there are initial conditions arbitrarily far from $\underline{\nu}$ (resp.
-$\overline{\nu}$), for which the solutions converge to $\underline{\nu}$ (resp.
-$\overline{\nu}$).
-Consider a family of probability measures $\{\mu_{m}\}_{m\in\mathbb{R}}$ on $\mathbb{R}$, whose densities are given by
-$\frac{\mathop{}\!\mathrm{d}\mu_{m}}{\mathop{}\!\mathrm{d}x}\propto\frac{1}{\sigma^{2}(x)}\exp\bigg{\{}-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg{\}}.$ (1.4)
-From a calculation of Fokker-Planck equations, the invariant measures of the equation (1.1) are those $\mu_{m}$ satisfying the following self-consistency equation,
-$\int_{\mathbb{R}}(x-m)\mathop{}\!\mathrm{d}\mu_{m}(x)=0.$
-Since $\mu_{m_{1}}\leq_{\mathrm{st}}\mu_{m_{2}}$ for $m_{1}\leq m_{2}$ (see Remark 4.3), it means $\underline{\nu}=\mu_{\underline{m}}$ and $\overline{\nu}=\mu_{\overline{m}}$ for some $\underline{m}\leq\overline{m}$.
-In this case, the solution starting from $\mu_{m}$ converges to $\underline{\nu}$ if $m\leq\underline{m}$, and converges to $\overline{\nu}$ if $m\geq\overline{m}$.
-Corollary 2.4 also gives a complete partition on limiting behaviour of solutions starting from the totally ordered arc $\{\mu_{m}\}_{m\in\mathbb{R}}$ in (1.4).
-Our method can even show the basins of attraction contain unbounded open sets in $\mathcal{P}_{2}(\mathbb{R})$.
-By translating probability measures, we obtain order related pairs under the stochastic order.
-Given a local neighbourhood attracted by the infimum invariant measure, its basin of attraction actually contains all translated copies of the neighbourhood along the descending direction of the stochastic order.
-Correspondingly, the basin of attraction of the supremum invariant measure contains every ascending copy of its attracting neighbourhood.
-It is worth mentioning the stochastic order is inconsistent with the Wasserstein topology.
-In fact, any order interval has empty interior in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition A.1), and resembles a “thin slice” as shown in Figure 1, 2.
+Due to their origin as infinite-particle limits, equations (1.1)-(1.2) provide natural mathematical models for large-scale phenomena.
+They are widely applied in various fields, such as battery models [21], optimization [6] and machine learning [40].
+In this paper, we focus on the typical attractive quadratic interaction potential, namely
+$W(x)=\frac{\theta}{2}x^{2},\quad\theta>0,$
+which arises naturally in a variety of application contexts, including muscle contraction [33], chemical kinetics [29], statistical physics [23], and large economic systems [3].
+A salient feature of this class of equations is the occurrence of multiple invariant measures if the noise is not too strong.
+For the double-well landscapes with additive noise, Dawson [13] and Tugaut [52] obtain phase transitions in the number of invariant measures.
+For the multi-well case, we refer to Alecio [1] and the references therein.
+Roughly speaking, this phenomenon can be understood through the geometry of the confining potential.
+It is common to consider well-shaped confining potentials $V$ in granular media equations, which means $V(x)\to\infty$ as $x\to\pm\infty$.
+If its derivative $V^{\prime}$ only has simple zeros, then basically $V^{\prime}$ has odd (2n1)(2n-1) zeros.
+In such a setting, the equation has exactly (2n1)(2n-1) invariant measures when noise is small enough, and has a unique invariant measure when noise is large.
+Consequently, the study of the dynamics of (1.1) is often divided into two regimes: the case of a unique invariant measure and the case of multiple invariant measures.
+In the former, convergence to the unique invariant measure has been extensively studied.
+For instance, Carrillo-McCann-Villani [7] employed generalized logarithmic Sobolev inequalities and mass transportation inequalities, while Cattiaux-Guillin-Malrieu [9] used a uniform-in-time propagation of chaos method to demonstrate convergence to the unique invariant measure.
+For more on this topic, see [8, 16, 22, 56] and the references therein.
+In the presence of multiple invariant measures, a complete characterization of the basins of attraction is in general a highly challenging problem.
+Current results in this direction are mainly restricted to establishing the local convergence around certain stable invariant measures, showing that solutions starting from initial distributions sufficiently close to them converge to the corresponding invariant measures.
+In Tugaut [53], local convergence is established under the assumptions that the initial condition admits a finite higher-order moment, a $C^{\infty}$-smooth density, and finite entropy.
+Monmarché-Reygner [43] applied a local version of the log-Sobolev inequality to obtain local convergence within a sharp range of the noise intensity $\sigma$ for double-well landscapes.
+Zhang [58] achieved local convergence by linearizing the nonlinear Markov semigroup, applicable to double-well landscapes.
+For early works using the free energy function, see Tamura [49, 50], while approaches relying on Lions derivatives can be found in Cormier [12].
+From a global perspective, Tugaut [51] showed that, for double-well granular media equations with qualitatively sufficiently small noise, solutions starting from initial conditions with finite higher-order moment, a $C^{\infty}$-smooth density, and finite entropy will converge to one of the three invariant measures.
+In [36], a monotone dynamical systems framework for cooperative McKean–Vlasov SDEs was developed.
+More precisely, under a local dissipativity assumption, the existence of multiple order related invariant measures in the Wasserstein space was established, together with monotone connecting (heteroclinic) orbits between them with respect to the stochastic order.
+Building on the main result of [36], the present paper goes further in the one-dimensional setting.
+By analyzing the local dissipativity condition and the number of invariant measures, we identify connecting orbits for a broader class of one-dimensional granular media equations with multiplicative noise, including, in particular, polynomial-type confining potentials of even degree.
+Relying on the existence of such connecting orbits together with a structural theorem for local attractors of order-preserving semigroups, we prove local convergence in the $2$-Wasserstein space.
+Furthermore, by combining a structural result on global attractors with an analysis of the stochastic order, we show that the basins of attraction of the minimal and maximal invariant measures contain unbounded open sets (see our main results Theorem 1.2-Corollary 1.5).
+To illustrate the applicability of our abstract results, we present below a theorem for the case where $V^{\prime}$ has finitely many simple zeros.
+Theorem 1.1.
+Suppose Assumption 1 holds, and $V^{\prime}$ has finite zeros and they are all simple.
+Then there exist $\theta_{0}>0,\ \sigma_{0}>0$ such that for all $\theta\geq\theta_{0}$ and $\overline{\sigma}\leq\sigma_{0}$,
+  1. (i)
+    (Total Orderedness) there are exactly (2n1)(2n-1) invariant measures $\{\nu_{k}\}_{k=1}^{2n-1}$ with $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$ for some integer $n\geq 1$;
+  2. (ii)
+    (Connecting Orbits) for all $k=1,2,\dots,n-1$, there are a decreasing connecting orbit $\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$ from $\nu_{2k}$ to $\nu_{2k-1}$ and an increasing connecting orbit $\{\nu^{k,\uparrow}_{t}\}_{t\in\mathbb{R}}$ from $\nu_{2k}$ to $\nu_{2k+1}$;
+  3. (iii)
+    (Locally Attracting) there exist $r>0$ and $a_{1}<\cdots<a_{k}$ such that for all $1\leq k\leq n$, $\nu_{2k-1}\in B(\delta_{a_{k}},r)$ and
+    $\mathcal{W}_{2}(P_{t}^{*}B(\delta_{a_{k}},r),\nu_{2k-1})=\sup_{\mu\in B(\delta_{a_{k}},r)}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0\text{ as }t\to\infty;$
+  4. (iv)
+    (Attracting Unbounded Open Neighbourhoods) if $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r)$, then
+    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\text{ as }t\to\infty.$
+    If $\mu\geq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{n}},r)$, then
+    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0\text{ as }t\to\infty.$
+    In particular,
+    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\text{ as }t\to\infty\ \text{ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{a_{1}+a},r)$},$
+    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0\text{ as }t\to\infty\ \text{ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{a_{n}+a},r)$}.$
+Our theorem clearly includes confining potentials of polynomial type, in particular those with
+$V^{\prime}(x)=a\prod_{k=1}^{2n-1}(x-\alpha_{k}),\qquad a>0,\ \alpha_{k}\in\mathbb{R}\ \text{and}\ \alpha_{i}\neq\alpha_{j}\ (i\neq j).$
+We stress that our approach is expected to extend to more general interaction potentials, for instance those satisfying $W^{\prime\prime}>0$, as well as to higher-dimensional systems, including a multi-species population models; see [14, 15].
+Our approach begins with the order-preserving property of the law-evolution semigroup generated by equation (1.1) on the $2$-Wasserstein space $\mathcal{P}_{2}(\mathbb{R})$ (see e.g., [30, 36, 45]).
+More precisely, we consider the stochastic order $\leq_{\mathrm{st}}$ defined by
+[…]
+Under a hyper-dissipativity assumption, we develop a dynamical result describing the global attractor structure of order-preserving semigroups (Proposition 3.2), based on a characterization of order-boundedness (Proposition 2.2) and the coming down from infinity property (Proposition 3.6).
+This yields a partial description of basins of attraction in the sense of the stochastic order (Theorem 1.2, Corollary 1.3).
+Building upon this global attractor analysis, and under the additional assumptions on local dissipativity and zero-crossing number, we further derive local attractor structures by tracing solutions backward along connecting orbits (Section 4.2).
+As a consequence, we obtain local convergence in the topological sense, as well as the existence of unbounded open subsets contained in the corresponding basins of attraction (Theorem 1.4, Corollary 1.5).
+It is worth noting that this phenomenon is far from obvious.
+Indeed, any order interval has empty interior in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition A.1), and resembles a ‘thin slice” as shown in Figure 1, 2.
 So attracting open sets by the method of preserving such an unsatisfactory partial order relation is counter-intuitive.
-As another necessary ingredient in our approach, we give the maximal number of invariant measures (see Proposition 5.3).
-Due to the analyticity and convolution form of the self-consistency function, we show the number of invariant measures of the equation (1.1) is finite and at most the zero-crossing number of the potential derivative $V^{\prime}$, regardless of the ranges of $\theta$ and $\sigma$.
-This helps to make our main theorems quantitative, and overcomes difficulties caused by qualitative results and missing parameter in the previous literature on phase transitions (see e.g., [1, 66]).
-One the other hand, this proposition allows $V^{\prime}$ vanishes on some interval, which is out of the settings in [67].
-Also, it provides an insight that the essential factor in the maximal number of invariant measures is the zero-crossing number of $V^{\prime}$, instead of the number of simple zeros of $V^{\prime}$.
-To demonstrate the versatility of our theory, we apply our main theorems to the following examples, including symmetric, flat-bottom, asymmetric double-well potentials, multi-well potentials, and double-well with perturbations (see Theorem 2.5-2.9).
-Take $W(x)=\frac{\theta}{2}x^{2}$ with $\theta>0$, and take the potential derivatives respectively as
-[…]
-[…]
-[…]
-[…]
-[…]
-In all these equations, all the dynamical results from Theorem 2.1, Corollary 2.2, Theorem 2.3 and Corollary 2.4 hold.
-The specific values for the ranges of $\theta$, $\sigma$, and explicit descriptions of attracting sets are given.
-For the phase diagrams, see Figure 1, 2.
+Finally, our results suggest the presence of a saddle-point structure in the double-well granular media equation.
+This leads us to formulate, at the end of the paper, a conjecture on the existence of an unordered invariant separatrix containing the unstable invariant measure, which separates the basins of attraction of another two stable invariant measures.
 This paper is organized as follows.
-In Section 2, we state our main results.
+In Section 1.1, we state the main results beyond Theorem 1.1.
 Section 2 is about properties of the stochastic order, including an equivalent characterization of order boundedness in $\mathcal{P}_{2}(\mathbb{R})$ and a generalization of the lattice property.
 Section 3 is devoted to the proof of Theorem 1.2 and Corollary 1.3.
-We obtain the structure of global attractors of order-preserving semigroups, the comparison theorem for extrinsic-law-dependent SDEs, and the coming down from infinity property of McKean-Vlasov SDEs.
-Section 5 serves the proof of Theorem 2.3 and Corollary 2.4.
-We deduce the information on the number of invariant measures, the existence of positively invariant neighbourhoods of invariant measures, and the structure of local attractors of order-preserving semigroups.
-Also, the order bounds of limit sets are found by tracing solutions backward along connecting orbits.
-In Section 6, we apply the main theorems to examples stated in Theorem 2.5-2.9.
+Section 4 serves the proof of the remaining theorems and the formulation of a saddle-point conjecture in the double-well granular media equation.
 In Appendix A, we show that, in $\mathcal{P}_{2}(\mathbb{R})$, any probability measure has no order relation with any open set, which denies the existence of strong ordered pairs for the stochastic order.
 Notations
-[…]
-[type=symbols,style=mylong]
-2 Assumptions and Main Results
+1.1 Assumptions and Main Results
 In this section, we state our assumptions, some necessary definitions and main results.
 Assumption 1.
-(i) The continuous function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is one-sided Lipschitz, i.e., there exists $L>0$ such that, for all $x,y\in\mathbb{R}$,
-$-(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq L\left\lvert x-y\right\rvert^{2}.$
+(i) The function $V\in C^{2}(\mathbb{R};\mathbb{R})$ such that the continuous function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is one-sided Lipschitz, i.e., there exists $L>0$ such that, for all $x,y\in\mathbb{R}$,
+$(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq L\left\lvert x-y\right\rvert^{2}.$
 (ii) The function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ has polynomial growth, i.e., there exist $L>0$, $\kappa\geq 1$ such that, for all $x\in\mathbb{R}$,
@@
 (iii) The function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is hyper-dissipative, i.e., there exist $\alpha>0$, $\beta>0$, $\delta>0$ such that, for all $x,y\in\mathbb{R}$,
-$-(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq-\alpha\left\lvert x-y\right\rvert^{2+\delta}+\beta.$
+$(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq-\alpha\left\lvert x-y\right\rvert^{2+\delta}+\beta.$
 (iv) $W(x)=\frac{\theta}{2}x^{2}$ with $\theta>0$.
@@
 Definition 1.1.
-The stochastic order “$\leq_{\mathrm{st}}$” on $\mathcal{P}_{2}(\mathbb{R})$ is a partial order, defined by
-$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\text{ for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
-Definition 2.2.
-The zero-crossing number of a continuous function $f:\mathbb{R}\to\mathbb{R}$, denoted by $Z(f)$, is the maximal integer $N$ such that, there exists an increasing sequence $x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}$ satisfying $f(x_{n})f(x_{n+1})<0$ for all $n=1,2,\dots,N$.
-Definition 2.3.
-The equation (1.1) is called locally dissipative at $a\in\mathbb{R}$ with configuration $(r,g)$, if
-  1. (i)
-    the function $g:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}$ satisfies, for all $x\in\mathbb{R}$, $\mu\in\mathcal{P}_{2}(\mathbb{R})$,
-    $-2xV^{\prime}(x+a)-2x(W^{\prime}*\mu)(x)+\sigma^{2}(x+a)\leq-g(x^{2},\|\mu\|_{2}^{2});$
-  2. (ii)
-    the number $r>0$ and the function $g$ satisfy
-    $\begin{split}&g(\cdot,r^{2})\text{ is continuous and convex};\\ &\inf_{0\leq w\leq r^{2}}g(z,w)=g(z,r^{2});\\ &\inf_{z\geq r^{2}}g(z,r^{2})>0.\end{split}$
+The stochastic order $\leq_{\mathrm{st}}$” on $\mathcal{P}_{2}(\mathbb{R})$ is a partial order, defined by
+[…]
 Definition 1.2.
@@
     $P^{\ast}_{t}\nu_{s}=\nu_{t+s}\text{ for any $t\geq 0$, $s\in\mathbb{R}$};$
-    $\mathcal{W}_{2}(\nu_{t},\nu)\to 0\text{ as $t\to-\infty$ and }\mathcal{W}_{2}(\nu_{t},\tilde{\nu})\to 0\text{ as $t\to\infty$}.$
+[…]
     Furthermore, if
@@
 Now we state our main theorems.
-Theorem 2.1, 2.3 and Corollary 2.2, 2.4 are abstract results for the equation (1.1), while Theorem 2.5-2.9 deal with concrete equations.
-In the sequel, we denote $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by the semigroup generated by the equation (1.1), see also (4.5).
+Theorem 1.2, 1.4 and Corollary 1.3, 1.5 are abstract results for the equation (1.1), while Theorem 1.6-1.8 deal with concrete equations.
+In the sequel, we denote by $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ the semigroup associated to the equation (1.1):
+$P_{t}^{*}\mu:=\mathcal{L}(X_{t}^{\xi})\ \text{ with }\ \mathcal{L}(\xi)=\mu,$ (1.4)
+where $X_{t}^{\xi}$ is the unique solution of (1.1) with initial data $\xi$ at time 0.
 Theorem 1.2.
@@
 (Global Convergence) Under Assumption 1, if the equation (1.1) has a unique invariant measure $\nu$, then for all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,\nu)\to 0$ as $t\to\infty$.
+Definition 1.3.
+The zero-crossing number of a continuous function $f:\mathbb{R}\to\mathbb{R}$, denoted by $Z(f)$, is the maximal integer $N$ such that, there exists an increasing sequence $x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}$ satisfying $f(x_{n})f(x_{n+1})<0$ for all $n=1,2,\dots,N$.
+Definition 1.4.
+The equation (1.1) is called locally dissipative at $a\in\mathbb{R}$ with configuration $(r,g)$, if
+  1. (i)
+    the function $g:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}$ satisfies, for all $x\in\mathbb{R}$, $\mu\in\mathcal{P}_{2}(\mathbb{R})$,
+    $2xV^{\prime}(x+a)-2x(W^{\prime}*\mu)(x)+\sigma^{2}(x+a)\leq-g(x^{2},\|\mu\|_{2}^{2});$
+  2. (ii)
+    the number $r>0$ and the function $g$ satisfy
+    $\begin{split}&g(\cdot,r^{2})\text{ is continuous and convex};\\ &\inf_{0\leq w\leq r^{2}}g(z,w)=g(z,r^{2});\\ &\inf_{z\geq r^{2}}g(z,r^{2})>0.\end{split}$
+Remark 1.2.
+It is straightforward to verify that Definition 1.4 provides an equivalent formulation of [36, Definition 1.2] in the one-dimensional setting.
+The concept of local dissipation was introduced by Zhang [57], and was later applied by Feng-Qu-Zhao [18] to the time-inhomogeneous McKean-Vlasov equation and by Bao-Wang [4] to the McKean–Vlasov SDEs with jumps.
 Theorem 1.4.
-Suppose Assumption 1 holds, and suppose the zero-crossing number $Z(V^{\prime})$ of $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is $(2n-1)$.
+Suppose Assumption 1 holds, and suppose the zero-crossing number $Z(V^{\prime})$ of $V^{\prime}:\mathbb{R}\to\mathbb{R}$ is (2n1)(2n-1).
 If the equation (1.1) is locally dissipative at $a_{1}<a_{2}<\cdots<a_{n}$ with configurations $\{(r_{k},g_{k})\}_{k=1}^{n}$, and if
-$r_{k}^{2}+r_{k+1}^{2}\leq\frac{1}{2}(a_{k+1}-a_{k})^{2}\ \text{ for all $k=1,2,\dots,n-1$},$ (2.1)
+$r_{k}+r_{k+1}\leq a_{k+1}-a_{k},\ \text{ for all $k=1,2,\dots,n-1$},$ (1.5)
 then
   1. (i)
-    (Total Orderedness) there are exactly $(2n-1)$ invariant measures $\{\nu_{k}\}_{k=1}^{2n-1}$ with $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$;
+    (Total Orderedness) there are exactly (2n1)(2n-1) invariant measures $\{\nu_{k}\}_{k=1}^{2n-1}$ with $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$;
   2. (ii)
@@
 So, the set of invariant measures cannot attract all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
-In other word, the order interval $[\underline{\nu},\overline{\nu}]$ in Theorem 2.1 (ii) cannot be replaced by the set of invariant measures.
-(iii). By the alternating arrangement of the invariant measures in Theorem 2.3, we have more information on the basins of attraction (see Theorem 4.2).
-  • if $\mu<_{\mathrm{st}}\nu^{1,\downarrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$ as $t\to\infty$;
-  • if $\nu^{k-1,\uparrow}_{t_{0}}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu^{k,\downarrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, $k=2,3,\dots,n-1$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$;
-  • if $\mu>_{\mathrm{st}}\nu^{n-1,\uparrow}_{t_{0}}$ for some $t_{0}\in\mathbb{R}$, then $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2n-1})\to 0$ as $t\to\infty$.
+In other word, the order interval $\underline{\nu},\overline{\nu}]$ in Theorem 1.2 (ii) cannot be replaced by the set of invariant measures.
+We next present a corollary providing the order position of invariant measures and also explicit expressions for certain densities in the basins of attraction.
+Consider a family of probability measures $\{\mu_{m}\}_{m\in\mathbb{R}}$ on $\mathbb{R}$, whose densities are given by
+$\frac{\mathop{}\!\mathrm{d}\mu_{m}}{\mathop{}\!\mathrm{d}x}\propto\frac{1}{\sigma^{2}(x)}\exp\bigg\{-2\int_{0}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg\}.$ (1.8)
 Corollary 1.5.
@@
 Remark 1.4.
-(i).
 For each $k=1,2,\dots,2n-2$, although the set $\{\mu_{m}\}_{m\in(m_{k},m_{k+1})}$ is a totally ordered arc with endpoints $\nu_{k}$, $\nu_{k+1}$, it is not a connecting orbit between invariant measures.
 One can check it from the Fokker-Planck equation.
-(ii).
-Corollary 2.4 also provides more information on basins of attraction.
-For each $k=1,2,\dots,n$, if $\mu_{m}\leq_{\mathrm{st}}\mu\leq_{\mathrm{st}}\mu_{\tilde{m}}$ for some $m_{2k-2}<m<\tilde{m}<m_{2k}$, then by the order-preservation of $P^{*}$ and the squeezing principle, we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2k-1})\to 0$ as $t\to\infty$.
-Besides abstract results, we have the following applications.
+As an application of our abstract theorems, we present below the classical double-well and multi-well examples.
+In contrast to Theorem 1.1, the confining potential $V$ is given explicitly in these cases, which allows us to determine explicit parameter ranges and to provide explicit descriptions of attracting sets.
+On the other hand, Proposition 4.3 provides a detailed analysis of the number of invariant measures under multiplicative noise, thereby verifying the a priori assumption in the main result of [36].
+Based on this analysis, the subsequent examples not only address basins of attraction with respect to both the partial order and the topology, but also strengthen the results on connecting orbits for the corresponding examples considered in [36].
+In particular, even under multiplicative noise, the classical double-well equation (1.9) admits connecting orbits; for the multi-well equation (1.11) we identify exactly five invariant measures together with a precise heteroclinic structure connecting them via monotone connecting orbits.
 Theorem 1.6 (Double-well landscapes, Figure 1).
@@
 If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
-$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{51\sqrt{17}-107}{256}\ \text{ for all $x\in\mathbb{R}$},$ (2.5)
-and let $r=\frac{9-\sqrt{17}}{8}$, then the following statements are true.
+$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\inf_{x\in\mathbb{R}}\sigma^{2}(x)\leq\sup_{x\in\mathbb{R}}\sigma^{2}(x)<\frac{51\sqrt{17}-107}{256},$ (1.10)
+and let $r=\frac{9-\sqrt{17}}{8}$, then
   1. (i)
@@
 Figure 1: Phase diagram for double-well landscapes
-Theorem 2.6 (Double-well with flat bottoms, Figure 1).
-Consider the following one-dimensional McKean-Vlasov SDE,
-$\mathop{}\!\mathrm{d}X_{t}=-\left[V^{\prime}(X_{t})+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.6)
-where
-$V^{\prime}(x)=\begin{cases}x\big{(}x-\frac{31}{32}\big{)}\big{(}x+\frac{31}{32}\big{)},&x\in\big{(}-\frac{31}{32},\frac{31}{32}\big{)},\\ 0,&x\in\big{(}-1,-\frac{31}{32}\big{]}\cup\big{[}\frac{31}{32},1\big{)},\\ x(x-1)(x+1),&x\in(-\infty,-1]\cup[1,\infty).\end{cases}$
-If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
-$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{51\sqrt{17}-171}{256}\ \text{ for all $x\in\mathbb{R}$},$ (2.7)
-and let $r=\frac{9-\sqrt{17}}{8}$, then the exactly same results as in Theorem 2.5 are valid.
-Theorem 2.7 (Asymmetric double-well, Figure 1).
-Consider the following one-dimensional McKean-Vlasov SDE,
-$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+2)+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.8)
-If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
-$\theta\geq\frac{4(3+\sqrt{3})}{3},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{3(39+55\sqrt{33})}{256}\ \text{ for all $x\in\mathbb{R}$},$
-and let $r_{-2}=\frac{(\sqrt{33}-15)^{2}}{64}$, $r_{1}=\frac{6-3\sqrt{3}}{2}$, then the following statements are true.
-  1. (i)
-    There are exactly three invariant measures, $\nu_{-2},\nu_{0},\nu_{1}\in\mathcal{P}_{\infty}(\mathbb{R})$ with $\nu_{-2}<_{\mathrm{st}}\nu_{0}<_{\mathrm{st}}\nu_{1}$;
-  2. (ii)
-    For all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\nu_{-2},\nu_{1}])\to 0$ as $t\to\infty$;
-  3. (iii)
-    There are a decreasing connecting orbit from $\nu_{0}$ to $\nu_{-2}$ and an increasing connecting orbit from $\nu_{0}$ to $\nu_{1}$;
-  4. (iv)
-    There hold $\nu_{-2}\in B(\delta_{-2},r_{-2})$, $\nu_{1}\in B(\delta_{1},r_{1})$, and
-    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{-2})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\leq 0}B(\delta_{-2+a},r_{-2})$},$
-    $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{1+a},r_{1})$}.$
-Figure 2: Phase diagram for multi-well landscapes
 Theorem 1.7 (Multi-well landscapes, Figure 2).
@@
 If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
-$\theta\geq 8\sqrt{5+\sqrt{13}},\quad 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\frac{4(13\sqrt{13}-35)}{27}\ \text{ for all $x\in\mathbb{R}$},$
-and let $r=\frac{5-\sqrt{13}}{3}$, then the following statements are true.
+$\theta\geq 8\sqrt{5+\sqrt{13}},\quad 0<\inf_{x\in\mathbb{R}}\sigma^{2}(x)\leq\sup_{x\in\mathbb{R}}\sigma^{2}(x)<\frac{4(13\sqrt{13}-35)}{27},$
+and let $r=\frac{\sqrt{15-3\sqrt{13}}}{3}$, then
   1. (i)
@@
     $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{2})\to 0\ \text{ as $t\to\infty$ for all $\mu\in\bigcup_{a\geq 0}B(\delta_{2+a},r)$}.$
-Theorem 2.9 (Double-well with perturbation).
+Figure 2: Phase diagram for multi-well landscapes
+Owing to our result on the number of invariant measures (Proposition 4.3), our framework is also capable of handling the case where $V^{\prime}$ vanishes on an interval.
+Theorem 1.8 (Double-well with flat bottoms, Figure 1).
 Consider the following one-dimensional McKean-Vlasov SDE,
-$\mathop{}\!\mathrm{d}X_{t}=-\left[X_{t}(X_{t}-1)(X_{t}+1)+f(X_{t})+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (2.10)
-If the parameter $\theta$ and the Lipschitz functions $f\colon\mathbb{R}\to\mathbb{R}$, $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
-$\theta\geq 0,\quad\left\lvert f(x)\right\rvert\leq L,\ \ 0<\underline{\sigma}^{2}<\sigma^{2}(x)<\overline{\sigma}^{2}\ \text{ for all $x\in\mathbb{R}$},$
-then the exactly same results as in Theorem 2.1 are valid.
+$\mathop{}\!\mathrm{d}X_{t}=-\left[V^{\prime}(X_{t})+\theta\left(X_{t}-\mathbf{E}[X_{t}]\right)\right]\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (1.12)
+where
+$V^{\prime}(x)=\begin{cases}x\big(x-\frac{31}{32}\big)\big(x+\frac{31}{32}\big),&x\in\big(-\frac{31}{32},\frac{31}{32}\big),\\ 0,&x\in\big(-1,-\frac{31}{32}\big]\cup\big[\frac{31}{32},1\big),\\ x(x-1)(x+1),&x\in(-\infty,-1]\cup[1,\infty).\end{cases}$
+If the parameter $\theta$ and the Lipschitz function $\sigma\colon\mathbb{R}\to\mathbb{R}$ satisfy
+$\theta\geq\frac{27(9+\sqrt{17})}{128},\quad 0<\inf_{x\in\mathbb{R}}\sigma^{2}(x)\leq\sup_{x\in\mathbb{R}}\sigma^{2}(x)<\frac{51\sqrt{17}-171}{256},$ (1.13)
+and let $r=\frac{9-\sqrt{17}}{8}$, then the exactly same results as in Theorem 1.6 are valid.
 2 Properties of Stochastic Order
@@
 2.1 Preliminaries of stochastic order
-Let $\mathcal{P}(\mathbb{R})$ be the set of probability measures on $\mathbb{R}$, and let $\mathcal{P}_{p}(\mathbb{R})$ be the set of those having finite $p$-th moments,
-$\mathcal{P}_{p}(\mathbb{R}):=\left\{\mu\in\mathcal{P}(\mathbb{R}):\left\lVert\mu\right\rVert_{p}=\left(\int_{\mathbb{R}}\left\lvert x\right\rvert^{p}\mathop{}\!\mathrm{d}\mu(x)\right)^{1/p}<\infty\right\}.$
-For $p\in[1,\infty)$, equip $\mathcal{P}_{p}(\mathbb{R})$ with the $p$-Wasserstein metric defined as below,
-$\mathcal{W}_{p}(\mu,\nu):=\inf\left\{\left(\mathbf{E}[\left\lvert X-Y\right\rvert^{p}]\right)^{1/p}:\mathcal{L}(X)=\mu,\,\mathcal{L}(Y)=\nu\right\},$
-where $\mathcal{L}(\cdot)$ represents the law of a random variable. Then $(\mathcal{P}_{p}(\mathbb{R}),\mathcal{W}_{p})$ is a complete metric space (see Villani [68, Theorem 6.18]).
-Next, we review the so-called “stochastic order”, which is also known as stochastic domination. By Definition 2.1, for $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$,
-$\mu\leq_{\mathrm{st}}\nu\text{ if and only if }\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\mu\leq\int_{\mathbb{R}}f\mathop{}\!\mathrm{d}\nu\text{ for all bounded increasing functions $f\colon\mathbb{R}\to\mathbb{R}$}.$
-Then “$\leq_{\mathrm{st}}$” is a partial order on $\mathcal{P}_{2}(\mathbb{R})$.
-Indeed, the reflexivity and transitivity are obvious, and the antisymmetry can be obtained by taking $f=\mathbf{1}_{(x,\infty)}$ for any $x\in\mathbb{R}$.
-Moreover, as shown in [48, Lemma 3.2], the stochastic order is a closed partial order relation in $\mathcal{P}_{2}(\mathbb{R})$, namely, the subset $\{(\mu,\nu):\mu\leq_{\mathrm{st}}\nu\}$ is closed in the product space $\mathcal{P}_{2}(\mathbb{R})\times\mathcal{P}_{2}(\mathbb{R})$.
+Let $\mathcal{P}(\mathbb{R}^{d})$ be the set of probability measures on $\mathbb{R}^{d}$, and let $\mathcal{P}_{p}(\mathbb{R}^{d})$ be the set of those having finite $p$-th moments,
+$\mathcal{P}_{p}(\mathbb{R}^{d}):=\left\{\mu\in\mathcal{P}(\mathbb{R}^{d}):\left\lVert\mu\right\rVert_{p}=\left(\int_{\mathbb{R}^{d}}\left\lvert x\right\rvert^{p}\mathop{}\!\mathrm{d}\mu(x)\right)^{1/p}<\infty\right\}.$
+For $p\in[1,\infty)$, equip $\mathcal{P}_{p}(\mathbb{R}^{d})$ with the $p$-Wasserstein metric defined as below,
+$\mathcal{W}_{p}(\mu,\nu):=\inf\Big\{\big(\mathbf{E}[\left\lvert X-Y\right\rvert^{p}]\big)^{1/p}:\mathcal{L}(X)=\mu,\,\mathcal{L}(Y)=\nu\Big\},$
+where $\mathcal{L}(\cdot)$ represents the law of a random variable. Then $(\mathcal{P}_{p}(\mathbb{R}^{d}),\mathcal{W}_{p})$ is a complete metric space (see Villani [54, Theorem 6.18]).
+Next, we review the so-called “stochastic order”, which is also known as stochastic domination; we refer to, e.g., Kamae-Krengel-O’Brien [32], Lindvall [34, Chapter IV] for a detailed discussion.
+Give $\mathbb{R}^{d}$ an order relation by coordinate-wise comparison, i.e., $x\leq y$ if and only if $x_{i}\leq y_{i}$ for all $i=1,2,\dots,d$.
+For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the stochastic order is defined by the following,
+[…]
+where “increasing” means $f(x)\leq f(y)$ whenever $x\leq y$.
+Then $\leq_{\mathrm{st}}$ is a partial order on $\mathcal{P}_{2}(\mathbb{R}^{d})$.
+Indeed, the reflexivity and transitivity are obvious, and the antisymmetry can be obtained by taking $f=\mathbf{1}_{(x,\infty)^{d}}$ for any $x\in\mathbb{R}^{d}$.
+Moreover, as shown in [36, Lemma 2.1], the stochastic order is a closed partial order relation in $\mathcal{P}_{2}(\mathbb{R}^{d})$, namely, the subset $\{(\mu,\nu):\mu\leq_{\mathrm{st}}\nu\}$ is closed in the product space $\mathcal{P}_{2}(\mathbb{R}^{d})\times\mathcal{P}_{2}(\mathbb{R}^{d})$.
 That is to say, $\mu\leq_{\mathrm{st}}\nu$ whenever $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ and $\mathcal{W}_{2}(\nu_{n},\nu)\to 0$ as $t\to\infty$ and $\mu_{n}\leq_{\mathrm{st}}\nu_{n}$ for all $n$.
 If $\mu\leq_{\mathrm{st}}\nu$ and $\mu\neq\nu$, we write $\mu<_{\mathrm{st}}\nu$.
-The reversed signs $\geq_{\mathrm{st}}$, $>_{\mathrm{st}}$ are used in the usual way.
-For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$ with $\mu\leq_{\mathrm{st}}\nu$, the order interval is defined by $[\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\mu\leq_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$, $(\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}):\mu<_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$ and similarly we can define $[\mu,\nu)$ and $(\mu,\nu)$.
-Measures $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R})$ are said to be order related, if $\mu\leq_{\mathrm{st}}\nu$ or $\nu\leq_{\mathrm{st}}\mu$ holds.
+The reversed signs $\geq_{\mathrm{st}}$, $_{\mathrm{st}}$ are used in the usual way.
+For $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with $\mu\leq_{\mathrm{st}}\nu$, the order interval is defined by $\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}^{d}):\mu\leq_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$, $(\mu,\nu]=\{\rho\in\mathcal{P}_{2}(\mathbb{R}^{d}):\mu<_{\mathrm{st}}\rho\leq_{\mathrm{st}}\nu\}$ and similarly we can define $\mu,\nu)$ and $(\mu,\nu)$.
+Measures $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ are said to be order related, if $\mu\leq_{\mathrm{st}}\nu$ or $\nu\leq_{\mathrm{st}}\mu$ holds.
 Otherwise, they are order unrelated.
-Given two subsets $A$ and $B$ of $\mathcal{P}_{2}(\mathbb{R})$, we write $A\leq_{\mathrm{st}}B$ ($A<_{\mathrm{st}}B$) when $\mu\leq_{\mathrm{st}}\nu$ ($\mu<_{\mathrm{st}}\nu$) holds for each choice of $\mu\in A$ and $\nu\in B$.
-Sets $A,B\subset\mathcal{P}_{2}(\mathbb{R})$ are said to be order related, if $A\leq_{\mathrm{st}}B$ or $B\leq_{\mathrm{st}}A$ holds.
-A totally ordered set (which is also referred as chain) $A$ in $\mathcal{P}_{2}(\mathbb{R})$ means that, any two measures $\mu$ and $\nu$ in $A$ are order related.
-A measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$ is an upper bound (lower bound) of a set $A\subset\mathcal{P}_{2}(\mathbb{R})$ if $\mu\geq_{\mathrm{st}}A$ ($\mu\leq_{\mathrm{st}}A$).
+Given two subsets $A$ and $B$ of $\mathcal{P}_{2}(\mathbb{R}^{d})$, we write $A\leq_{\mathrm{st}}B$ ($A<_{\mathrm{st}}B$) when $\mu\leq_{\mathrm{st}}\nu$ ($\mu<_{\mathrm{st}}\nu$) holds for each choice of $\mu\in A$ and $\nu\in B$.
+Sets $A,B\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ are said to be order related, if $A\leq_{\mathrm{st}}B$ or $B\leq_{\mathrm{st}}A$ holds.
+A totally ordered set (which is also referred as chain) $A$ in $\mathcal{P}_{2}(\mathbb{R}^{d})$ means that, any two measures $\mu$ and $\nu$ in $A$ are order related.
+A measure $\mu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ is an upper bound (lower bound) of a set $A\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ if $\mu\geq_{\mathrm{st}}A$ ($\mu\leq_{\mathrm{st}}A$).
 If the set $A$ has an upper (lower) bound, it is said to be bounded from above (below).
@@
 There are some complementary material of the stochastic order that we will use later.
-Firstly, a useful equivalent characterization of the stochastic order is Strassen’s theorem (see e.g., Lindvall [47, equation (3)]), which says $\mu\leq_{\mathrm{st}}\nu$ if and only if there exist two random variables $X,Y$ such that
+Firstly, a useful equivalent characterization of the stochastic order is Strassen’s theorem (see e.g., Lindvall [34, Theorem IV.2.4] or [35, equation (3)]), which says $\mu\leq_{\mathrm{st}}\nu$ if and only if there exist two random variables $X,Y$ such that
 $\mathcal{L}(X)=\mu,\quad\mathcal{L}(Y)=\nu,\quad X\leq Y\ \mathrm{a.s.}$
-Secondly, according to Hiai-Lawson-Lim [40, Proposition 3.11], test functions $f$ in Definition 2.1 can be restricted as bounded continuous increasing.
-If $\mu$, $\nu$ both have finite first moments, it is equivalent for test functions ranging over all $1$-Lipschitz increasing functions (see Fritz-Perrone [36, Theorem 4.2.1]), where a $1$-Lipschitz function $f$ means $\left\lvert f(x)-f(y)\right\rvert\leq\left\lvert x-y\right\rvert$ for all $x,y\in\mathbb{R}$.
+Secondly, according to Hiai-Lawson-Lim [25, Proposition 3.11], test functions $f$ in Definition 1.1 can be restricted as bounded continuous increasing.
+If $\mu$, $\nu$ both have finite first moments, it is equivalent for test functions ranging over all $1$-Lipschitz increasing functions (see Fritz-Perrone [19, Theorem 4.2.1]), where a $1$-Lipschitz function $f$ means $\left\lvert f(x)-f(y)\right\rvert\leq\left\lvert x-y\right\rvert$ for all $x,y\in\mathbb{R}^{d}$.
 We may confine test functions in such different sets whenever convenient.
-Additionally, an important observation regarding the definition of $\mu\leq_{\mathrm{st}}\nu$, is that when we take the test function as $f=-\mathbf{1}_{(-\infty,x]}$ or $f=\mathbf{1}_{(x,\infty)}$, we have $\mu\big{(}(-\infty,x]\big{)}\geq\nu\big{(}(-\infty,x]\big{)}$ and $\mu\big{(}(x,\infty)\big{)}\leq\nu\big{(}(x,\infty)\big{)}$, for any $x\in\mathbb{R}$.
+Additionally, an important observation regarding the definition of $\mu\leq_{\mathrm{st}}\nu$, is that when we take the test function as $f=-\mathbf{1}_{(-\infty,x]^{d}}$ or $f=\mathbf{1}_{(x,\infty)^{d}}$, we have $\mu\big((-\infty,x]^{d}\big)\geq\nu\big((-\infty,x]^{d}\big)$ and $\mu\big((x,\infty)^{d}\big)\leq\nu\big((x,\infty)^{d}\big)$, for any $x\in\mathbb{R}^{d}$.
 2.2 Equivalent characterization of order boundedness
-The following lemma gives an equivalent characterization of the stochastic order by the comparison of cumulative distribution functions.
+We begin with an equivalent characterization of the stochastic order in $\mathcal{P}_{2}(\mathbb{R})$ by comparison of cumulative distribution functions.
 Lemma 2.1.
@@
 Take the test function $f=-\mathbf{1}_{(-\infty,x]}$ in the definition of $\mu_{1}\leq_{\mathrm{st}}\mu_{2}$, it follows that $F_{1}(x)\geq F_{2}(x)$, for any $x\in\mathbb{R}$.
-($\Leftarrow$) By Hiai-Lawson-Lim [40, Proposition 3.11], it suffices to show $\int f\mathop{}\!\mathrm{d}\mu_{1}\leq\int f\mathop{}\!\mathrm{d}\mu_{2}$ for all bounded continuous increasing functions $f$.
+($\Leftarrow$) By Hiai-Lawson-Lim [25, Proposition 3.11], it suffices to show $\int f\mathop{}\!\mathrm{d}\mu_{1}\leq\int f\mathop{}\!\mathrm{d}\mu_{2}$ for all bounded continuous increasing functions $f$.
 Let $f:\mathbb{R}\to\mathbb{R}$ be such a function, and we may assume $f$ is non-negative.
@@
 and we have $f_{n}\uparrow f$ as $n\to\infty$.
-Note that the set $\{f>k/2^{n}\}$ is an open interval of the form $(x_{k},\infty)$, where (,):=assign($\infty$,$\infty$):$\emptyset$( ∞ , ∞ ) := ∅, since $f$ is increasing continuous.
-So we know that
+Note that the set $\{f>k/2^{n}\}$ is an open interval of the form $(x_{k},\infty)$, where $(\infty,\infty):=\emptyset$, since $f$ is increasing continuous.
+So we have
 $\int f_{n}\mathop{}\!\mathrm{d}\mu_{i}=\sum_{k=1}^{n2^{n}}\frac{1}{2^{n}}\mu_{i}\big((x_{k},\infty)\big),\ \ i=1,2.$
@@
 It remains to show that $\nu\geq_{\mathrm{st}}M$.
-Since $F_{\nu}=F=\inf_{\mu\in M}F_{\mu}$, we know that $F_{\nu}\leq F_{\mu}$ for all $\mu\in M$.
+Since $F_{\nu}=F=\inf_{\mu\in M}F_{\mu}$, we have $F_{\nu}\leq F_{\mu}$ for all $\mu\in M$.
 Lemma 2.1 yields that $\nu\geq_{\mathrm{st}}\mu$ for all $\mu\in M$, which completes the proof.
@@
 Likewise, the condition (2.2) implies the existence of $\inf_{\leq_{\mathrm{st}}}M$ in $\mathcal{P}_{2}(\mathbb{R})$.
+Remark 2.3.
+We remark here that in higher dimensional case, there is also an equivalent characterization of order boundedness for a subset of $\mathcal{P}_{2}(\mathbb{R}^{d})$:
+A subset $M\subset\mathcal{P}_{2}(\mathbb{R}^{d})$ is order bounded from above (below) in $\mathcal{P}_{2}(\mathbb{R}^{d})$ if and only if the marginals of $M$ on each component are order bounded from above (below).
+We give a brief proof here.
+($\Rightarrow$) is obvious, and we only show ($\Leftarrow$).
+We will find a lower bound $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$ of $M$, and the construction of an upper bound is similar, which is omitted.
+[…]
+$\hat{\mu}\big((-\infty,u]\big)=\mu\big(\mathbb{R}^{d}\setminus(u,\infty)^{d}\big)\leq\sum_{i=1}^{d}\mu\big(\mathbb{R}^{i-1}\times(-\infty,u]\times\mathbb{R}^{d-i}\big).$ (2.11)
+Given that marginals of $M$ on each component are order bounded below, Proposition 2.2 (ii) implies, for each $i=1,2,\dots,d$,
+$\int_{0}^{\infty}\bigg(\sup_{\mu\in M}\mu\big(\mathbb{R}^{i-1}\times(-\infty,-u]\times\mathbb{R}^{d-i}\big)\bigg)u\mathop{}\!\mathrm{d}u<\infty.$ (2.12)
+Combining (2.11)(2.12) yields
+$\int_{0}^{\infty}\bigg(\sup_{\mu\in M}\hat{\mu}\big((-\infty,-u]\big)\bigg)u\mathop{}\!\mathrm{d}u\leq\sum_{i=1}^{d}\int_{0}^{\infty}\bigg(\sup_{\mu\in M}\mu\big(\mathbb{R}^{i-1}\times(-\infty,-u]\times\mathbb{R}^{d-i}\big)\bigg)u\mathop{}\!\mathrm{d}u<\infty,$
+so there is some $\nu\in\mathcal{P}_{2}(\Delta)$ such that $\nu\leq_{\mathrm{st}}\hat{\mu}$ for all $\mu\in M$.
+Now we observe $\nu\in\mathcal{P}_{2}(\mathbb{R}^{d})$, and $\nu\leq_{\mathrm{st}}\hat{\mu}\leq_{\mathrm{st}}\mu$ for all $\mu\in M$.
+This finishes the proof.
 Corollary 2.3.
@@
 If $M$ is an order bounded subset of $\mathcal{P}_{2}(\mathbb{R})$, then $M$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$.
-In particular, any order interval $[\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$.
+In particular, any order interval $\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$.
 Proof.
@@
 Note that for any $\mu\in M$, by choosing $f^{+}_{N}(x)=|x|^{2}\mathbf{1}_{\{x:x\geq N\}}$ and $f^{-}_{N}(x)=-|x|^{2}\mathbf{1}_{\{x:x\leq-N\}}$ respectively, we deduce
-$\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\overline{\mu}(x),\ \text{ and }\ \int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\underline{\mu}(x).$
+[…]
 Thus, for all $\mu\in M$,
 $\int_{\{x:|x|\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\overline{\mu}(x)+\int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\underline{\mu}(x).$
-Then we know that
+Then we have
 $\lim_{N\to\infty}\sup_{\mu\in M}\int_{\{x:|x|\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\mu(x)\leq\lim_{N\to\infty}\Bigg(\int_{\{x:x\geq N\}}|x|^{2}\mathop{}\!\mathrm{d}\overline{\mu}(x)+\int_{\{x:x\leq-N\}}|x|^{2}\mathop{}\!\mathrm{d}\underline{\mu}(x)\Bigg)=0.$
 Hence, $M$ is $2$-uniformly integrable in $\mathcal{P}_{2}(\mathbb{R})$.
-In particular, an order interval $[\mu,\nu]$ is closed due to the closedness of the stochastic order, so it is compact in $\mathcal{P}_{2}(\mathbb{R})$. ∎
+In particular, an order interval $\mu,\nu]$ is closed due to the closedness of the stochastic order, so it is compact in $\mathcal{P}_{2}(\mathbb{R})$. ∎
 Remark 2.4.
@@
 $\mu_{n}=\frac{1}{n^{2}\log n}\delta_{-n}+\left(1-\frac{2}{n^{2}\log n}\right)\delta_{0}+\frac{1}{n^{2}\log n}\delta_{n}.$
-Since $\mathcal{W}_{2}(\mu_{n},\delta_{0})=(2/\log n)^{1/2}\to 0$ as $n\to\infty$, the set $\{\mu_{n}\}_{n=3}^{\infty}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. Note that
-$\inf_{n\geq 3}F_{\mu_{n}}(x)=\left\{\begin{aligned} &0,&&x<0,\\ &1-\frac{1}{3^{2}\log 3},&&0\leq x<3,\\ &1-\frac{1}{k^{2}\log k},&&k-1\leq x<k,\ k\geq 4.\end{aligned}\right.$
+Since $\mathcal{W}_{2}(\mu_{n},\delta_{0})=(2/\log n)^{1/2}\to 0$ as $n\to\infty$, the set $\{\mu_{n}\}_{n=2}^{\infty}$ is relatively compact in $\mathcal{P}_{2}(\mathbb{R})$. Note that
+$\inf_{n\geq 2}F_{\mu_{n}}(x)=\left\{\begin{aligned} &0,&&x<0,\\ &1-\frac{1}{2^{2}\log 2},&&0\leq x<2,\\ &1-\frac{1}{k^{2}\log k},&&k-1\leq x<k,\ k\geq 3.\end{aligned}\right$.
 It gives
-$\int_{0}^{\infty}\bigg{(}\sup_{n\geq 3}\mu_{n}\big{(}(u,\infty)\big{)}\bigg{)}u\mathop{}\!\mathrm{d}u=\int_{0}^{\infty}\big{[}1-\inf_{n\geq 3}F_{\mu_{n}}(u)\big{]}u\mathop{}\!\mathrm{d}u\geq\sum_{k=3}^{\infty}\frac{k}{(k+1)^{2}\log(k+1)}=\infty,$
+$\int_{0}^{\infty}\bigg(\sup_{n\geq 2}\mu_{n}\big((u,\infty)\big)\bigg)u\mathop{}\!\mathrm{d}u=\int_{0}^{\infty}\big[1-\inf_{n\geq 2}F_{\mu_{n}}(u)\big]u\mathop{}\!\mathrm{d}u\geq\sum_{k=2}^{\infty}\frac{k}{(k+1)^{2}\log(k+1)}=\infty,$
 so the condition (2.1) does not hold, and $M$ has no upper bound in $\mathcal{P}_{2}(\mathbb{R})$ with respect to the stochastic order.
@@
 Corollary 2.5.
-Let $[\mu,\nu]$ be an order interval and $\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, $\rho\in[\mu,\nu]$. Then the following convergences are equivalent.
+Let $\mu,\nu]$ be an order interval and $\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, $\rho\in[\mu,\nu]$. Then the following convergences are equivalent.
   1. (i)
@@
   2. (ii)
-    $\mathcal{W}_{p}(\rho_{n},\rho)\to 0$ as $n\to\infty$, for all $21\leq p<$ 2;
+    $\mathcal{W}_{p}(\rho_{n},\rho)\to 0$ as $n\to\infty$, for all $1\leq p<2$;
   3. (iii)
@@
 Proof.
-It is obvious that (i)$\Rightarrow$(ii)$\Rightarrow$(iii), and we only need to show (iii)$\Rightarrow$(i).
-By Proposition 3.4, we know that $[\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$.
-Then for any subsequence $\{\rho_{n_{k}}\}_{k\geq 1}\subset\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, there exists a subsequence $\{\rho_{n_{k_{j}}}\}_{j\geq 1}$ of $\{\rho_{n_{k}}\}_{k\geq 1}$ and a $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R})$ such that $\mathcal{W}_{2}(\rho_{n_{k_{j}}},\rho_{0})\to 0$ as $j\to\infty$.
+[…]
+By Proposition 2.4, we have $\mu,\nu]$ is compact in $\mathcal{P}_{2}(\mathbb{R}^{d})$.
+Then for any subsequence $\{\rho_{n_{k}}\}_{k\geq 1}\subset\{\rho_{n}\}_{n\geq 1}\subset[\mu,\nu]$, there exists a subsequence $\{\rho_{n_{k_{j}}}\}_{j\geq 1}$ of $\{\rho_{n_{k}}\}_{k\geq 1}$ and a $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ such that $\mathcal{W}_{2}(\rho_{n_{k_{j}}},\rho_{0})\to 0$ as $j\to\infty$.
 Thus, $\rho_{n_{k_{j}}}\to\rho_{0}$ weakly as $j\to\infty$.
@@
 In Section 3.1, we target on the structure of global attractors of general semigroups on $\mathcal{P}_{2}(\mathbb{R})$ that preserve the stochastic order (namely, order-preserving semigroups).
-In Section 4.2, we prove a comparison theorem for extrinsic-law-dependent SDEs, and obtain the generation of order-preserving semigroup by the equation (1.1).
 The coming down from infinity property of hyper-dissipative McKean-Vlasov SDEs is shown in Section 3.2.
@@
 3.1 Global attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
-In this subsection, we utilize the concepts of dissipative dynamical systems (see e.g., Hale [39]) to study the asymptotic behaviour of a semigroup on $\mathcal{P}_{2}(\mathbb{R})$ that preserves the stochastic order.
-With a Lyapunov condition, [14, 34] study the unique ergodicity of order-preserving Markov processes.
-Here, we focus on the structure of the global attractor of an order-preserving semigroup.
-Based on the existence of a globally attracting order bounded set, we prove that a global attractor of an order-preserving semigroup must be encompassed by two order related invariant measures, and these two invariant measures attract all solutions initiated from the regions above and below them with respect to the stochastic order, respectively (see Theorem 4.2).
+The notion of order-preserving has already appeared in the study of uniqueness of invariant measures, ergodicity, and synchronization phenomena for certain classes of Markov processes; see, for instance, [5, 37, 46].
+In this subsection, we instead utilize concepts from the theory of dissipative dynamical systems (see, e.g., Hale [24]) to study the asymptotic behaviour of a general semigroup on $\mathcal{P}_{2}(\mathbb{R})$ that preserves the stochastic order, allowing in particular for the presence of multiple invariant measures.
+Based on the existence of a globally attracting order bounded set, we prove that a global attractor of an order-preserving semigroup must be encompassed by two order related invariant measures, and these two invariant measures attract all solutions initiated from the regions above and below them with respect to the stochastic order, respectively (see Proposition 3.2).
 Firstly, we give the definition of a semiflow on $\mathcal{P}_{2}(\mathbb{R})$.
@@
 (i) is a direct consequence of the definition of $\omega$-limit sets.
-(ii)-(iv) are standard results for dissipative systems on complete metric spaces, for their proof one can see Hale [39, Lemma 3.1.1, Lemma 3.2.1 and Theorem 3.4.2 (ii)] respectively.
+(ii)-(iv) are standard results for dissipative systems on complete metric spaces, for their proof one can see Hale [24, Lemma 3.1.1, Lemma 3.2.1 and Theorem 3.4.2 (ii)] respectively.
 ∎
 Clearly, Lemma 3.1 (iii) implies that, if the global attractor $A$ exists, then $\omega(B)\subset A$ for any bounded set $B\subset\mathcal{P}_{2}(\mathbb{R})$.
-In the following, we give the definition of order-preserving semigroups.
+In the following, we give the definition of order-preserving semiflows.
 Definition 3.5.
@@
 Now, we give our main theorem in this subsection concerning the structure of the global attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$.
-Theorem 4.2.
+Proposition 3.2.
 (Global attractor).
@@
 Since $P^{*}$ is order-preserving and $\rho\geq_{\mathrm{st}}\mu$, we have $P^{*}_{t_{k}}\rho\geq_{\mathrm{st}}P^{*}_{t_{k}}\mu$ for any $k\geq 1$.
-Together with the closedness of the stochastic order, $\mathcal{W}_{2}(P^{*}_{t_{k}}\mu,\overline{\nu})\to 0$ as $k\to\infty$ implies that $\pi\geq_{\mathrm{st}}\overline{\nu}$.
+Together with the closedness of the stochastic order, $\mathcal{W}_{2}(P^{*}_{t_{k}}\mu,\overline{\nu})\to 0$ as $k\to\infty$, one has $\pi\geq_{\mathrm{st}}\overline{\nu}$.
 Since the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$, one has $\pi\in\omega(\rho)\subset\omega(K)\subset[\underline{\nu},\overline{\nu}]$.
@@
 (i).
-If we consider a semiflow preserving the stochastic order acting only on a closed subspace $B$ of $\mathcal{P}_{2}(\mathbb{R})$ (rather than the entire space $\mathcal{P}_{2}(\mathbb{R})$), by repeating the same proof, Theorem 4.2 still holds, where the order interval within $B$ is the restriction of the order interval to $B$.
-In fact, for semiflow preserving some closed partial order relation on a general complete metric space, by assuming the atttracting set $K$ is compact and order bounded, Theorem 4.2 still holds.
+If we consider a semiflow preserving the stochastic order acting only on a closed subspace $B$ of $\mathcal{P}_{2}(\mathbb{R})$ (rather than the entire space $\mathcal{P}_{2}(\mathbb{R})$), by repeating the same proof, Proposition 3.2 still holds, where the order interval within $B$ is the restriction of the order interval to $B$.
+In fact, for semiflow preserving some closed partial order relation on a general complete metric space, by assuming the atttracting set $K$ is compact and order bounded, Proposition 3.2 still holds.
 (ii).
-For semiflows preserving some closed partial order relation on a strongly ordered space, Hirsch [41, Theorem 3.3] implies the unique equilibrium must be globally attracting (Theorem 4.2 (vi)).
+For semiflows preserving some closed partial order relation on a strongly ordered space, Hirsch [27, Theorem 3.3] implies the unique equilibrium must be globally attracting (Proposition 3.2 (vi)).
 However, Proposition A.1 shows that the strong ordering relation generated by $\leq_{\mathrm{st}}$ is empty.
-So, [41, Theorem 3.3] is not applicable here.
-4.2 Comparison principle for extrinsic-law-dependent SDEs
+So, [27, Theorem 3.3] is not applicable here.
+3.2 Coming down from infinity
+In this subsection, we prove hyper-dissipative McKean-Vlasov SDEs have the coming down from infinity property.
+It means, there exists a compact subset of $\mathcal{P}_{2}(\mathbb{R})$ attracting all bounded sets.
+As a byproduct, we show $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is a compact map.
 Now let us go back to the McKean-Vlasov equation (1.1).
@@
 When the starting time $s=0$, we denote the solution $\{X_{t}^{0,\xi}\}_{t\geq 0}$ by $\{X_{t}^{\xi}\}_{t\geq 0}$.
-Next, we present a comparison theorem for extrinsic-law-dependent SDEs with respect to the stochastic order.
-Let $C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ be the space of continuous functions from $[s,\infty)$ to $\mathcal{P}_{2}(\mathbb{R})$.
-For any $\mu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $\xi\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$, we denote $\{X_{t}^{s,\mu,\xi}\}_{t\geq s}$ by the solution of the following extrinsic-law-dependent SDE
-$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mu_{t})(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},\ \ X_{s}=\xi,\ \ t\geq s.$ (4.1)
-Similarly, we denote the solution $\{X_{t}^{0,\mu,\xi}\}_{t\geq 0}$ by $\{X_{t}^{\mu,\xi}\}_{t\geq 0}$.
-Then we have the following comparison principle.
-Theorem 4.3.
-Suppose Assumption 1 holds. Let $\mu,\nu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $T>0$. Assume
-$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{t}(x)\leq\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{t}(x),\ \text{ for all }\ t\in[s,s+T].$ (4.2)
-If $\xi,\eta\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ and $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, then $\mathcal{L}(X_{t}^{s,\mu,\xi})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{s,\nu,\eta})$ for all $t\in[s,s+T]$.
-Proof.
-Since $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, Strassen’s theorem (see Lindvall [47, equation (3)]) yields that there exist $\tilde{\xi},\tilde{\eta}\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ such that
-[…]
-It suffices to show that $\mathcal{L}(X_{t}^{s,\mu,\tilde{\xi}})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{s,\nu,\tilde{\eta}})$ for all $t\in[s,s+T]$.
-Note that
-$\mathop{}\!\mathrm{d}X^{s,\mu,\tilde{\xi}}_{t}$ $=\Big{[}-V^{\prime}(X^{s,\mu,\tilde{\xi}}_{t})-\theta X^{s,\mu,\tilde{\xi}}_{t}+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{t}(x)\Big{]}\mathop{}\!\mathrm{d}t+\sigma(X^{s,\mu,\tilde{\xi}}_{t})\mathop{}\!\mathrm{d}B_{t};$ (4.4)
-$\mathop{}\!\mathrm{d}Y^{s,\nu,\tilde{\eta}}_{t}$ $=\Big{[}-V^{\prime}(Y^{s,\nu,\tilde{\eta}}_{t})-\theta Y^{s,\nu,\tilde{\eta}}_{t}+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{t}(x)\Big{]}\mathop{}\!\mathrm{d}t+\sigma(Y^{s,\nu,\tilde{\eta}}_{t})\mathop{}\!\mathrm{d}B_{t}.$
-Since $\theta>0$, according to the comparison theorem for usual SDEs [37, Theorem 1.2], we know that
-[…]
-The desired result is proved. ∎
-Remark 4.2.
-Theorem 4.3 is not only used in the proof of the order-preservation for semigroups of McKean-Vlasov SDEs, but also crucial in the proof of our main result Theorem 2.3.
-Associated to the McKean-Vlasov SDE (1.1), there is a semigroup $P_{t}^{*}$ acting on $\mathcal{P}_{2}(\mathbb{R})$ by
-$P_{t}^{*}\mu:=\mathcal{L}(X_{t}^{\xi})\ \text{ with }\ \mathcal{L}(\xi)=\mu.$ (4.5)
-In the following, we show that $P_{t}^{*}$ is an order-preserving semiflow.
-Theorem 4.4.
-Under Assumption 1, the semigroup $P^{*}:[0,\infty)\times\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ is an order-preserving semiflow.
-Proof.
-We first prove $P^{*}$ is order-preserving.
-It suffices to show that $P^{*}_{t}\rho_{1}\leq_{\mathrm{st}}P^{*}_{t}\rho_{2}$ for any $\rho_{1}<_{\mathrm{st}}\rho_{2}$ and all $t\geq 0$.
-By Strassen’s theorem, there exist $\xi_{1},\xi_{2}\in L^{2}(\Omega,\mathcal{F}_{0},\mathbf{P})$ such that
-$\mathcal{L}(\xi_{1})=\rho_{1},\ \ \mathcal{L}(\xi_{2})=\rho_{2},\ \ \text{ and}\ \ \xi_{1}\leq\xi_{2},\ \mathbf{P}-a.s.$
-Since $\rho_{1}\neq\rho_{2}$, we have $\mathbf{P}\{\xi_{1}<\xi_{2}\}>0$ and hence
-$\mathbf{E}[\xi_{1}]<\mathbf{E}[\xi_{2}].$ (4.6)
-Let
-$T:=\sup\{t>0:\mathbf{E}[X_{s}^{\xi_{1}}]<\mathbf{E}[X_{s}^{\xi_{2}}],\ \text{ for all }s\in[0,t]\}.$
-Note that $\mathbf{E}[X_{t}^{\xi_{i}}]$ is continuous in $t$ for $i=1,2$, then it follows that $T>0$. The proof is divided into the following two cases:
-$\textbf{Case 1}.\ \ T=\infty,\ \ \textbf{Case 2}.\ \ T<\infty.$
-In case 1, let $\mu^{i}_{t}:=\mathcal{L}(X_{t}^{\xi_{i}}),t\in[0,\infty)$, $i=1,2$. Then we know that
-$X_{t}^{\mu^{i},\xi_{i}}=X_{t}^{\xi_{i}},\ \text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{t}(x)\leq\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{t}(x),\ \text{ for all }\ t\geq 0.$
-Then we conclude from Theorem 4.3 that
-$P^{*}_{t}\rho_{1}=\mathcal{L}(X_{t}^{\mu^{1},\xi_{1}})\leq_{\mathrm{st}}\mathcal{L}(X_{t}^{\mu^{2},\xi_{2}})=P^{*}_{t}\rho_{2},\ \text{ for all }t\geq 0.$
-In case 2, we have $0<T<\infty$. Let $\mu^{i}$ be defined as in case 1, we have $X_{t}^{\mu^{i},\xi_{i}}=X_{t}^{\xi_{i}}$ for all $t\geq 0$ and
-$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{t}(x)<\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{t}(x),\ t\in[0,T)\ \text{ and }\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{T}(x)=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{T}(x).$
-Using Theorem 4.3 again, we deduce
-$P^{*}_{t}\rho_{1}\leq_{\mathrm{st}}P^{*}_{t}\rho_{2},\ \text{ for all }t\in[0,T].$
-Note that $\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{1}_{T}(x)=\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu^{2}_{T}(x)$, by the proof of (4.6), we conclude that $P^{*}_{T}\rho_{1}=P^{*}_{T}\rho_{2}$.
-Then we have $P^{*}_{t}\rho_{1}=P^{*}_{t}\rho_{2}$ for all $t\geq T$.
-So the order-preservation of $P^{*}$ is proved. It remains to show that $P^{*}$ is jointly continuous, which follows from [48, Proposition 2.3] directly. ∎
-4.3 Coming down from infinity
-In this subsection, we prove hyper-dissipative McKean-Vlasov SDEs have the coming down from infinity property.
-It means, there exists a compact subset of $\mathcal{P}_{2}(\mathbb{R})$ attracting all bounded sets.
-As a byproduct, we show $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is a compact map.
 Suppose Assumption 1 (ii)-(iv) hold. Then for any $x\in\mathbb{R}$, $\mu\in\mathcal{P}_{2}(\mathbb{R})$, we have
@@
 ∎
-Indeed, we want to show the solutions are $p$-integrable for all $p\geq 2$ if the initial conditions are just square integrable. For this purpose, we need the following extrinsic-law case.
+Indeed, we want to show the solutions are $p$-integrable for all $p\geq 2$ if the initial conditions are just square integrable.
+For this purpose, we need the following extrinsic-law case.
+Let $C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ be the space of continuous functions from $s,\infty)$ to $\mathcal{P}_{2}(\mathbb{R})$.
+For any $\mu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $\xi\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$, we denote $\{X_{t}^{s,\mu,\xi}\}_{t\geq s}$ by the solution of the following extrinsic-law-dependent SDE
+$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mu_{t})(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t},\ \ X_{s}=\xi,\ \ t\geq s.$ (3.4)
+Similarly, we denote the solution $\{X_{t}^{0,\mu,\xi}\}_{t\geq 0}$ by $\{X_{t}^{\mu,\xi}\}_{t\geq 0}$.
 Lemma 3.4.
@@
 The proof of the claim is divided into the following two cases.
-$\begin{split}&\textbf{Case 1:}\ y_{t}\geq\frac{\alpha\delta}{2}t,\ \text{ if }\ y_{s}\leq A_{t},\ \text{ for all }\ s\in[0,t];\\ &\textbf{Case 2:}\ y_{t}\geq A_{t},\ \text{ if there exists $s\in[0,t]$ such that }\ y_{s}\geq A_{t}.\end{split}$
+[…]
 Case 1 is a direct consequence of (3.5). To prove Case 2, we assume otherwise that $y_{t}<A_{t}$. Note that $y_{\cdot}$ is continuous in $[s,t]$ and $y_{s}\geq A_{t}$, let
-$r:=\inf\{u\in[s,t]:y_{v}\leq A_{t},\text{ for all }v\in[u,t]\},$
+$r:=\inf\{v\in[s,t]:y_{u}\leq A_{t},\text{ for all }u\in[v,t]\},$
 we conclude that $s\leq r<t$ and
@@
 which gives a contradiction to (3.5) and the fact that $y_{u_{0}}\leq A_{t}$. We complete the proof of the claim.
-By (4.11), we know that, for any $p\geq 1$ and $t>0$,
+By (3.6), we have, for any $p\geq 1$ and $t>0$,
 $\begin{split}|X_{t}^{\mu,\xi}|^{p}&\leq 2^{p-1}(|Y_{t}|^{p}+|Z_{t}|^{p})\\ &\leq 2^{p-1}(y_{t}^{-\frac{p}{\delta}}+(Z_{t}^{*})^{p})\\ &\leq 2^{p-1}\max\bigg\{\Big(\frac{\alpha\delta}{2}t\Big)^{-\frac{p}{\delta}},\Big(\frac{\theta}{\alpha}(\mu_{t}^{*})^{2}+\frac{2\beta+\gamma}{\alpha}+\frac{\gamma}{\alpha}(Z_{t}^{*})^{2\kappa}\Big)^{\frac{p}{2+\delta}}\bigg\}+2^{p-1}(Z_{t}^{*})^{p}\\ &\leq C(p)\Big(t^{-\frac{p}{\delta}}+(\mu_{t}^{*})^{\frac{2p}{2+\delta}}+1+(Z_{t}^{*})^{\kappa p}\Big).\end{split}$
@@
 Proof.
-Denote $\mu_{t}:=\mathcal{L}(X_{t}^{\xi})$ for $t\geq 0$, and by Lemma 4.5 and Lemma 4.6, we know that for any $t>0$,
+Denote $\mu_{t}:=\mathcal{L}(X_{t}^{\xi})$ for $t\geq 0$, and by Lemma 3.3 and Lemma 3.4, we have for any $t>0$,
 $\big(\mathbf{E}\big[|X_{t}^{\xi}|^{p}\big]\big)^{1/p}\leq C(p)\Big(t^{-\frac{1}{\delta}}+t^{\frac{\kappa}{2}}+\big(\mathbf{E}[|\xi|^{2}]\big)^{\frac{1}{2+\delta}}\Big).$ (3.7)
@@
 Now we are ready to prove the coming down from infinity under the hyper-dissipative condition.
-Theorem 4.8.
+Proposition 3.6.
 Suppose Assumption 1 holds.
@@
 $K_{p}:=\Big\{\mu\in\mathcal{P}_{p}(\mathbb{R}):\|\mu\|_{p}\leq\frac{e}{e-1}C(p)\Big\},\ \text{ and }\ K:=\bigcap_{p\geq 2}K_{p},$ (3.8)
-where the constant $C(p)$ come from Lemma 4.7.
+where the constant $C(p)$ comes from Lemma 3.5.
 Clearly, $\delta_{0}\in K$, so $K$ is nonempty.
@@
 Suppose a $\mathcal{P}_{2}$-Cauchy sequence $\{\mu_{n}\}_{n\geq 1}\subset K$ converges to a probability measure $\mu\in\mathcal{P}_{2}(\mathbb{R})$, and it remains to check $\mu\in K$, i.e., $\mu\in K_{p}$ for all $p\geq 2$.
-By [68, Theorem 6.9], $\mu_{n}$ weakly converges to $\mu$.
+By [54, Theorem 6.9], $\mu_{n}$ weakly converges to $\mu$.
 For any $p\geq 2$, since $\{\mu_{n}\}_{n\geq 1}\subset K_{p}$, then
@@
 3.3 Proof of Theorem 1.2 and Corollary 1.3
-We first give the following lemma, which is a direct consequence of Feng-Qu-Zhao [33, Theorem 4.8].
+We first give the following lemma, which is a direct consequence of Feng-Qu-Zhao [17, Theorem 4.8].
 Let $C_{b}(\mathbb{R};\mathcal{P}_{2}(\mathbb{R}))$ be the collection of all bounded continuous paths $\{\nu_{t}\}_{t\in\mathbb{R}}$ in $\mathcal{P}_{2}(\mathbb{R})$.
@@
 Obviously, $\mu_{i}$ is an invariant measures of (3.11).
-Denote by $X_{t}^{i,x}$ the solution of (4.16) with initial value $x$ and starting time 00, $i=1,2$.
-Since $m_{1}\leq m_{2}$, it follows from the comparison theorem (Theorem 4.3) that
+Denote by $X_{t}^{i,x}$ the solution of (3.11) with initial value $x$ and starting time 0, $i=1,2$.
+Since $m_{1}\leq m_{2}$, it follows from the comparison theorem for usual SDEs [20, Theorem 1.2] that
 $\mathcal{L}(X_{t}^{1,0})\leq_{\mathrm{st}}\mathcal{L}(X_{t}^{2,0})\ \text{ for all $t\geq 0$}.$ (3.12)
-By Lemma 4.9 (ii), $\mu_{i}$ is the unique invariant measure of (4.16) and
+By Lemma 3.7, $\mu_{i}$ is the unique invariant measure of (3.11) and
 $d_{2}(\mathcal{L}(X_{t}^{i,0}),\mu_{i})\to 0\ \text{ as $t\to\infty$, for $i=1,2$},$ (3.13)
-Next, we want to show the convergence in $\mathcal{P}_{2}(\mathbb{R})$. Observe that, for any bounded continuous function $f:\mathbb{R}\to\mathbb{R}$, with $|f|_{\infty}:=\sup_{x\in\mathbb{R}}|f(x)|$,
-$\Big{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{1}(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{2}(x)\Big{|}\leq|f|_{\infty}d_{2}\big{(}\nu_{1},\nu_{2}\big{)},$ (4.19)
-and
-$\Big{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1}(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{2}(x)\Big{|}\leq d_{2}\big{(}\nu_{1},\nu_{2}\big{)}.$ (4.20)
-Combine (4.18)-(4.20), and it follows from the equivalent characterization of $\mathcal{W}_{2}$-convergence (see Villani [68, Theorem 6.9]) that
+Note that $\mathcal{W}_{2}(\mu,\nu)\leq\sqrt{2d_{2}(\mu,\nu)}$ (see e.g. [54, Theorem 6.15]). Then we conclude that
 $\mathcal{W}_{2}(\mathcal{L}(X_{t}^{i,0}),\mu_{i})\to 0\ \text{ as $t\to\infty$, for $i=1,2$}.$ (3.14)
@@
 By Lemma 3.8, we see $\mathcal{M}$ is totally ordered with respect to the stochastic order.
-Take $K\subset\mathcal{P}_{\infty}(\mathbb{R})$ as the set in (4.13), and Theorem 4.8 (iv) gives $\mathcal{M}\subset K\subset\mathcal{P}_{\infty}(\mathbb{R})$.
+Take $K\subset\mathcal{P}_{\infty}(\mathbb{R})$ as the set in (3.8), and Proposition 3.6 (iv) gives $\mathcal{M}\subset K\subset\mathcal{P}_{\infty}(\mathbb{R})$.
 The finiteness of $\mathcal{M}$ is postponed to the end of the proof.
 (ii).
-By Theorem 4.4, the semigroup $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is an order-preserving semiflow.
-Theorem 4.8 gives a eventually positively invariant, order bounded, closed set $K\subset\mathcal{P}_{2}(\mathbb{R})$, and this set $K$ attracts all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
-Now, apply Theorem 4.2, whose statements (i)(ii) say that, there are two invariant measures $\underline{\nu},\overline{\nu}\in\mathcal{P}_{2}(\mathbb{R})$ such that the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$.
+By [36, Theorem 3.5], the semigroup $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is an order-preserving semiflow.
+Proposition 3.6 gives a eventually positively invariant, order bounded, closed set $K\subset\mathcal{P}_{2}(\mathbb{R})$, and this set $K$ attracts all bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
+Now, apply Proposition 3.2, whose statements (i)(ii) say that, there are two invariant measures $\underline{\nu},\overline{\nu}\in\mathcal{P}_{2}(\mathbb{R})$ such that the global attractor $\omega(K)\subset[\underline{\nu},\overline{\nu}]$.
 By the definition of global attractor, one has $\mathcal{M}\subset\omega(K)$.
 Obviously, $\underline{\nu}=\inf_{\leq_{\mathrm{st}}}\mathcal{M}$ and $\overline{\nu}=\sup_{\leq_{\mathrm{st}}}\mathcal{M}$.
-Next, we need to show the global convergence to the order interval $[\underline{\nu},\overline{\nu}]$.
+Next, we need to show the global convergence to the order interval $\underline{\nu},\overline{\nu}]$.
 Since $\omega(K)$ is the global attractor, the inclusion $\omega(K)\subset[\underline{\nu},\overline{\nu}]$ tells us, for all bounded sets $B\subset\mathcal{P}_{2}(\mathbb{R})$, $\mathcal{W}_{2}(P_{t}^{*}B,[\underline{\nu},\overline{\nu}])\to 0$ as $t\to\infty$.
-(iii). This is a consequence of Theorem 4.2 (iii).
+(iii). This is a consequence of Proposition 3.2 (iii).
 (i)-finiteness. Now we prove that $\mathcal{M}$ is a finite set. Let $\{\mu_{m}\}_{m\in\mathbb{R}}$ be probability measures given in (1.8), i.e.,
@@
 We assume there are infinitely many invariant measures, and argue by contradiction.
-By the statement (ii), the set of invariant measures $\mathcal{M}$ is contained in an order interval $[\underline{\nu},\overline{\nu}]$.
-So the boundedness of $[\underline{\nu},\overline{\nu}]$ in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition 3.4) entails boundedness of the zero set of $F$ in $\mathbb{R}$.
+By the statement (ii), the set of invariant measures $\mathcal{M}$ is contained in an order interval $\underline{\nu},\overline{\nu}]$.
+So the boundedness of $\underline{\nu},\overline{\nu}]$ in $\mathcal{P}_{2}(\mathbb{R})$ (see Proposition 2.4) entails boundedness of the zero set of $F$ in $\mathbb{R}$.
 Since invariant measures are infinitely many, the zero set of $F$ has an accumulation point.
@@
 In Remark 1.3, we show that the set of invariant measures cannot attract all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
-Here, by Theorem 4.2 (i)(ii) and Theorem 4.8 (i), there exists a connected compact invariant set (the global attractor) within $[\underline{\nu},\overline{\nu}]$ which attracts all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
+Here, by Proposition 3.2 (i)(ii) and Proposition 3.6 (i), there exists a connected compact invariant set (the global attractor) within $\underline{\nu},\overline{\nu}]$ which attracts all the bounded subsets of $\mathcal{P}_{2}(\mathbb{R})$.
 Proof of Corollary 1.3.
 The uniqueness of invariant measure enatils that $\underline{\nu}=\nu=\overline{\nu}$.
-Then the antisymmetry of stochastic order implies that $[\underline{\nu},\overline{\nu}]=\{\nu\}$.
+Then the antisymmetry of stochastic order implies that $\underline{\nu},\overline{\nu}]=\{\nu\}$.
 Hence, this corollary is a direct consequence of Theorem 1.2 (ii).
 ∎
-5 Unbounded Open Neighbourhood Attracted by Supremum (Infimum) Invariant Measure
-This section attempts to prove Theorem 2.3 and Corollary 2.4.
-In Section 5.1, we find the number of invariant measures of the equation (1.1) is bounded by the zero-crossing number of the function $V^{\prime}$.
-In Section 5.2, we show the existence of positively invariant neighbourhoods of invariant measures, for both the semigroup $P_{t}^{*}$ and the measure-iterating map $\Psi$.
-After providing a structure theorem of local attractors of order-preserving semigroups in Section 5.3, we proceed to prove Theorem 2.3 and Corollary 2.4 by tracing solutions of (1.1) backward along connecting orbits.
-5.1 Number of invariant measures of granular media equations
+4 Basins of attraction
+This section attempts to prove Theorem 1.1, Theorem 1.4, Corollary 1.5, Theorem 1.6, Theorem 1.7 and Theorem 1.8.
+In Section 4.1, we prove that the number of invariant measures of the equation (1.1) is bounded by the zero-crossing number of the function $V^{\prime}$.
+Using the fact that solutions can be traced backward along connecting orbits, we establish a local attractor theorem in Section 4.2 to show that the forward limit of a connecting orbit is locally attracting.
+Then, we finish the proof of our main results, especially the verification of the local dissipative condition under the assumptions of Theorem 1.1.
+4.1 Number of invariant measures
 In this subsection, we provide the prior information on the number of invariant measures.
-Assuming $V^{\prime}$ has $(2n-1)$ simple zeros, Alecio [1] shows that the equation (1.1) has only one invariant measure if $\left\lvert\sigma\right\rvert$ is above a certain upper critical threshold, and has $(2n-1)$ invariant measures if $\left\lvert\sigma\right\rvert$ is below another lower critical threshold.
-There is a gap between these two critical thresholds, and the information is missing in the intermediate stage.
-Here, for any non-degenerate $\sigma$, we prove that the number of invariant measures of the equation (1.1) does not exceed the zero-crossing number of $V^{\prime}$ in Propoaition 5.3.
-This facilitates the quantitative results in our main theorems.
+We show the number of invariant measures is at most the zero-crossing number of the derivative $V^{\prime}$ of confining potential, regardless of the ranges of $\theta$ and $\sigma$ (see Proposition 4.3).
+This helps to make our main theorems quantitative (Theorem 1.6-1.8), and overcomes difficulties caused by qualitative results and missing parameter in the previous literature on phase transitions (see e.g., [1, 52]).
+One the other hand, this proposition allows $V^{\prime}$ vanishes on some interval (Theorem 1.8), which is out of the settings of many existing works.
+Also, it provides an insight that the essential factor in the maximal number of invariant measures is the zero-crossing number of $V^{\prime}$, instead of the number of simple zeros of $V^{\prime}$.
 Let $f:\mathbb{R}\to\mathbb{R}$ be a function.
 Recall that its zero-crossing number $Z(f)$ is the maximal integer $N$ such that, there exists an increasing sequence $x_{1}<x_{2}<\cdots<x_{N}<x_{N+1}$ satisfying $f(x_{n})f(x_{n+1})<0$ for all $n=1,2,\dots,N$.
-It is easy to check that, if $V^{\prime}$ has $(2n-1)$ simple zeros, then $Z(V^{\prime})=2n-1$.
+It is easy to check that, if $V^{\prime}$ has (2n1)(2n-1) simple zeros, then $Z(V^{\prime})=2n-1$.
 Next, we will see the zero-crossing number of the function $V^{\prime}$ is essential to determine the number of invariant measures.
@@
 $f(x_{-})<0,\ \ \ f(x_{+})>0.$
-Now define another function $\tilde{F}:\mathbb{R}\to\mathbb{R}$ by
+Now define another function $\tilde{F}$$\mathbb{R}\to\mathbb{R}$ by
 $\tilde{F}(m):=\int_{\mathbb{R}}f(x)\exp\bigg\{-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg\}\mathop{}\!\mathrm{d}x,$
@@
 Assume the conclusion is true for $n$.
-If the function $f$ is continuous and $Z(f)=n+1$, then we may assume there is an increasing sequence $-\infty=x_{-1}<x_{0}<\dots<x_{n}<x_{n+1}=\infty$ such that, $f$ changes its sign on any two adjacent intervals $(x_{k-1},x_{k})$, $k=0,1,\dots,n+1$.
-Define $\tilde{F}:\mathbb{R}\to\mathbb{R}$ by
+If the function $f$ is continuous and $Z(f)=n+1$, then we may assume there is an increasing sequence $\infty=x_{-1}<x_{0}<\dots<x_{n}<x_{n+1}=\infty$ such that, $f$ changes its sign on any two adjacent intervals $(x_{k-1},x_{k})$, $k=0,1,\dots,n+1$.
+Define $\tilde{F}$$\mathbb{R}\to\mathbb{R}$ by
 $\tilde{F}(m):=\int_{\mathbb{R}}f(x)\exp\bigg\{-2\int_{x_{0}}^{x}\frac{V^{\prime}(y)+\theta y-\theta m}{\sigma^{2}(y)}\mathop{}\!\mathrm{d}y\bigg\}\mathop{}\!\mathrm{d}x.$
@@
 Proposition 4.3.
-Suppose Assumption 1 holds. If the function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ has zero-crossing number $Z(V^{\prime})=2n-1$, then the equation (1.1) has at most $(2n-1)$ invariant measures.
+Suppose Assumption 1 holds. If the function $V^{\prime}:\mathbb{R}\to\mathbb{R}$ has zero-crossing number $Z(V^{\prime})=2n-1$, then the equation (1.1) has at most (2n1)(2n-1) invariant measures.
 Proof.
@@
 Assumption 1 (iv)(v) tell us $\theta>0$ and $\sigma^{2}(x)>0$ for all $x\in\mathbb{R}$, so $Z(V^{\prime}/\theta\sigma^{2})=Z(V^{\prime})=2n-1$.
-Lemma 5.2 implies that $F$ has at most $(2n-1)$ zeros, and thus, the equation (1.1) has at most $(2n-1)$ invariant measures.
+Lemma 4.2 implies that $F$ has at most (2n1)(2n-1) zeros, and thus, the equation (1.1) has at most (2n1)(2n-1) invariant measures.
 ∎
-5.2 Positively invariant neighbourhoods of invariant measures
-In this subsection, we prove the existence of positively invariant $\mathcal{P}_{2}(\mathbb{R})$-neighbourhoods of invariant measures under $P^{*}$ when assuming the locally dissipative condition.
-The existence of positively invariant $\mathcal{P}_{2}(\mathbb{R})$-neighbourhoods is also obtained for the measure-iterating map $\Psi$ defined in (5.13).
-Theorem 5.4.
-Suppose that Assumption 1 holds.
-If the locally dissipative condition holds at $a\in\mathbb{R}$ with configuration $(r_{a},g_{a})$, then the equation (1.1) has an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$.
-Moreover, $B(\delta_{a},r_{a})$ is positively invariant sets, i.e.,
-$P_{t}^{*}B(\delta_{a},r_{a})\subset B(\delta_{a},r_{a}),\ \text{ for all }\ t\geq 0.$ (5.3)
-Proof.
-By (ii) in the definition of local dissipation (Definition 2.3), we know that there exists $0<\tilde{r}_{a}<r_{a}$ such that
-$g_{a}(z,r_{a}^{2})>0,\ \text{ for all }z\geq\tilde{r}_{a}^{2}.$ (5.4)
-To prove (5.3), it suffices to prove that $\overline{B(\delta_{a},r)}$ is positively invariant for all $\tilde{r}_{a}\leq r<r_{a}$, which implies that $B(\delta_{a},r)$ is positively invariant for all $\tilde{r}_{a}<r\leq r_{a}$.
-For any fixed $\tilde{r}_{a}\leq r<r_{a}$ and $\mu\in\overline{B(\delta_{a},r)}$, let
-$\mathcal{T}_{\mu}:=\sup\big{\{}t\geq 0:P_{s}^{*}\mu\in\overline{B(\delta_{a},r)}\ \text{ for all }\ 0\leq s\leq t\big{\}}.$
-To prove that $\overline{B(\delta_{a},r)}$ is positively invariant, it suffices to show that $\mathcal{T}_{\mu}=\infty$.
-By Theorem 4.4, we know that $P^{*}:\mathbb{R}\times\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ is jointly continuous, so $\alpha_{t}:=\mathcal{W}_{2}^{2}(P^{*}_{t}\mu,\delta_{a})$ is continuous in $t$.
-Then it is equivalent to show that
-$\mathcal{T}_{\mu}=\sup\big{\{}t\geq 0:\alpha_{s}\leq r^{2}\ \text{ for all }\ 0\leq s\leq t\big{\}}=\infty.$
-Otherwise, if $\mathcal{T}_{\mu}<\infty$, then it follows from the continuity of $\alpha$ and definition of $\mathcal{T}_{\mu}$ that there exists $T_{0}\geq\mathcal{T}_{\mu}$ and $\epsilon_{0}>0$ such that
-$\alpha_{T_{0}}=r^{2},\ \text{ and }\ r^{2}<\alpha_{t}\leq r_{a}^{2}\ \text{ for all }\ T_{0}<t\leq T_{0}+\epsilon_{0}.$ (5.5)
-Let $X_{t}^{\xi}$ be the solution of (1.1) with intial distribution $\mathcal{L}(\xi)=\mu$ and starting time 00.
-Then $P_{t}^{*}\mu=\mathcal{L}(X_{t}^{\xi})$ for all $t\geq 0$.
-For any $a\in\mathbb{R}$, denote $\tau^{a}:\mathbb{R}\to\mathbb{R}$ by the translation $x\mapsto x+a$ and $\tau^{a}_{\#}:\mathcal{P}(\mathbb{R})\to\mathcal{P}(\mathbb{R})$ by $\tau^{a}_{\#}\mu=\mu\circ(\tau^{a})^{-1}$.
-Applying Itô’s formula to $|X_{t}^{\xi}-a|^{2}$ on $[T_{0},T_{0}+\epsilon_{0}]$, we have for any $T_{0}<t\leq T_{0}+\epsilon_{0}$,
-$\begin{split}|X_{t}^{\xi}-a|^{2}&=|X_{T_{0}}^{\xi}-a|^{2}+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s}\\ &\ \ \ \ +\int_{T_{0}}^{t}\big{(}-2(X_{s}^{\xi}-a)V^{\prime}(X_{s}^{\xi})-2(X_{s}^{\xi}-a)(W^{\prime}*P^{*}_{s}\mu)(X_{s}^{\xi})+\sigma^{2}(X_{s}^{\xi})\big{)}\mathop{}\!\mathrm{d}s\\ &=|X_{T_{0}}^{\xi}-a|^{2}+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s}\\ &\ \ \ \ +\int_{T_{0}}^{t}\big{(}-2(X_{s}^{\xi}-a)V^{\prime}(X_{s}^{\xi})-2(X_{s}^{\xi}-a)(W^{\prime}*\tau^{-a}_{\#}P^{*}_{s}\mu)(X_{s}^{\xi}-a)+\sigma^{2}(X_{s}^{\xi})\big{)}\mathop{}\!\mathrm{d}s\\ &\leq|X_{T_{0}}^{\xi}-a|^{2}-\int_{T_{0}}^{t}g_{a}\big{(}|X_{s}^{\xi}-a|^{2},r_{a}^{2}\big{)}\mathop{}\!\mathrm{d}s+2\int_{T_{0}}^{t}(X_{s}^{\xi}-a)\sigma(X_{s}^{\xi})\mathop{}\!\mathrm{d}B_{s},\end{split}$ (5.6)
-where we have used the fact
-$\|\tau^{-a}_{\#}P^{*}_{s}\mu\|_{2}^{2}=\int_{\mathbb{R}}|x-a|^{2}\mathop{}\!\mathrm{d}P^{*}_{s}\mu(x)=\mathcal{W}_{2}^{2}(P^{*}_{s}\mu,\delta_{a})=\alpha_{s}\in[r^{2},r_{a}^{2}]\ \text{ for all $s\in[T_{0},T_{0}+\varepsilon]$}.$
-Taking expectation on both sides and by the convexity of $g_{a}(\cdot,r_{a}^{2})$, we get
-$\alpha_{T_{0}+\epsilon_{0}}\leq r^{2}-\int_{T_{0}}^{T_{0}+\epsilon_{0}}g_{a}(\alpha_{t},r_{a}^{2})\mathop{}\!\mathrm{d}t.$ (5.7)
-Then by (5.4), (5.5) and (5.7), we conclude that $\alpha_{T_{0}+\epsilon_{0}}<r^{2}$ which contradicts (5.5). Hence,
-$P_{t}^{*}\overline{B(\delta_{a},r)}\subset\overline{B(\delta_{a},r)},\ \text{ for all }\ \tilde{r}\leq r<r_{a},$ (5.8)
-and (5.3) follows.
-It remains to prove there exists an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$. We first give the following two claims.
-Claim 1: Let $S\subset\mathcal{P}_{2}(\mathbb{R})$ be a bounded closed convex positively invariant set under $P^{*}$.
-If $P_{t}^{*}$ is compact for some $t>0$, then there exists a fixed point $\mu_{t}$ of $P_{t}^{*}$ in $S$, i.e., $P^{*}_{t}\mu_{t}=\mu_{t}$.
-Claim 2: Let $S\subset\mathcal{P}_{2}(\mathbb{R})$ be a bounded closed convex positively invariant set under $P^{*}$, and $P_{t}^{*}$ is compact for all $t>0$, then there exists an invariant measure $\mu$ in $S$, i.e., $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
-Proof of Claim 1: Let $S_{t}=P_{t}^{*}S$ and $\overline{\mathrm{Conv}(S_{t})}$ be the closed convex hull of $S_{t}$.
-Since $P_{t}^{*}$ is compact, it follows $S_{t}$ is relatively compact.
-Noting also that $S$ is closed convex positively invariant, we have $\overline{\mathrm{Conv}(S_{t})}\subset S$ is compact in $\mathcal{P}_{2}(\mathbb{R})$, and
-$P_{t}^{*}\overline{\mathrm{Conv}(S_{t})}\subset P_{t}^{*}S=S_{t}\subset\overline{\mathrm{Conv}(S_{t})}.$
-Note that $\overline{\mathrm{Conv}(S_{t})}$ being compact convex in $\mathcal{P}_{2}(\mathbb{R})$ entails that it is compact convex in $\mathcal{P}_{1}(\mathbb{R})$.
-Let us consider the space $\mathcal{M}_{1}(\mathbb{R})$ of finite signed Borel measures on $\mathbb{R}$ with finite first moments,
-$\mathcal{M}_{1}(\mathbb{R}):=\left\{\mu\in\mathcal{M}(\mathbb{R}):\int_{\mathbb{R}}\left\lvert x\right\rvert\mathop{}\!\mathrm{d}\left\lvert\mu\right\rvert(x)<\infty\right\}$
-with the following norm:
-[…]
-where a $1$-Lipschitz function $f$ means $|f(x)-f(y)|\leq|x-y|$ for all $x,y\in\mathbb{R}$. By [48, Lemma 3.4 and Remark 3.2], we have the following properties of $\mathcal{M}_{1}(\mathbb{R})$:
-$(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$ is a normed vector space; (5.10)
-$\mathcal{P}_{1}(\mathbb{R})$ is a closed convex subset of $(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$. (5.11)
-Hence, $\overline{\mathrm{Conv}(S_{t})}$ is a compact convex subset of a normed vector space.
-Then the Tychonoff fixed-point theorem yields that there exists $\mu_{t}\in\overline{\mathrm{Conv}(S_{t})}\subset S$ such that $P^{*}_{t}\mu_{t}=\mu_{t}$.
-Proof of Claim 2: Fix a positive sequence $\{t_{n}\}_{n\geq 1}$ such that $t_{n}\downarrow 0$ as $n\to\infty$.
-For any $n\geq 1$, consider the continuous compact map $P^{*}_{t_{n}}:S\to S$.
-Then Claim 1 yields that there exists $\mu_{n}\in S$ such that $P^{*}_{t_{n}}\mu_{n}=\mu_{n}$.
-Let $k_{n}$ be an integer such that $k_{n}t_{n}>t_{1}$ for all $n\geq 1$.
-Then
-$\mu_{n}=P_{k_{n}t_{n}}^{*}\mu_{n}=P_{t_{1}}^{*}(P_{k_{n}t_{n}-t_{1}}^{*}\mu_{n})\in P_{t_{1}}^{*}(P_{k_{n}t_{n}-t_{1}}^{*}S)\subset P_{t_{1}}^{*}S\subset S\ \text{ for all $n\geq 1$}.$
-Since $P_{t_{1}}^{*}S\subset S$ is relatively compact and $S$ is closed, there exists a subsequence of $\{\mu_{n}\}_{n\geq 1}$, which still denotes by $\{\mu_{n}\}_{n\geq 1}$, and a $\mu\in S$ such that $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ as $n\to\infty$.
-Then we show that $P^{*}_{t}\mu=\mu$ for any $t\geq 0$.
-In fact, for any $t>0$, we have $t=k_{n,t}t_{n}+\tau_{n,t}$, $k_{n,t}\in\mathbb{N}$ and $\tau_{n,t}\in[0,t_{n})$. Then,
-$\mathcal{W}_{2}(P^{*}_{t}\mu,\mu)$ $\leq\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{t}\mu_{n},\mu)$
-$=\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}(P^{*}_{k_{n,t}t_{n}}\mu_{n}),\mu)$
-$=\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})+\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}\mu_{n},\mu).$
-Since $\tau_{n,t}\to 0$ and $\mathcal{W}_{2}(\mu_{n},\mu)\to 0$ as $n\to\infty$, by virtue of the joint continuity of the $P^{*}$, we have $\mathcal{W}_{2}(P^{*}_{t}\mu,P^{*}_{t}\mu_{n})\to 0$ and $\mathcal{W}_{2}(P^{*}_{\tau_{n,t}}\mu_{n},\mu)\to 0$ as $n\to\infty$.
-Hence, we obtain that $P^{*}_{t}\mu=\mu$ and the claim is proved.
-Now we are in a position to show that there exists an invariant measure in $B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$.
-Since $P^{*}_{t}$ is compact for all $t>0$ (see Theorem 4.8) and $\overline{B(\delta_{a},\tilde{r}_{a})}$ is bounded closed convex positively invariant under $P^{*}$, by Claim 2, there exists an invariant measure $\mu\in\overline{B(\delta_{a},\tilde{r}_{a})}\subset B(\delta_{a},r_{a})$.
-Moreover, Lemma 4.7 shows that $\mu\in\mathcal{P}_{\infty}(\mathbb{R})$.
-Now, we conclude that the invariant measure $\mu\in B(\delta_{a},r_{a})\cap\mathcal{P}_{\infty}(\mathbb{R})$.
-∎
-Now for any fixed $\mu\in\mathcal{P}_{2}(\mathbb{R})$, we consider the following SDE:
-$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mu)(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (5.12)
-Under Assumption 1, we know that SDE (5.12) has a unique invariant measure $\rho^{\mu}\in\mathcal{P}_{2}(\mathbb{R})$ (see e.g., Meyn-Tweedie [54, 55] or Lemma 4.9 (ii)).
-Define a measure-iterating map $\Psi:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by
-[…]
-Then we have the following theorem.
-Theorem 5.5.
-Suppose that Assumption 1 holds.
-If the equation (1.1) is locally dissipative at $a\in\mathbb{R}$ with configuration $(r_{a},g_{a})$, then we have $\Psi\big{(}\overline{B(\delta_{a},r_{a})}\big{)}\subset B(\delta_{a},r_{a})$.
-Proof.
-Note that there is some $\tilde{r}_{a}\in(0,r_{a})$ such that (5.4) holds.
-Now fix any $\mu\in\overline{B(\delta_{a},r_{a})}$.
-Let $\{X_{t}^{\mu,x}\}_{t\geq 0}$ be the solution of (5.12) with initial value $x\in\mathbb{R}$ and starting time 0.
-Similar to (5.6), applying Itô’s formula to $|X_{t}^{\mu,a}-a|^{2}$ yields
-$|X_{t}^{\mu,a}-a|^{2}\leq-\int_{0}^{t}g_{a}\big{(}|X_{s}^{\mu,a}-a|^{2},r_{a}^{2}\big{)}\mathop{}\!\mathrm{d}s+2\int_{0}^{t}(X_{s}^{\mu,a}-a)\sigma(X_{s}^{\mu,a})\mathop{}\!\mathrm{d}B_{s}.$
-Taking expectation on both sides and by the convexity of $g_{a}(\cdot,r_{a}^{2})$, we have that for any $t>0$,
-$g_{a}\bigg{(}\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}\mathop{}\!\mathrm{d}s,r_{a}^{2}\bigg{)}\leq\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}g_{a}\big{(}|X_{s}^{\mu,a}-a|^{2},r_{a}^{2}\big{)}\big{]}\mathop{}\!\mathrm{d}s\leq-\frac{1}{t}\mathbf{E}[|X_{t}^{\mu,a}-a|^{2}]\leq 0.$ (5.14)
-According to Lemma 4.9 (ii), we know that there exist $C_{\mu}>0,\lambda_{\mu}>0$ such that for any $t\geq 0$,
-$d_{2}\big{(}\mathcal{L}(X_{t}^{\mu,a}),\Psi(\mu)\big{)}\leq C_{\mu}(1+|a|^{2})e^{-\lambda_{\mu}t},$
-where
-$d_{2}(\nu_{1},\nu_{2})=\int_{\mathbb{R}}(1+|x|^{2})\mathop{}\!\mathrm{d}|\nu_{1}-\nu_{2}|(x),$ (5.15)
-Then
-$\big{|}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}-\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}\big{|}=\Big{|}\int_{\mathbb{R}}|x-a|^{2}\mathop{}\!\mathrm{d}\big{(}\mathcal{L}(X_{s}^{\mu,a})-\Psi(\mu)(x)\big{)}\Big{|}\leq 2C_{\mu}(1+|a|^{2})^{2}e^{-\lambda_{\mu}s}.$
-Thus
-$\lim_{t\to\infty}\frac{1}{t}\int_{0}^{t}\mathbf{E}\big{[}|X_{s}^{\mu,a}-a|^{2}\big{]}\mathop{}\!\mathrm{d}s=\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}.$
-Then the continuity of $g_{a}(\cdot,r_{a}^{2})$ and (5.14) yield $g_{a}(\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2},r_{a}^{2})\leq 0$.
-Hence, by (5.4), $\|\tau^{-a}_{\#}\Psi(\mu)\|_{2}^{2}\leq\tilde{r}_{a}^{2}$, i.e., $\Psi(\mu)\in\overline{B(\delta_{a},\tilde{r}_{a})}$.
-Hence,
-$\Psi\big{(}\overline{B(\delta_{a},r_{a})}\big{)}\subset\overline{B(\delta_{a},\tilde{r}_{a})}\subset B(\delta_{a},r_{a}).$
-The proof is completed. ∎
-Remark 5.1.
-Note that $\mu$ is an invariant measure of the equation (1.1) if and only if $\Psi(\mu)=\mu$, and then Theorem 5.5 indicates that there is no invariant measure on $\partial\overline{B(\delta_{a},r_{a})}:=\overline{B(\delta_{a},r_{a})}\setminus B(\delta_{a},r_{a})$.
 4.2 Local attractors of order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$
-To investigate the local dynamics near the invariant measures of McKean-Vlasov SDEs, we establish a connecting orbit theorem and a structure theorem on local attractors of order-preserving semigroups.
-Firstly, the following proposition concerns about connecting orbits on an order interval in $\mathcal{P}_{2}(\mathbb{R})$.
-It is a $\mathcal{P}_{2}$-version of the general connecting orbit theorem for an abstract monotone semiflow on a convex compact subset of a Hausdorff locally convex topological vector space (see Liu-Qu-Yao-Zhi [48, Theorem 4.2]).
-Proposition 5.6.
-Let $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ be an order-preserving semiflow, and suppose $\mu_{1}<_{\mathrm{st}}\mu_{2}$ are two order related invariant measures in $\mathcal{P}_{2}(\mathbb{R})$.
-Then at least one of the following holds:
-  1. (a)
-    $P_{t}^{*}$ has a third invariant measure $\nu\in\mathcal{P}_{2}(\mathbb{R})$ with $\mu_{1}<_{\mathrm{st}}\nu<_{\mathrm{st}}\mu_{2}$;
-  2. (b)
-    there exists an increasing connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\mu_{1}$ to $\mu_{2}$;
-  3. (c)
-    there exists a decreasing connecting orbit $\{\nu_{t}\}_{t\in\mathbb{R}}$ from $\mu_{2}$ to $\mu_{1}$.
-Proof.
-To prove the proposition, we only need to verify the assumptions (M1)-(M5) in [48, Theorem 4.2].
-  1. (M1)
-    Recall $(\mathcal{M}_{1}(\mathbb{R}),\left\lVert\cdot\right\rVert_{\mathcal{W}_{1}})$ is a normed space, see (5.10). A cone
-    $C:=\bigg{\{}\mu\in\mathcal{M}_{1}(\mathbb{R}):\int_{\mathbb{R}^{d}}f\mathop{}\!\mathrm{d}\mu\geq 0\text{ for all nonnegative 1-Lipschitz increasing functions $f$}\bigg{\}}$
-    in $\mathcal{M}_{1}(\mathbb{R})$ is given by [48, Lemma 3.5].
-  2. (M2)
-    By Proposition 3.4, we see that $S:=[\mu_{1},\mu_{2}]$ is compact in $\mathcal{P}_{2}(\mathbb{R})$, and hence, $S$ is compact in $\mathcal{P}_{1}(\mathbb{R})$.
-Recall $\mathcal{P}_{1}(\mathbb{R})$ is closed convex in $\mathcal{M}_{1}(\mathbb{R})$, see (5.11).
-Then $S$ is compact in $\mathcal{M}_{1}(\mathbb{R})$.
-The convexity and non-singleton of $S$ are obvious.
-By Corollary 3.5, (5.9) and the dual representation of $\mathcal{W}_{1}$, the 2-Wasserstein metric on $[\mu,\nu]$ induces the relative topology in $\mathcal{M}_{1}(\mathbb{R})$.
-  3. (M3)
-    By [48, Lemma 3.5], the partial order induced by the cone $C$ in the normed space $\mathcal{M}_{1}(\mathbb{R})$ coincides with the stochastic order when restricted on $\mathcal{P}_{2}(\mathbb{R})$.
-The infimum and the supremum of $S$ in $\mathcal{M}_{1}(\mathbb{R})$ are clearly endpoints $\mu_{1}$, $\mu_{2}$, of the order interval.
-  4. (M4)
-    The semiflow $P_{t}^{*}$ is order-preserving.
-  5. (M5)
-    $\mu_{1}$ and $\mu_{2}$ are invariant measures of $P^{\ast}_{t}$.
-Based on the existence of connecting orbits, we will present a theorem for order-preserving semigroups that ensures the local attraction of the forward limit of a connecting orbit.
+To investigate the local dynamics near the invariant measures, we will present a theorem for order-preserving semigroups that ensures the local attraction of the forward limit of a connecting orbit.
+We first apply Proposition 4.3 and the main result of [36] to obtain the existence of connecting orbits.
+Proof of (i)-(iii) in Theorem 1.4.
+By Proposition 4.3, the equation (1.1) has at most (2n1)(2n-1) invariant measures.
+Then, by applying [36, Theorem 1.1], it remains only to show the pairwise disjointness of $\{B(\delta_{a_{k}},r_{k})\}_{k=1}^{n}$.
+In fact, (1.5) implies
+$r_{i}+r_{j}\leq|a_{i}-a_{j}|,\ \text{ for all }\ 1\leq i\neq j\leq n.$ (4.3)
+Suppose on the contrary that there exists $\nu\in B(\delta_{a_{i}},r_{i})\cap B(\delta_{a_{j}},r_{j})$. Then,
+$|a_{i}-a_{j}|=\mathcal{W}_{2}(\delta_{a_{i}},\delta_{a_{j}})\leq\mathcal{W}_{2}(\delta_{a_{i}},\nu)+\mathcal{W}_{2}(\nu,\delta_{a_{j}})<r_{i}+r_{j},$
+which gives a contradiction to (4.3). ∎
+Based on the existence of connecting orbits, we give the following theorem ensuring the local attraction of the forward limit of a connecting orbit.
 Theorem 4.4.
@@
 Hence, $\mu_{0}$ is the unique invariant measure in $S$.
-Observing that $S$ has a lower bound $\nu^{-1,0}_{t_{-1}}\in S$ and an upper bound $\mu_{0}\in S$, Theorem 4.2 (i) (vi) and Remark 4.1 (i) entail that $\omega(S)=\{\mu_{0}\}$ is the global attractor of $P^{*}|_{S}$, where $P^{*}|_{S}$ means the semiflow $P^{*}$ restricted to $S$.
+Observing that $S$ has a lower bound $\nu^{-1,0}_{t_{-1}}\in S$ and an upper bound $\mu_{0}\in S$, Proposition 3.2 (i) (vi) and Remark 3.1 (i) entail that $\omega(S)=\{\mu_{0}\}$ is the global attractor of $P^{*}|_{S}$, where $P^{*}|_{S}$ means the semiflow $P^{*}$ restricted to $S$.
 The invariance of $\omega(B)$ and $\omega(B)\subset S$ entail that
@@
 with lower bound $\nu^{-1,0}_{t_{-1}}\in S$ and upper bound $\nu^{1,0}_{t_{1}}\in S$, the proof of (iii) is analogous to (i). ∎
-5.4 Proof of Theorem 2.3 and Corollary 2.4
-We first prove Theorem 2.3 (i)-(iii).
-Then we use the structure theorems of global attractors (Theorem 4.2) and local attractors (Theorem 5.7) to show local convergence and attracting unbounded open sets (Theorem 2.3 (iv)(v)).
-To this end, we prepare two crucial propositions by tracing solutions backward along connecting orbits (see Proposition 5.8-5.9).
-Proof of (i)-(iii) in Theorem 2.3.
-(ii).
-By Theorem 5.4, we know that the open ball $B(\delta_{a_{k}},r_{k})$ is positively invariant under $P_{t}^{*}$, and there is some invariant measure $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$.
-Next, we show that the open balls $\{B(\delta_{a_{k}},r_{k})\}_{k=1}^{n}$ are pairwise disjoint.
-Note that (2.1) implies
-$r_{a_{i}}^{2}+r_{a_{j}}^{2}\leq 2^{-1}|a_{i}-a_{j}|^{2},\ \text{ for all }\ 1\leq i\neq j\leq n.$ (5.16)
-Suppose on the contrary that there exists $\nu\in B(\delta_{a_{i}},r_{a_{i}})\cap B(\delta_{a_{j}},r_{a_{j}})$. Then,
-$\mathcal{W}_{2}^{2}(\nu,\delta_{a_{i}})=\int_{\mathbb{R}}|x-a_{i}|^{2}\mathop{}\!\mathrm{d}\nu(x)\geq 2^{-1}|a_{i}-a_{j}|^{2}-\int_{\mathbb{R}}|x-a_{j}|^{2}\mathop{}\!\mathrm{d}\nu(x)=2^{-1}|a_{i}-a_{j}|^{2}-\mathcal{W}_{2}^{2}(\nu,\delta_{a_{j}}).$
-Hence, $2^{-1}|a_{i}-a_{j}|^{2}\leq\mathcal{W}_{2}^{2}(\nu,\delta_{a_{i}})+\mathcal{W}_{2}^{2}(\nu,\delta_{a_{j}})<r_{a_{i}}^{2}+r_{a_{j}}^{2}$, which gives a contradiction to (5.16).
-This proves (ii) except (2.3), and (2.3) will be obtained later.
-(i).
-For each $k=1,2,\dots,n-1$, notice $\frac{1}{2}(a_{k+1}-a_{k})^{2}\geq r_{k}^{2}+r_{k+1}^{2}\geq\frac{1}{2}(r_{k}+r_{k+1})^{2}$, so we have $r_{k}+r_{k+1}\leq a_{k+1}-a_{k}$.
-Together with $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$ and $\nu_{2k+1}\in B(\delta_{a_{k+1}},r_{k+1})$, we have
-$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k-1}(x)<a_{k}+r_{k}\leq a_{k+1}-r_{k+1}<\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{2k+1}(x).$
-Then Lemma 4.10 entails that $\nu_{1}<_{\mathrm{st}}\nu_{3}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$.
-In the order interval $[\nu_{2k-1},\nu_{2k+1}]$, by Proposition 5.6, at least one of the three situations happens:
-  • there is another invariant measure $\nu_{2k}$ with $\nu_{2k-1}<_{\mathrm{st}}\nu_{2k}<_{\mathrm{st}}\nu_{2k+1}$;
-  • there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$;
-  • there is a decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k-1}$.
-We want to exclude the existence of connecting orbits, and we argue by contradiction.
-Assume there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$.
-Since $\nu_{2k-1}\in B(\delta_{a_{k}},r_{k})$ and $\nu_{2k+1}\in B(\delta_{a_{k+1}},r_{k+1})$, we have $P_{t_{0}}^{*}\mu\in B(\delta_{a_{k+1}},r_{k+1})$ for some $\mu\in B(\delta_{a_{k}},r_{k})$ and some $t_{0}>0$.
-But Theorem 5.4 and disjointness in (ii) tell us $P_{t}^{*}\mu\in B(\delta_{a_{k}},r_{k})$ for all $t\geq 0$, and $B(\delta_{a_{k}},r_{k})\cap B(\delta_{a_{k+1}},r_{k+1})=\emptyset$.
-This is a contradiction, so there is no increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k+1}$.
-In the same manner, we can show there is no decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k-1}$.
-Hence, we obtain $(2n-1)$ order related invariant measures, say $\nu_{1}<_{\mathrm{st}}\nu_{2}<_{\mathrm{st}}\cdots<_{\mathrm{st}}\nu_{2n-1}$.
-Proposition 5.3 shows that there are no other invariant measures.
-(ii)-(2.3) We only need to show that $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$, and $\nu_{2k}\notin\overline{B(\delta_{a_{k+1}},r_{k+1})}$ can be proved in the same way.
-We first show that $\nu_{2k}\notin B(\delta_{a_{k}},r_{k})$.
-Assume otherwise, $\nu_{2k}\in B(\delta_{a_{k}},r_{k})$, and apply Proposition 5.6 to the order interval $[\nu_{2k},\nu_{2k+1}]$.
-Then at least one of the following statements holds:
-  • there is another invariant measure $\mu$ with $\nu_{2k}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu_{2k+1}$;
-  • there is an increasing connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$;
-  • there is a decreasing connecting orbit from $\nu_{2k+1}$ to $\nu_{2k}$.
-The statement (i) excludes the existence of $\mu$.
-Repeating the argument in the proof of (i), we can invoke Theorem 5.4 and disjointness in (ii) to show there are no connecting orbits, regardless increasing or decreasing.
-This is impossible, and thus, we must have $\nu_{2k}\notin B(\delta_{a_{k}},r_{k})$.
-By Remark 5.1, $\nu_{2k}\notin\partial\overline{B(\delta_{a_{k}},r_{k})}$, and hence, we have $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$.
-(iii).
-For each $k=1,2,\dots,n-1$, we prove there is a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$, and the existence of an increasing connecting orbit from $\nu_{2k}$ to $\mu_{2k+1}$ can be shown similarly.
-In the order interval $[\nu_{2k-1},\nu_{2k}]$, we use again Proposition 5.6.
-Then at least one of the following three is true:
-  • there is another invariant measure $\mu$ with $\nu_{2k-1}<_{\mathrm{st}}\mu<_{\mathrm{st}}\nu_{2k}$;
-  • there is an increasing connecting orbit from $\nu_{2k-1}$ to $\nu_{2k}$;
-  • there is a decreasing connecting orbit from $\nu_{2k}$ to $\nu_{2k-1}$.
-The existence of $\mu$ is excluded by the statement (i).
-The statement (ii) suggests that $\nu_{2k-1}$ has an positively invariant neighbourhood $B(\delta_{a_{k}},r_{k})$.
-By (ii)-(2.3), $\nu_{2k}\notin\overline{B(\delta_{a_{k}},r_{k})}$.
-By repeating the proof in (i), an increasing connecting orbit does not exist.
-The only option left is a decreasing connecting orbit $\{\nu_{t}^{k,\downarrow}\}_{t\in\mathbb{R}}$ from $\nu_{2k}$ to $\nu_{2k-1}$.
-∎
-Before the proof of (iv)(v) in Theorem 2.3, we need to verify a core condition in Theorem 5.7 that, $\omega(B(\delta_{a_{k}},r_{k}))$ has upper/lower bounds in connecting orbits.
-In Proposition 5.8-5.9, we will fix the connecting orbit as the laws of the solution flow of (1.1), which gives a time-inhomogeneous SDE.
-Noting that the solution of this SDE can be traced backward along the connecting orbit, and by the comparison theorem, we can find a measure in the connecting orbit at a point of time sufficiently close to $\infty$- ∞, which is a lower bound or upper bound of $\omega(B(\delta_{a_{k}},r_{k}))$ with respect to stochastic order.
+In the following, we will give key results to verify the core condition in Theorem 4.4 that, $\omega(B(\delta_{a_{k}},r_{k}))$ has upper/lower bounds in connecting orbits (Proposition 4.6 and Proposition 4.8).
+In Proposition 4.6 and Proposition 4.8, we will fix the connecting orbit as the laws of the solution flow of (1.1), which gives a time-inhomogeneous SDE.
+Noting that the solution of this SDE can be traced backward along the connecting orbit, and by the comparison theorem, we can find a measure in the connecting orbit at a point of time sufficiently close to -$\infty$, which is a lower bound or upper bound of $\omega(B(\delta_{a_{k}},r_{k}))$ with respect to stochastic order.
+Firstly, we need to present a comparison principle for extrinsic-law-dependent SDEs (3.4).
+Lemma 4.5.
+Suppose Assumption 1 holds. Let $\mu,\nu\in C([s,\infty);\mathcal{P}_{2}(\mathbb{R}))$ and $T>0$. Assume
+$\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{t}(x)\leq\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{t}(x),\ \text{ for all }\ t\in[s,s+T].$ (4.4)
+If $\xi,\eta\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ and $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, then $\mathcal{L}(X_{t}^{s,\mu,\xi})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{s,\nu,\eta})$ for all $t\in[s,s+T]$.
+Proof.
+Since $\mathcal{L}(\xi)\leq_{\mathrm{st}}\mathcal{L}(\eta)$, Strassen’s theorem (see Lindvall [34, Theorem IV.2.4] or [35, equation (3)]) yields that there exist $\tilde{\xi},\tilde{\eta}\in L^{2}(\Omega,\mathcal{F}_{s},\mathbf{P})$ such that
+$\mathcal{L}(\tilde{\xi})=\mathcal{L}(\xi),\ \ \mathcal{L}(\tilde{\eta})=\mathcal{L}(\eta),\ \ \text{ and }\ \ \tilde{\xi}\leq\tilde{\eta},\ \text{ $\mathbf{P}$-a.s.}$ (4.5)
+[…]
+Note that
+$\mathop{}\!\mathrm{d}X^{s,\mu,\tilde{\xi}}_{t}$ $=\Big[-V^{\prime}(X^{s,\mu,\tilde{\xi}}_{t})-\theta X^{s,\mu,\tilde{\xi}}_{t}+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu_{t}(x)\Big]\mathop{}\!\mathrm{d}t+\sigma(X^{s,\mu,\tilde{\xi}}_{t})\mathop{}\!\mathrm{d}B_{t};$ (4.6)
+$\mathop{}\!\mathrm{d}Y^{s,\nu,\tilde{\eta}}_{t}$ $=\Big[-V^{\prime}(Y^{s,\nu,\tilde{\eta}}_{t})-\theta Y^{s,\nu,\tilde{\eta}}_{t}+\theta\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu_{t}(x)\Big]\mathop{}\!\mathrm{d}t+\sigma(Y^{s,\nu,\tilde{\eta}}_{t})\mathop{}\!\mathrm{d}B_{t}.$
+Since $\theta>0$, according to the comparison theorem for usual SDEs [20, Theorem 1.2], we have
+$X_{t}^{s,\mu,\tilde{\xi}}\leq Y^{s,\nu,\tilde{\eta}}_{t}\ \text{ for all }\ t\in[s,s+T],\text{ $\mathbf{P}$-a.s.}$
+The desired result is proved. ∎
 Proposition 4.6.
@@
 $\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu(x)<\inf_{\mu\in B}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x),\ \ \Big(resp.\ \int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\nu(x)>\sup_{\mu\in B}\int_{\mathbb{R}}x\mathop{}\!\mathrm{d}\mu(x)\Big),$ (4.7)
-then there exists $t_{0}\in\mathbb{R}$ such that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ ((((resp. $\tilde{\nu}_{t}\geq_{\mathrm{st}}\omega(B)$)))) for all $t\leq t_{0}$.
+then there exists $t_{0}\in\mathbb{R}$ such that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ ((resp. $\tilde{\nu}_{t}\geq_{\mathrm{st}}\omega(B)$)) for all $t\leq t_{0}$.
 Proof.
-We first prove the “$<$” case in (5.17), and the case “$>$” can be proved similarly. Now consider the following time-inhomogeneous SDE
+We first prove the “<<” case in (4.7), and the case “>>” can be proved similarly. Now consider the following time-inhomogeneous SDE
 $\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\tilde{\nu}_{t})(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (4.8)
@@
 Since $\{\tilde{\nu}_{t}\}_{t\in\mathbb{R}}$ is a connecting orbit, then Lemma 3.7 (i) yields that it is an entrance measure of equation (4.8), i.e., $P^{*}(t,s)\tilde{\nu}_{s}=\tilde{\nu}_{t}$ for all $t\geq s$.
-Note that $\sup_{t\in\mathbb{R}}\|\tilde{\nu}_{t}\|_{2}<\infty$ (by definition, the connecting orbit is bounded in $\mathcal{P}_{2}(\mathbb{R})$), by Lemma 4.9, we know that there exist $C>0,\lambda>0$ such that for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$ and $t\geq s$,
+Note that $\sup_{t\in\mathbb{R}}\|\tilde{\nu}_{t}\|_{2}<\infty$ (by definition, the connecting orbit is bounded in $\mathcal{P}_{2}(\mathbb{R})$), by Lemma 3.7, we have there exist $C>0,\lambda>0$ such that for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$ and $t\geq s$,
 $d_{2}\big(\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big)\leq C(1+\|\mu\|_{2}^{2})e^{-\lambda(t-s)},$ (4.9)
-where $d_{2}$ is given as in (5.15). Note that
-$\Big{|}\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\mathcal{L}(X_{t}^{s,\mu})(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)\Big{|}\leq|f|_{\infty}d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big{)},\ \text{ for all }\ f\in C_{b}(\mathbb{R}),$ (5.20)
-and
-$\Big{|}\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\mathcal{L}(X_{t}^{s,\mu})(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\tilde{\nu}_{t}(x)\Big{|}\leq d_{2}\big{(}\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big{)}.$ (5.21)
-By (5.19) and (5.20), we conclude that $\mathcal{L}(X_{t}^{s,\mu_{s}})\to\tilde{\nu}_{t}$ weakly as $s\to-\infty$ for all bounded sets $\{\mu_{s}\}_{s\leq t}$ in $\mathcal{P}_{2}(\mathbb{R})$.
-Then (5.19), (5.21) together with [68, Theorem 6.9] entail that, for all $t\in\mathbb{R}$,
+where $d_{2}$ is given by
+$d_{2}(\nu_{1},\nu_{2})=\int_{\mathbb{R}}(1+|x|^{2})\mathop{}\!\mathrm{d}|\nu_{1}-\nu_{2}|(x).$ (4.10)
+According to [54, Theorem 6.15], we conclude that
+$\mathcal{W}_{2}\big(\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big)\leq\sqrt{2d_{2}\big(\mathcal{L}(X_{t}^{s,\mu}),\tilde{\nu}_{t}\big)}\leq\sqrt{2C(1+\|\mu\|_{2}^{2})e^{-\lambda(t-s)}},\ \text{ for all }\ \mu\in\mathcal{P}_{2}(\mathbb{R}),\ t\geq s.$
+Hence, for all bounded sets $\{\mu_{s}\}_{s\leq t}$ in $\mathcal{P}_{2}(\mathbb{R})$, we have
 $\lim_{s\to-\infty}\mathcal{W}_{2}\big(\mathcal{L}(X_{t}^{s,\mu_{s}}),\tilde{\nu}_{t}\big)=0,\ \text{ for any bounded set $\{\mu_{s}\}_{s\leq t}$ in $\mathcal{P}_{2}(\mathbb{R})$}.$ (4.11)
@@
 Next, we prove that $\tilde{\nu}_{t}\leq_{\mathrm{st}}\omega(B)$ for all $t\leq t_{0}$.
-By Theorem 2.1 (ii), the order interval $[\underline{\nu},\overline{\nu}]$ attracts the bounded set $B$.
+By Theorem 1.2 (ii), the order interval $\underline{\nu},\overline{\nu}]$ attracts the bounded set $B$.
 Then Proposition 2.4 and Lemma 3.1 (iii) tell us $\omega(B)$ is nonempty.
@@
 $\mathop{}\!\mathrm{d}Y_{s}$ $=-V^{\prime}(Y_{s})\mathop{}\!\mathrm{d}t-(W^{\prime}*P^{*}_{s-t+t_{k}}\rho_{k})(Y_{s})\mathop{}\!\mathrm{d}s+\sigma(Y_{s})\mathop{}\!\mathrm{d}B_{s},$ $\mathcal{L}(Y_{t-t_{k}})$ $=\rho_{k}.$
-Note that $\rho_{k}\in B$ and $B$ is positively invariant, then Assumption 1 (iv), (5.23) and Theorem 4.3 yield
+Note that $\rho_{k}\in B$ and $B$ is positively invariant, then Assumption 1 (iv), (4.12) and Lemma 4.5 yield
 $\mathcal{L}(X_{t}^{t-t_{k},\rho_{k}})\leq_{\mathrm{st}}\mathcal{L}(Y_{t}^{t-t_{k},\rho_{k}})=P^{*}_{t_{k}}\rho_{k}.$
@@
 ∎
+Now for any fixed $\mu\in\mathcal{P}_{2}(\mathbb{R})$, we consider the following SDE:
+$\mathop{}\!\mathrm{d}X_{t}=-V^{\prime}(X_{t})\mathop{}\!\mathrm{d}t-(W^{\prime}*\mu)(X_{t})\mathop{}\!\mathrm{d}t+\sigma(X_{t})\mathop{}\!\mathrm{d}B_{t}.$ (4.14)
+Under Assumption 1, we have SDE (4.14) has a unique invariant measure $\rho^{\mu}\in\mathcal{P}_{2}(\mathbb{R})$ (see e.g., Meyn-Tweedie [41, 42] or Lemma 3.7 (ii)).
+Define a measure-iterating map $\Psi:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ by
+[…]
+According to [36, (67)], we have the following property for $\Psi$.
+Lemma 4.7.
+Suppose that Assumption 1 holds.
+If the equation (1.1) is locally dissipative at $a\in\mathbb{R}$ with configuration $(r_{a},g_{a})$, then we have $\Psi\big(\overline{B(\delta_{a},r_{a})}\big)\subset B(\delta_{a},r_{a})$.
+Remark 4.1.
+Note that $\mu$ is an invariant measure of the equation (1.1) if and only if $\Psi(\mu)=\mu$.
+Then Lemma 4.7 implies that there is no invariant measure in $\partial\overline{B(\delta_{a},r_{a})}:=\overline{B(\delta_{a},r_{a})}\setminus B(\delta_{a},r_{a})$.
 Proposition 4.8.
 Assume all assumptions in Theorem 1.4 hold.
-Let $\{\nu^{k,\uparrow}_{t}\}_{t\in\mathbb{R}}$ ((((resp.
-$\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$)))) be the increasing ((((resp.
-decreasing)))) connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$ ((((resp.
-$\nu_{2k-1}$)))) in Theorem 2.3 (iii).
+Let $\{\nu^{k,\uparrow}_{t}\}_{t\in\mathbb{R}}$ ((resp.
+$\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$)) be the increasing ((resp.
+decreasing)) connecting orbit from $\nu_{2k}$ to $\nu_{2k+1}$ ((resp.
+$\nu_{2k-1}$)) in Theorem 1.4 (iii).
 Then there exists $t_{k}\in\mathbb{R}$ such that, for all $t\leq t_{k}$,
-$\nu^{k,\uparrow}_{t}\leq_{\mathrm{st}}\omega\big{(}B(\delta_{a_{k+1}},r_{k+1})\big{)},\ \text{ {$\big{(}$}resp.
-$\nu^{k,\downarrow}_{t}\geq_{\mathrm{st}}\omega\big{(}B(\delta_{a_{k}},r_{k})\big{)}${$\big{)}$}}.$
+$\nu^{k,\uparrow}_{t}\leq_{\mathrm{st}}\omega\big(B(\delta_{a_{k+1}},r_{k+1})\big),\ \text{ {$\big($}resp. $\nu^{k,\downarrow}_{t}\geq_{\mathrm{st}}\omega\big(B(\delta_{a_{k}},r_{k})\big)${$\big)$}}.$
 Proof.
@@
 there exists $\mu\in\overline{B(\delta_{a_{k+1}},r_{k+1})}$ such that $\int x\mathop{}\!\mathrm{d}\mu(x)=m_{2k}$.
-By the definition of $\Psi$ (see (5.13)) and Theorem 5.5, it follows that $\Psi(\mu)=\mu_{m_{2k}}=\nu_{2k}\in\Psi(\overline{B(\delta_{a_{k+1}},r_{k+1})})\subset B(\delta_{a_{k+1}},r_{k+1})$.
+By the definition of $\Psi$ (see (4.15)) and Lemma 4.7, it follows that $\Psi(\mu)=\mu_{m_{2k}}=\nu_{2k}\in\Psi(\overline{B(\delta_{a_{k+1}},r_{k+1})})\subset B(\delta_{a_{k+1}},r_{k+1})$.
 This gives a contradiction to (1.7).
@@
 This finishes the proof. ∎
+4.3 Proof of Main Theorems
+We have obtained Theorem 1.2 and Corollary 1.3 in Section 3.3, and proved (i)–(iii) of Theorem 1.4 in Section 4.2.
+In this section, we complete the proof of the remaining main theorems and conclude with a conjecture on a saddle-point structure in the double-well granular media equation.
 Proof of (iv)(v) in Theorem 1.4.
@@
 By Theorem 1.4 (i)(ii) and Remark 4.1, there is only one invariant measure $\nu_{1}$ in $\overline{B(\delta_{a_{1}},r_{1})}$.
-Theorem 2.1 (ii) and Theorem 2.3 (i) entail that $[\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{1}},r_{1})$ and hence $\omega(B(\delta_{a_{1}},r_{1}))\subset[\nu_{1},\nu_{2n-1}]$.
+Theorem 1.2 (ii) and Theorem 1.4 (i) entail that $\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{1}},r_{1})$ and hence $\omega(B(\delta_{a_{1}},r_{1}))\subset[\nu_{1},\nu_{2n-1}]$.
 Then $\nu_{1}\leq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$.
@@
 Proposition 4.8 shows that there exists $t_{1}\in\mathbb{R}$ such that $\nu_{t_{1}}^{1,\downarrow}\geq_{\mathrm{st}}\omega(B(\delta_{a_{1}},r_{1}))$.
+Note that, by [36, Theorem 3.5], the semigroup $P_{t}^{*}:\mathcal{P}_{2}(\mathbb{R})\to\mathcal{P}_{2}(\mathbb{R})$ generated by the equation (1.1) is an order-preserving semiflow.
 Then Theorem 4.4 (ii) implies that $\nu_{1}$ attracts $B(\delta_{a_{1}},r_{1})$ under $P^{*}$, which means $\sup_{\mu\in B(\delta_{a_{k}},r_{k})}\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$ as $t\to\infty$.
@@
 By Theorem 1.4 (i)(ii) and Remark 4.1, there is only one invariant measure $\nu_{2k-1}$ in $\overline{B(\delta_{a_{k}},r_{k})}$.
-We also see the order interval $[\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{k}},r_{k})$ by Theorem 2.1 (ii) and Theorem 2.3 (i).
+We also see the order interval $\nu_{1},\nu_{2n-1}]$ attracts $B(\delta_{a_{k}},r_{k})$ by Theorem 1.2 (ii) and Theorem 1.4 (i).
 By Proposition 4.8, we can find some $t_{k}\in\mathbb{R}$ such that $\nu_{t_{k}}^{k,\uparrow}\leq_{\mathrm{st}}\omega(B(\delta_{a_{k}},r_{k}))\leq_{\mathrm{st}}\nu_{t_{k}}^{k,\downarrow}$.
@@
 (v).
-Noticing the conditions in Theorem 4.2 are verified in the proof of Theorem 2.1, we apply Theorem 4.2 (iv) to finish the proof.
-If $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r_{1})$, then from Theorem 4.2 (iv), we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$.
+Noticing the conditions in Proposition 3.2 are verified in the proof of Theorem 1.2, we apply Proposition 3.2 (iv) to finish the proof.
+If $\mu\leq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{1}},r_{1})$, then from Proposition 3.2 (iv), we have $\mathcal{W}_{2}(P_{t}^{*}\mu,\nu_{1})\to 0$.
 If $\mu\geq_{\mathrm{st}}\nu$ for some $\nu\in B(\delta_{a_{n}},r_{n})$, the convergence result follows as well.
@@
 $\lim_{s\to-\infty}\mathcal{W}_{2}\big(\mathcal{L}(X_{t_{0}}^{m;s,0}),\mu_{m}\big)=0,\ \text{ and }\ \lim_{s\to-\infty}\mathcal{W}_{2}\big(\mathcal{L}(Y^{s,0}_{t_{0}}),\nu^{1,\downarrow}_{t_{0}}\big)=0.$ (4.19)
-On the other hand, Theorem 4.3 yields
+On the other hand, Lemma 4.5 yields
 $\mathcal{L}(X_{t_{0}}^{m;s,0})\leq_{\mathrm{st}}\mathcal{L}(Y^{s,0}_{t_{0}}),\ \text{ for all }\ s\leq t_{0}.$
@@
 $\mathcal{W}_{2}(P_{t}^{*}\nu^{1,\downarrow}_{t_{0}},\nu_{1})=\mathcal{W}_{2}(\nu^{1,\downarrow}_{t_{0}+t},\nu_{1})\to 0,\ \text{ as }\ t\to\infty,$
-then Theorem 4.2 (v) shows that $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{1})\to 0$ as $t\to\infty$.
+then Proposition 3.2 (v) shows that $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{1})\to 0$ as $t\to\infty$.
 Case $1<k<n$: Let $\{\nu^{k,\downarrow}_{t}\}_{t\in\mathbb{R}}$ and $\{\nu^{k-1,\uparrow}_{t}\}_{t\in\mathbb{R}}$ be the corresponding decreasing and increasing connecting orbits from $\nu_{2k}=\mu_{m_{2k}}$ to $\nu_{2k-1}$, and $\nu_{2k-2}=\mu_{m_{2k-2}}$ to $\nu_{2k-1}$ respectively in Theorem 1.4 (iii).
@@
 Similar to Case $k=1$, there exists $t_{0}\in\mathbb{R}$ such that $\mu_{m}\geq_{\mathrm{st}}\nu^{n-1,\uparrow}_{t_{0}}$.
-Then applying Theorem 4.2 (v) again, we conclude $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{2n-1})\to 0$ as $t\to\infty$.
+Then applying Proposition 3.2 (v) again, we conclude $\mathcal{W}_{2}(P^{*}_{t}\mu_{m},\nu_{2n-1})\to 0$ as $t\to\infty$.
 ∎
-6 Applications to Classical Examples: Proof of Theorem 2.5-2.9
+Before proving Theorem 1.1, we first give the following key proposition to verify the local dissipativity condition.
+Proposition 4.9.
+Assume that $W(x)=\frac{\theta}{2}x^{2}$ and the confining potential function $V\in C^{2}(\mathbb{R};\mathbb{R})$ satisfy the following conditions:
+  1. (i)
+    There exists $a\in\mathbb{R}$ such that $V^{\prime}(a)=0$ and $V^{\prime\prime}(a)>0$;
+  2. (ii)
+    $\liminf_{|x|\to\infty}\frac{V^{\prime}(x)}{x|x|^{\delta}}>0$ for some $\delta>0$.
+Then, there exists $r_{a}>0$ such that for any $0<r\leq r_{a}$, there are $\theta_{r}>0$ and $\sigma_{r}>0$ such that for all $\theta\geq\theta_{r}$ and $\sigma^{2}(x)\leq\sigma_{r}$ for all $x\in\mathbb{R}$, the equation (1.1) is locally dissipative at $a$ with radius $r$.
+Proof.
+Without loss of generality, we assume $0<\delta<1$. Then we first show that there exist $C_{1},C_{2},C_{3}>0$ such that
+$\frac{V^{\prime}(x+a)}{x}\geq C_{1}|x|^{\delta}-C_{2}|x|^{\frac{\delta}{2}}+C_{3},\ \ \text{for any $x\neq 0$}.$ (4.20)
+By (ii), it holds that $\liminf_{|x|\to\infty}\frac{V^{\prime}(x+a)}{x|x|^{\delta}}>0$. Thus, there exists $R>0$ and $C_{1}>0$ such that
+$\frac{V^{\prime}(x+a)}{x}\geq C_{1}|x|^{\delta},\ \text{ for all }\ |x|\geq R.$
+Let $C_{3}=\frac{V^{\prime\prime}(a)}{2}>0$. Notice that $\lim_{x\to 0}\frac{V^{\prime}(x+a)}{x}=V^{\prime\prime}(a)>0$, there exist $0<\epsilon_{a}<R$ and such that
+$\frac{V^{\prime}(x+a)}{x}\geq C_{3},\ \text{ for all }\ |x|\leq\epsilon_{a}.$
+Take
+$C_{2}=\epsilon_{a}^{-\frac{\delta}{2}}\bigg(\sup_{\epsilon_{a}\leq|x|\leq R}\bigg|\frac{V^{\prime}(x+a)}{x}\bigg|+C_{1}R^{\delta}+C_{3}\bigg).$
+One can verify (4.20) in three cases: $\left\lvert x\right\rvert\geq R$, $\epsilon_{a}<|x|<R$ and $|x|\leq\epsilon_{a}$, separately. Then by (4.20) we have
+$2xV^{\prime}(x+a)=-2|x|^{2}\frac{V^{\prime}(x+a)}{x}\leq-2C_{1}|x|^{2+\delta}+2C_{2}|x|^{2+\frac{\delta}{2}}-2C_{3}|x|^{2},\ \text{ for all }\ x\in\mathbb{R}.$ (4.21)
+Let $\overline{\sigma}:=\sup_{x\in\mathbb{R}}|\sigma(x)|$. Choose
+$g_{a}(z,w)=2C_{1}z^{1+\frac{\delta}{2}}-2C_{2}z^{1+\frac{\delta}{4}}+(2C_{3}+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2}.$
+Obviously, $g_{a}$ is continuous and $g_{a}(z,\cdot)$ is decreasing for any $z\geq 0$ and hence
+$\inf_{0\leq w\leq r^{2}}g_{a}(z,w)=g_{a}(z,r^{2}),\ \text{ for any }\ r>0.$ (4.22)
+Also, (4.21) implies that
+$2xV^{\prime}(x+a)-2x(W^{\prime}*\mu)(x)+|\sigma(x+a)|^{2}\leq-g_{a}(|x|^{2},\|\mu\|_{2}^{2}).$ (4.23)
+For any $r>0$, a direct computation yields
+$g_{a}(r^{2},r^{2})=2r^{2}\big(C_{1}r^{\delta}-C_{2}r^{\frac{\delta}{2}}+C_{3}\big)-\overline{\sigma}^{2},$
+$\frac{\partial g_{a}(z,r^{2})}{\partial z}\bigg|_{z=r^{2}}=2\big(1+\frac{\delta}{2}\big)C_{1}r^{\delta}-2\big(1+\frac{\delta}{4}\big)C_{2}r^{\frac{\delta}{2}}+2C_{3}+\theta,$
+and for all $z\geq 0$,
+$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}=\delta\big(1+\frac{\delta}{2}\big)C_{1}z^{-(1-\frac{\delta}{2})}-\frac{\delta}{2}\big(1+\frac{\delta}{4}\big)C_{2}z^{-(1-\frac{\delta}{4})}+\frac{\theta r}{2}z^{-\frac{3}{2}}.$ (4.24)
+Then there exists $r_{a}>0$ such that for any $0<r\leq r_{a}$, we have
+$2r^{2}\big(C_{1}r^{\delta}-C_{2}r^{\frac{\delta}{2}}+C_{3}\big)>0,$ (4.25)
+and
+$\frac{\partial g_{a}(z,r^{2})}{\partial z}\bigg|_{z=r^{2}}=2\big(1+\frac{\delta}{2}\big)C_{1}r^{\delta}-2\big(1+\frac{\delta}{4}\big)C_{2}r^{\frac{\delta}{2}}+2C_{3}+\theta>0.$ (4.26)
+Now, we fix $0<r\leq r_{a}$. By (4.24), we have $\lim_{z\downarrow 0}\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}=\infty$. Hence, there exists $\epsilon_{r}>0$ such that
+$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}>0,\ \text{ for all }\ 0<z<\epsilon_{r}.$
+On the other hand,
+$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}=\frac{\delta}{8}z^{(\frac{\delta}{4}-1)}\big(4(2+\delta)C_{1}z^{\frac{\delta}{4}}-(4+\delta)C_{2}\big)+\frac{\theta r}{2}z^{-\frac{3}{2}}>0,\ \text{ for all }\ z\geq C_{\delta}:=\Big(\frac{(4+\delta)C_{2}}{4(2+\delta)C_{1}}\Big)^{\frac{4}{\delta}}.$
+Let
+$\theta_{r}=\frac{\delta}{r}\big(1+\frac{\delta}{4}\big)C_{2}C_{\delta}^{\frac{3}{2}}\epsilon_{r}^{-(1-\frac{\delta}{4})}.$ (4.27)
+For $\theta\geq\theta_{r}$, by (4.24), one has
+$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}\geq\delta\big(1+\frac{\delta}{2}\big)C_{1}z^{-(1-\frac{\delta}{2})}>0,\ \text{ for all }\ \epsilon_{r}\leq z\leq C_{\delta}.$
+Thus, for all $\theta\geq\theta_{r}$, we conclude that
+$\frac{\partial^{2}g_{a}(z,r^{2})}{\partial z^{2}}>0,\ \text{ for all }\ z>0.$ (4.28)
+So $g_{a}(\cdot,r^{2})$ is convex. Thus, by (4.26),
+$\frac{\partial g_{a}(z,r^{2})}{\partial z}\geq\frac{\partial g_{a}(r^{2},r^{2})}{\partial z}>0,\ \text{ for all }\ z\geq r^{2}.$ (4.29)
+(4.25) implies that there exists $\sigma_{r}$ such that
+$0<\sigma_{r}<2r^{2}\big(C_{1}r^{\delta}-C_{2}r^{\frac{\delta}{2}}+C_{3}\big).$ (4.30)
+Then by (4.29), we have
+$g_{a}(z,r^{2})\geq g_{a}(r^{2},r^{2})>\sigma_{r}-\overline{\sigma}^{2}\geq 0,\ \text{ for all }\ z\geq r^{2}.$ (4.31)
+Therefore, for any $0<r\leq r_{a}$, let $\theta_{r},\sigma_{r}$ be given as in (4.27) and (4.30).
+It follows from (4.22), (4.23), (4.28), (4.31) that for all $\theta\geq\theta_{r},\ \overline{\sigma}^{2}\leq\sigma_{r}$, equation (1.1) is locally dissipative at $a$ with radius $r$.
+∎
+Remark 4.2.
+If $V^{\prime}$ is hyper-dissipative (see Assumption 1 (iii)), i.e., there exist $\alpha>0$, $\beta>0$, $\delta>0$ such that, for all $x,y\in\mathbb{R}$,
+$(x-y)(V^{\prime}(x)-V^{\prime}(y))\leq-\alpha\left\lvert x-y\right\rvert^{2+\delta}+\beta.$
+Then we have
+$xV^{\prime}(x)\leq-\alpha\left\lvert x\right\rvert^{2+\delta}+|x||V^{\prime}(0)|+\beta,$
+and hence
+$\frac{V^{\prime}(x)}{x|x|^{\delta}}=\frac{xV^{\prime}(x)}{|x|^{2+\delta}}\geq\alpha-\frac{|V^{\prime}(0)|}{|x|^{1+\delta}}-\frac{\beta}{|x|^{2+\delta}}.$
+Thus, $\lim_{|x|\to\infty}\frac{V^{\prime}(x)}{x|x|^{\delta}}=\alpha>0$ and the condition (ii) in Proposition 4.9 holds true.
 Proof of Theorem 1.1.
-According to Theorem 2.3, it suffices to show that equation (2.4) are locally dissipative at $-1,1$ with configurations $(r,g_{-1}),(r,g_{1})$ for $r=\frac{9-\sqrt{17}}{8}$ and for some $g_{\pm 1}$.
+According to Remark 4.2, we have
+$\lim_{x\to\infty}V^{\prime}(x)=\infty,\ \ \lim_{x\to-\infty}V^{\prime}(x)=-\infty.$ (4.32)
+Since $V^{\prime}$ has only finitely many zeros and each of them is simple, the number of zeros must be odd.
+Denoting them by $a_{1}<b_{1}<a_{2}<b_{2}<\cdots<a_{n-1}<b_{n-1}<a_{n}$, it then follows that
+$V^{\prime\prime}(a_{i})>0,\ 1\leq i\leq n,\ \text{ and }\ V^{\prime\prime}(b_{j})<0,\ 1\leq j\leq n-1.$ (4.33)
+By virtue of Proposition 4.9, Remark 4.2 and (4.33) that, there exist
+$0<r<\frac{1}{2}\min_{1\leq i\leq n-1}\left\lvert a_{i+1}-a_{i}\right\rvert,\ \ \theta_{0}>0,\ \text{ and }\ \sigma_{0}>0$
+such that, for all $\theta\geq\theta_{0}$ and $\overline{\sigma}\leq\sigma_{0}$, equation (1.1) is locally dissipative at $a_{k}$ with radius $r$ for all $1\leq k\leq n$.
+Then the desired result follows from Theorem 1.4.
+∎
+Proof of Theorem 1.6.
+According to Theorem 1.4, it suffices to show that equation (1.9) are locally dissipative at 1,1-1,1 with configurations $(r,g_{-1}),(r,g_{1})$ for $r=\frac{9-\sqrt{17}}{8}$ and for some $g_{\pm 1}$.
 Note that
@@
 Construction of $g_{1}$: For any $x\in\mathbb{R},\mu\in\mathcal{P}_{2}(\mathbb{R})$, by a simple calculation, we have
-$-2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big{(}2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}\big{)},$
+$2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big(2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}\big),$
 where
@@
 $\frac{\partial^{2}g_{1}(z,w)}{\partial z^{2}}=4-\frac{9}{2}z^{-\frac{1}{2}}+\frac{1}{2}\theta w^{\frac{1}{2}}z^{-\frac{3}{2}}.$
-Then by calculation, we know that for any $w>0$, $g_{1}(\cdot,w)$ is convex if and only if
+Then by calculation, we have for any $w>0$, $g_{1}(\cdot,w)$ is convex if and only if
 $\theta\geq\frac{27}{16w^{\frac{1}{2}}}.$ (4.35)
-Note that for any $10<r<$ 1,
+Note that for any $0<r<1$,
 $g_{1}(r^{2},r^{2})=2r^{2}(r-1)(r-2)-\overline{\sigma}^{2},$
@@
 $\frac{\partial g_{1}}{\partial z}(r^{2},r^{2})=4r^{2}-9r+4+\theta>\theta-1.$
-Now for any fixed $10<r<$ 1, if we choose
+Now for any fixed $0<r<1$, if we choose
 $\theta\geq\frac{27}{16r}>1,\ \text{ and }\ \overline{\sigma}^{2}<2r^{2}(r-1)(r-2),$
@@
 $g_{1}(r^{2},r^{2})>0,\ \text{ and }\ \frac{\partial g_{1}}{\partial z}(r^{2},r^{2})>0.$
-By (6.2), we know that $g_{1}(\cdot,r^{2})$ is convex. Hence,
+By (4.35), we have $g_{1}(\cdot,r^{2})$ is convex. Hence,
 $\frac{\partial g_{1}}{\partial z}(z,r^{2})\geq\frac{\partial g_{1}}{\partial z}(r^{2},r^{2})>0\ \text{ for all }\ z\geq r^{2},$
@@
 $g_{1}(z,r^{2})\geq g_{1}(r^{2},r^{2})>0\ \text{ for all }\ z\geq r^{2}.$
-Hence, for any $10<r<$ 1 and choosing $g_{1}$ as in (6.1), equation (2.4) is locally dissipative at $1$ with configuration $(r,g_{1})$ whenever
+Hence, for any $0<r<1$ and choosing $g_{1}$ as in (4.34), equation (1.9) is locally dissipative at $1$ with configuration $(r,g_{1})$ whenever
 $\theta\geq\frac{27}{16r},\ \text{ and }\ \overline{\sigma}^{2}<2r^{2}(r-1)(r-2).$ (4.36)
@@
 Combining (4.36)(4.37), under the condition (1.10) in Theorem 1.6, equation (1.9) is locally dissipative at $1$ with configuration $(r,g_{1})$ where $r=\frac{9-\sqrt{17}}{8}$ and $g_{1}$ is given as in (4.34).
-Construction of $g_{-1}$: Similarly, let $g_{-1}=g_{1}$ be given as in (6.1), we can show that equation (2.4) is also locally dissipative at $-1$ with configuration $\big{(}\frac{9-\sqrt{17}}{8},g_{-1}\big{)}$.
-∎
-Proof of Theorem 2.6.
-By a simple calculation, we know that
-$-2xV^{\prime}(x+1)\leq-2x^{2}(x+1)(x+2)+\frac{1}{4}.$
-Then we have
-$-2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big{(}2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}-\frac{1}{4}\big{)}.$
-Choose $g_{1}$ as follows: for any $z,w\geq 0$,
-$g_{1}(z,w)=2z^{2}-6z^{\frac{3}{2}}+(4+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-|\overline{\sigma}|^{2}-\frac{1}{4}.$ (6.5)
-Repeating the proof of that in Theorem 2.5, we know that the equation (2.6) is locally dissipative at $1$ with configuration $(r,g_{1})$ with $r=\frac{9-\sqrt{17}}{8}$ whenever
-$\theta\geq\frac{27(9+\sqrt{17})}{128},\ \text{ and }\ \overline{\sigma}^{2}+\frac{1}{4}<\frac{51\sqrt{17}-107}{256},$
-i.e., (2.7) holds.
-Similarly, choose $g_{-1}=g_{1}$ as in (6.5), the equation (2.6) is locally dissipative at $-1$ with configuration $\big{(}\frac{9-\sqrt{17}}{8},g_{-1}\big{)}$.
-Then the desired results follow from Theorem 2.3 directly.
+Construction of $g_{-1}$: Similarly, let $g_{-1}=g_{1}$ be given as in (4.34), we can show that equation (1.9) is also locally dissipative at 1-1 with configuration $\big(\frac{9-\sqrt{17}}{8},g_{-1}\big)$.
 ∎
@@
 Take
-$g_{1}(z,w)$ $=2z^{2}-8z^{\frac{3}{2}}+(6+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{1}$ $=\frac{6-3\sqrt{3}}{2},$
-$g_{-2}(z,w)$ $=2z^{2}-10z^{\frac{3}{2}}+(12+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{-2}$ $=\frac{(\sqrt{33}-15)^{2}}{64}.$
-Through calculations similar in Theorem 2.5, local dissipative conditions hold at $1,-2$ for the equation (2.8). We apply Theorem 2.3 to get the desired results. ∎
+$g_{0}(z,w)$ $=2z^{3}-10z^{2}+(8+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{0}$ $=\frac{\sqrt{15-3\sqrt{13}}}{3},$
+$g_{2}(z,w)$ $=2z^{3}-20z^{\frac{5}{2}}+70z^{2}-100z^{\frac{3}{2}}+(48+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{2}$ $=r_{0},$
+$g_{-2}(z,w)$ $=g_{2}(z,w),$ $r_{-2}$ $=r_{0},$
+where $\overline{\sigma}:=\sup_{x\in\mathbb{R}}|\sigma(x)|$.
+Similar to the calculation in the proof of Theorem 1.6, we can prove that the equation (1.11) is locally dissipative at $0,\pm 2$.
+The proof is finished by applying Theorem 1.4.
+∎
 Proof of Theorem 1.8.
-Take
-$g_{0}(z,w)$ $=2z^{3}-10z^{2}+(8+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{0}$ $=\frac{5-\sqrt{13}}{3},$
-$g_{2}(z,w)$ $=2z^{3}-20z^{\frac{5}{2}}+70z^{2}-100z^{\frac{3}{2}}+(48+2\beta)z-2\beta z^{\frac{1}{2}}w^{\frac{1}{2}}-\overline{\sigma}^{2},$ $r_{2}$ $=r_{0},$
-$g_{-2}(z,w)$ $=g_{2}(z,w),$ $r_{-2}$ $=r_{0}.$
-Then the equation (2.9) is locally dissipative at $0,\pm 2$. The proof is finished by applying Theorem 2.3. ∎
-Proof of Theorem 2.9.
-The conclusions directly follow from Theorem 2.1. ∎
+By a simple calculation, we have
+$2xV^{\prime}(x+1)\leq-2x^{2}(x+1)(x+2)+\frac{1}{4}.$
+Then we have
+$2xV^{\prime}(x+1)-2x(W^{\prime}*\mu)(x)+|\sigma(x+1)|^{2}\leq-\big(2|x|^{4}-6|x|^{3}+(4+2\theta)|x|^{2}-2\theta|x|\|\mu\|_{2}-\overline{\sigma}^{2}-\frac{1}{4}\big),$
+where $\overline{\sigma}:=\sup_{x\in\mathbb{R}}|\sigma(x)|$. Choose $g_{1}$ as follows: for any $z,w\geq 0$,
+$g_{1}(z,w)=2z^{2}-6z^{\frac{3}{2}}+(4+2\theta)z-2\theta z^{\frac{1}{2}}w^{\frac{1}{2}}-|\overline{\sigma}|^{2}-\frac{1}{4}.$ (4.38)
+Repeating the proof of that in Theorem 1.6, we have the equation (1.12) is locally dissipative at $1$ with configuration $(r,g_{1})$ with $r=\frac{9-\sqrt{17}}{8}$ whenever
+$\theta\geq\frac{27(9+\sqrt{17})}{128},\ \text{ and }\ \overline{\sigma}^{2}+\frac{1}{4}<\frac{51\sqrt{17}-107}{256},$
+i.e., (1.13) holds.
+Similarly, choose $g_{-1}=g_{1}$ as in (4.38), the equation (1.12) is locally dissipative at 1-1 with configuration $\big(\frac{9-\sqrt{17}}{8},g_{-1}\big)$.
+Then the desired results follow from Theorem 1.4 directly.
+∎
+Motivated by our results, we conclude the paper with the following conjecture.
+In this work, by exploiting the order-preserving property of the associated semiflow, we have obtained a partial characterization of the basins of attraction of invariant measures, both in the sense of the stochastic order and the topology.
+In particular, our results reveal the existence of unbounded open subsets contained in the basins of attraction of minimal/maximum invariant measures.
+In the theory of monotone dynamical systems, a classical saddle-point structure (or saddle-type behavior) arises when a system possesses three equilibria $a<b<c$, where $a$ and $b$ are stable.
+More precisely, the state space can be decomposed into three disjoint invariant sets: the basin of attraction $B_{1}$ of $a$, the basin of attraction $B_{2}$ of $b$, and an unordered invariant set $M$ containing $c$.
+The set $M$ separates the two basins $B_{1}$ and $B_{2}$ and is usually referred to as the separatrix (see e.g., Jiang-Liang-Zhao [31]).
+Conjecture.
+In the double-well granular media equation, the three invariant measures admit an analogous saddle-point structure.
+More precisely, the unstable invariant measure $\mu_{0}$ is expected to be contained in an unordered invariant separating set that acts as a separatrix between the basins of attraction of the two stable invariant measures.
 Acknowledgement
@@
 NIF/R1/221003) and National Natural Science Foundation of China (No.
+12501184, No.
 12171280).
 Appendix A Nonexistence of Strong Ordered Pairs for Stochastic Order
-In addition to closedness of the stochastic order mentioned in Section 3.1, almost all classical results in the theory of systems that preserve a certain partial order (monotone dynamical systems) are established on the fundamental assumption that the state space $X$ is not only endowed with a partial order relation $\leq$”, but is also a “strongly ordered space” (see e.g., Hirsch [41, 42]).
+In addition to closedness of the stochastic order, almost all classical results in the theory of monotone dynamical systems are established on the fundamental assumption that the state space $X$ is not only endowed with a partial order relation $\leq$”, but is also a “strongly ordered space” (see e.g., Hirsch [27, 28]).
 A strongly ordered space implies that the interior of the partial order relation $\leq$” (referred to as the strong ordering $\ll$”) is a nonempty relation on $X$, and its closure is $\leq$”.
-The strong ordering $\ll$” is a relation for which two points “$x\ll y$” means that there are open neighbourhoods $U,V$ in $X$ of $x,y$ respectively such that $U\leq V$, where the term “$U\leq V$” refers to $u\leq v$ for any $u\in U$ and $v\in V$.
+The strong ordering $\ll$” is a relation for which two points $x\ll y$” means that there are open neighbourhoods $U,V$ in $X$ of $x,y$ respectively such that $U\leq V$, where the term $U\leq V$” refers to $u\leq v$ for any $u\in U$ and $v\in V$.
 If $x\ll y$, we call $(x,y)$ is a strong ordered pair.
-Based on the tool of order topology, especially the topology generated by the interiors of order intervals enclosed by strong ordered pairs, the theory of monotone dynamical systems and its applications to cooperative irreducible ODEs (coordinate order in $\mathbb{R}^{d}$), delay equations and parabolic equations (pointwise order in functional spaces) have undergone extensively investigation (see e.g., [60, 43, 71]).
-Unfortunately, in the appendix, we will show that the strong ordering relation generated by the stochastic order “$\leq_{\mathrm{st}}$” in $\mathcal{P}_{2}(\mathbb{R})$ is empty, meaning that NO strong ordered pairs exist.
+Based on the tool of order topology, especially the topology generated by the interiors of order intervals enclosed by strong ordered pairs, the theory of monotone dynamical systems and its applications to cooperative irreducible ODEs, delay equations and parabolic equations have undergone extensively investigation (see e.g., [47, 26, 59]).
+Unfortunately, in the appendix, we will show that the strong ordering relation generated by the stochastic order $\leq_{\mathrm{st}}$” in $\mathcal{P}_{2}(\mathbb{R})$ is empty, meaning that NO strong ordered pairs exist.
 We actually prove a stronger conclusion – for any $\mu\in\mathcal{P}_{2}(\mathbb{R})$, every open set $B\subset\mathcal{P}_{2}(\mathbb{R})$ has an element that is order unrelated with $\mu$ (see Proposition A.1).
 This implies that classical theory of monotone dynamical systems are inapplicable to order-preserving semigroups on $\mathcal{P}_{2}(\mathbb{R})$.
-Consequently, exploration of properties of the stochastic order and establishment of the theory for systems preserving the stochastic order, as discussed in Section 3.2, 4.1 and 5.3, are indispensable.
+Consequently, exploration of properties of the stochastic order and the results of monotone dynamical systems, as discussed in Section 2.2, 3.1 and 4.2, are indispensable.
 Proposition A.1.
@@
 $\begin{split}&\ \ \ \ \bigg|\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{1,\epsilon}(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{0}(x)\bigg|\\ &\leq\frac{\alpha_{\epsilon}}{\nu_{0}((-c,c))}\int_{(-c,c)}|f(x)|\mathop{}\!\mathrm{d}\nu_{0}(x)+\int_{(-\infty,-c]\cup[c,\infty)}|f(x)|\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x\\ &\leq 2|f|_{\infty}\alpha_{\epsilon}\\ &\leq\frac{2|f|_{\infty}}{c^{2}}\epsilon^{2}\to 0\ \text{ as }\ \epsilon\to 0,\end{split}$ (A.2)
-[…]
+where $|f|_{\infty}=\sup_{x\in\mathbb{R}}|f(x)|$.
+Hence $\nu_{1,\epsilon}$ weakly converges to $\nu_{0}$ as $\epsilon\to 0$, which is denoted by $\nu_{1,\epsilon}\xrightarrow{w}\nu_{0}$ as $\epsilon\to 0$.
+Moreover, similar to (A.2), we have
 $\begin{split}&\ \ \ \ \lim_{\epsilon\to 0}\bigg|\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{1,\epsilon}(x)-\int_{\mathbb{R}}|x|^{2}\mathop{}\!\mathrm{d}\nu_{0}(x)\bigg|\\ &\leq\lim_{\epsilon\to 0}\Big(c^{2}\alpha_{\epsilon}+\int_{(-\infty,-c]\cup[c,\infty)}|x|^{2}\rho_{\epsilon}(x)\mathop{}\!\mathrm{d}x\Big)\\ &\leq\lim_{\epsilon\to 0}2\epsilon^{2}=0.\end{split}$ (A.3)
-[…]
-If $\mu$ and $\nu_{1}$ are order unrelated, the proof is completed by choosing $\nu=\nu_{1}$. Otherwise, without loss of generality, we assume $\mu\leq_{\mathrm{st}}\nu_{1}$. Then we know that
+By Theorem 6.9 in [54], we have $\lim_{\epsilon\to 0}\mathcal{W}_{2}(\nu_{1,\epsilon},\nu_{0})=0$ due to $\nu_{1,\varepsilon}\xrightarrow{w}\nu_{0}$ and (A.3).
+Then there exists $0<\epsilon_{1}<\epsilon_{0}$ such that $\nu_{1}:=\nu_{1,\epsilon_{1}}\in B(\nu_{0},\delta)$ satisfying (A.1).
+Thus, we have proved the Claim.
+If $\mu$ and $\nu_{1}$ are order unrelated, the proof is completed by choosing $\nu=\nu_{1}$. Otherwise, without loss of generality, we assume $\mu\leq_{\mathrm{st}}\nu_{1}$. Then we have
 $\mu((-\infty,-r])\geq\nu_{1}((-\infty,-r]),\ \ \mu([r,\infty))\leq\nu_{1}([r,\infty)),\ \text{ for all }\ r\geq c.$ (A.4)
@@
 $\mathop{}\!\mathrm{d}\nu_{2,n}(x)=\begin{cases}\mathop{}\!\mathrm{d}\nu_{1}(x),&\text{on }(-n,\infty),\\ \big(\nu_{1}((-\infty,-n])-2\mu((-\infty,-r_{n}])\big)\mathop{}\!\mathrm{d}x,&\text{on }(-(n+1),-n],\\ 0,&\text{on }(-r_{n},-(n+1)],\\ 2\mathop{}\!\mathrm{d}\mu(x),&\text{on }(-\infty,-r_{n}].\end{cases}$ (A.6)
-[…]
+It is easy to check that $\nu_{2,n}\in\mathcal{P}_{2}(\mathbb{R})$ for all $n\geq c$. Moreover, $\nu_{2,n}\xrightarrow{w}\nu_{1}$ as $n\to\infty$ since for any $f\in C_{b}(\mathbb{R})$,
 $\begin{split}&\ \ \ \ \bigg|\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{2,n}(x)-\int_{\mathbb{R}}f(x)\mathop{}\!\mathrm{d}\nu_{1}(x)\bigg|\\ &\leq\int_{(-\infty,-n]}|f(x)|\mathop{}\!\mathrm{d}\nu_{2,n}(x)+\int_{(-\infty,-n]}|f(x)|\mathop{}\!\mathrm{d}\nu_{1}(x)\\ &\leq 2|f|_{\infty}\nu_{1}((-\infty,-n])\to 0\ \text{ as }\ n\to\infty.\end{split}$
@@
 Hence, it follows from Theorem 6.9 in [54] again that $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{2,n},\nu_{1})=0.$ Thus, there exists $n_{0}\geq c$ such that $\nu:=\nu_{2,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$.
-By (A.5), the construction of $\nu_{2,n_{0}}$ in (A.6) and the choice of $r_{0}$ in Case 1, we know that
+By (A.5), the construction of $\nu_{2,n_{0}}$ in (A.6) and the choice of $r_{0}$ in Case 1, we have
 $\begin{split}0<\mu&((-\infty,-r])<2\mu((-\infty,-r])=\nu((-\infty,-r]),\text{ for all }\ r\geq r_{n_{0}},\\ &\mu([r_{0},\infty))<\nu_{1}([r_{0},\infty))=\nu([r_{0},\infty)).\end{split}$
@@
 $\mathop{}\!\mathrm{d}\nu_{3,n}(x)=\begin{cases}\mathop{}\!\mathrm{d}\nu_{1}(x),&\text{on }(-n,n),\\ \big(\nu_{1}([n,\infty))-2\mu([r_{n},\infty))\big)\mathop{}\!\mathrm{d}x,&\text{on }[n,n+1),\\ \big(\nu_{1}((-\infty,-n])-2\mu((-\infty,-r_{n}])\big)\mathop{}\!\mathrm{d}x,&\text{on }(-(n+1),-n],\\ 0,&\text{on }(-r_{n},-(n+1)]\cup[n+1,r_{n}),\\ 2\mathop{}\!\mathrm{d}\mu(x),&\text{on }(-\infty,-r_{n}]\cup[r_{n},\infty).\end{cases}$
-Similarly, we can conclude that $\nu_{3,n}\in\mathcal{P}_{2}(\mathbb{R})$ and $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{3,n},\nu_{1})=0.$ Choose $n_{0}\geq c$ large enough such that $\nu:=\nu_{3,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$, then we know that for all $r\geq r_{n_{0}}$,
+Similarly, we can conclude that $\nu_{3,n}\in\mathcal{P}_{2}(\mathbb{R})$ and $\lim_{n\to\infty}\mathcal{W}_{2}(\nu_{3,n},\nu_{1})=0.$ Choose $n_{0}\geq c$ large enough such that $\nu:=\nu_{3,n_{0}}\in B(\nu_{1},\delta)\subset B(\nu_{0},2\delta)$, then we have for all $r\geq r_{n_{0}}$,
 $\begin{split}0<&\mu((-\infty,-r])<2\mu((-\infty,-r])=\nu((-\infty,-r]),\ \text{ and }\\ 0<&\mu([r,\infty))<2\mu([r,\infty))=\nu([r,\infty)).\end{split}$